Specifications and Notation for Lexicon Encoding

COP Project 106 multext-East

Work Package WP1 – Task 1.1 Deliverable D1.1 F

Final Report

29 August 1997

Workpackage Coordinators: CNRS - Nancy Ide and Jean Veronis

Task Leader: LJUBLJANA - Tomaž Erjavec

PISA - Monica Monachini

Contributors:

Bulgarian: R.Pavlov, L.Dimitrova, L.Sinapova, K.Simov Czech: V.Petkevič Estonian: H.J.Kaalep Hungarian: L.Tihanyi

Romanian: D.Tufiş, A.M.Barbu Slovene: T.Erjavec, P.Holozan

Contents

1 Introduction 1 1.1 Background work . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Multext-East work and approach . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Description of the proposal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Lexical lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.5 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6 Organization of the language-specific sections . . . . . . . . . . . . . . . . . . . . 6

2 Morphosyntactic Specifications for Central & Eastern Lexicons 7 2.1 Table of categories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2 Tables of attribute-values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.2.1 Noun (N) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.2 Verb (V) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.2.3 Adjective (A) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.2.4 Pronoun (P) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2.5 Determiner (D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.2.6 Article (T) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.2.7 Adverb (R) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.2.8 Adposition (S) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.2.9 Conjunction (C) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.2.10 Numeral (M) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.2.11 Interjection (I) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.2.12 Residual (X) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.2.13 Abbreviation (Y) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.2.14 Particle (Q) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

2.3 List of values with respective codes . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.4 List of attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 COP project 106 multext-East Deliverable D1.1 FLexicon specifications and notation iii

3 Language Specific Applications 34 3.1 Application to Bulgarian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.2 Application to Czech . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 3.3 Application to Estonian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 3.4 Application to Hungarian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

3.4.1 Derivation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 3.4.2 Compounding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

3.5 Application to Romanian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 3.6 Application to Slovene . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

4 References 165

Chapter 1

Introduction

The purpose of this document is to (i) provide harmonized lexical specifications for the six languages involved in the multext-East project — Bulgarian, Czech, Estonian, Hungarian, Romanian, Slovene — and (ii) formulate the relevant notation to be used in the lexicons contributed by each language group.

The two proposals of lexicon specifications presented in the multext D1-6-1B Deliverable of the multext Project (Bel, Calzolari and Monachini eds. 1995) and in the eagles document of the Lexicon sub-group on Morphosyntactic annotation (Monachini and Calzolari 1995) — which is the basis of the previous one — constitute together the starting point and the model of the work and results presented here.

The multext-East partners have evaluated the two proposals from the point of view of the coverage with respect to their languages, have added the specifications needed to encode the peculiarities of their languages and have produced concrete applications of the proposed specifications. The work has been done through a cyclical process of adjustments and re-application, giving rise to three versions of this document: the initial IM1 version under the coordination from pisa, a subsequent, significantly revised and validated version at project milestone M, and this final version, incorporating the last changes to the language specific tables, the latter two coordinated by ljubljana.

The result is the common proposal for lexicon specifications of the Central & Eastern European languages contained in Chapter 2. On the basis of the eagles and multext models, they are presented as sets of attribute-values — displayed in tabular format — and the notation proposed follows the “string of characters in fixed positions” strategy.

The applications of the proposal to the six languages are given in Chapter 3.They have been contributed respectively by:

Bulgarian: R.Pavlov, L.Dimitrova, L.Sinapova and K.Simov;

Czech: V.Petkevič;

Estonian: H.J.Kaalep;

Hungarian: L.Tihanyi;

Romanian: D.Tufiş and A.M.Barbu; Slovene: T.Erjavec and P.Holozan.

The language application parts present the lexicon specifications category by category and are structured as follows:

(a) section which describes the features and values pertinent to a given category in the form of tables with examples from the language in question;

(b) a second section which provides the combinations of values: the way in which different values combine together, giving rise to all the possible lexicon specifications for the items belonging to that category, is displayed.

Pisa, October 1995 Ljubljana, August 1997

1.1 Background work

The multext D1-6-1B Deliverable (Bel, Calzolari and Monachini eds. 1995) and the eagles proposal of specifications to be encoded in lexica (Monachini and Calzolari 1995) have constituted the basis of the results presented in what follows.

The two documents together collect the outcomes of the main efforts in Europe towards standardization in large linguistic resources and, among other fields, in linguistic annotation. The relationships and the interdependencies between the two actions are reported in the introduction of the eagles document, where the multext contribution to the eagles effort is described in detail and commented also in the introduction of the multext Deliverable. Summing up, the two actions have proceeded with the same bottom-up approach — that is also followed here —, producing similar sets of specifications: the eagles one, aiming at a general description of the languages, is intended to offer a more general proposal and cover a vast range of purposes, while the multext set has to be considered in the context of the project itself, whose aim was to produce resources to be used within the specific application of automatic corpus tagging. The multext set of lexicon specifications is, hence, task-oriented, being designed with a special look to the final purpose.

1.2 multext-East work and approach

The multext-East partners have analysed and evaluated the set of multext specifications with respect to their languages; when the scheme has not been sufficient, then, the eagles inventory has been consulted. If necessary, new distinctions for language-specific information have been proposed.

The work proceeded in a cyclical process: firstly, the partners have circulated the results of their evaluation work in applications that have been, in turn, evaluated by Pisa, and later by Ljubljana. These applications have constituted the basis for the proposal of a set of Central & Eastern specifications.

The nucleus of common features, already isolated within multext, proved to be applicable to the Central & Eastern languages but the selection of distinctions needed for encoding the information peculiar to the multext-East languages and the harmonisation of proposals has required much further work. Some of the features proposed by the partners in their first applications have been dropped, as they pertained to a not-purely morphosyntactic level (i.e. Transitivity); other distinctions have been kept out from the set since considered too fine.

The emerged set has been again circulated among the partners for new cycles of revisions and re-applications, until the specifications have been considered acceptable for all the language groups and stable enough for the first version (IM1) of this deliverable.

However, there were still significant problems with the IM1 this version of the deliverable, in particular:

• different attributes or values were used to describe the same phenomena with different categories / languages (e.g. 1st and 1 for first person);

• especially the three Slavic languages at times described the same phenomena in different

ways;

• language independent aspects of the specifications (e.g. the ’form’ of numerals: digit, roman, etc.) were treated differently for different languages;

• the ordering of attributes in some categories was sub-optimal, often necessitating long morphosyntactic descriptions.

• the attributes and values used different ’punctuation’ (e.g. full art, Modific.Type,

     Pron-Form, SubType;

• the common tables were formated differently for different categories.

Therefore it was decided to produce, for milestone M, version 2 of the deliverable; this effort was led by the Ljubljana site. This harmonisation led to a more motivated and — on the average — more compact morpho-syntactic descriptions for the multext-East languages. The formalisation of the tables and descriptions had an added benefit: a simple Perl program (mtems-expand) was written, which could, working directly with common tables of the morpho-syntactic descriptions of this report, either expand or validate lexical morphosyntactic descriptions. This program was used to validate the word-form lexica of the project, thus ensuring that all the morphosyntactic descriptions in the lexica of the particular languages are well-formed. Another Perl program, mtems-split was also written, which, again working on the common tables, produces language specific tables. These tables were circulated to the partners, thus ensuring that the language specific section do in fact reflect the common tables.

In the tagging of the multext-East corpus, it was found that the initial lexica still contained errors or non-optimal choices; this lead to a revised set of lexica, and this to new language applications of the morpho-syntactic tables. These were incorporated in this Final report.

1.3 Description of the proposal

The proposal has been prepared in the usual multext table format, which displays the specifications (as sets of attribute-values, see below for further details about the notation), with their respective codes used to mark them in the lexicons. Two types of features are distingished:

(i) the minimal core features

These are shared by most of the languages and are highlighted in the tables with asterisks (*). We tried to keep this set in common to all the multext and multext-East languages. In such a way, the comparability across the information encoded in the lexical lists of Central & Eastern and of the six original multext languages is ensured to a certain extent.

• (ii) the multext-East language-specific features

These are shared by the multext-East languages, but differ from the multext ones. The formulation of this set has been, as already mentioned, highly delicate, due also to the fact that many language-specific values were presented in the applications and sometimes the same (or ’similar enough’) morphosyntactic phenomenon was referred to with two different attribute or value names. The phase of recognition and harmonization of semantics of some attributes, values and naming conventions has, hence, required much effort.

If a feature-value is used by only one multext-East language (i.e. if a value is language-specific) it has been marked with l.s.. This marking is used only when a subset of the feature’s values is language specific — in case a whole attribute (i.e. all its features) is language specific, then the mark l.s. is not used.

This representation, with the concrete applications which display and exemplify the attributes and values and provide their internal constraints and relationships, makes the proposal selfexplanatory. Other groups can easily test the specifications on their language, simply by following the method of the applications. The possibility of incorporating idiosyncratic classes and distinctions after the common core features makes the proposal relativelly adaptable and flexible, without ruining compatibility.

1.4 Lexical lists

The specifications presented here have been used in the encoding the word-form lexica of the project. These lexcion contain one entry per line, where an entry has the following form:

Word-form, lemma, morphosyntactic description, (TAG)

This reports describes the morphosyntactic descriptions used in the lexica and the tags developed by the partners. It should be noted that the TAG field is optional in the lexica, partly because it easier and more compact to provide the mapping between the morphosyntactic information and the tags in separate tables and due to the preliminary nature of the tagsets developed. In other words, the lexica with the morphosyntactic descriptions provide a much more mature resource than do the tags.

1.5 Notation

In multext, the notation has been chosen following current practices for NLP, where information is represented in attribute-value formalisms and following the idea that it should also be self-informative for human understanding. At the same time, a relativelly compact encoding was maintained. The notation format suggested has the following main characteristics:

• attributes are marked by positions;

• values are represented by a single character;

• a special marker reflects the non applicability of a given attribute.

These characteristics make the proposed lexical notation similar to attribute-value pairs used in unification based formalisms (see the multext D1-6-1B Deliverable for further details).

The linear strings of characters representing the morphosyntactic descriptions are constructed following the philosophy of the Intermediate Format proposed in the eagles Corpus proposal (Leech and Wilson, 1994), i.e. of having agreed symbols in predefined and fixed positions: the positions of a string of characters are numbered 0, 1, 2, etc. in the following way:

a. the agreed character at position 0 encodes part-of-speech;

b. each character at position 1, 2, n, encodes the value of one attribute (person, gender, number, etc.);

c. if an attribute does not apply, the corresponding position in the string contains a special marker, in our case ’-’ (hyphen).

Example: Ncms- (Noun, common, masculine, singular, nocase)

This notation adopts the eagles Intermediate Format with a small revision: the Intermediate Format encodes information by means of digits, while in multext characters of a mnemonic nature are preferred.

The marker ’-’ has a special semantics and it means ’not-applicable’. As stated above, its function is just to keep the relationship established between attributes and values. It might be used for the following cases:

(a) not relevant to a particular language, e.g. Gender to Estonian.

(b) not applicable to a particular combination of attribute-values, i.e. although the attribute is used by a category in a given language it does not apply to a particular subclass of the category; e.g., Person applies to Pronouns, but not to the Type demonstrative.

(c) not applicable to a particular lexical item, i.e. although the attribute applies to the rest of its paradigm, e.g., Gender in the paradigm of English Personal Pronouns applies only to the 3rd person, I, you vs. she, he).

1.6 Organization of the language-specific sections

Following the procedure of the multext Deliverable, the language-specific sections consist of two distinct parts:

• a descriptive section, where the features and values relevant to the lexicon of the language are displayed in tabular form, with examples taken from the language;

• the combination section, where the way in which the different values combine together is shown.

Two different strategies of displaying combinations have been followed:

(i) all the admitted combinations are provided together with an example. This has the disadvantage of producing big lists, given the high number of features and values to combine together, but the advantage of providing the only legal combinations with the relevant constraints in the application of some features/values in presence of other features or values or combination of them (see e.g Gender in Personal Pronouns).

(ii) a mathematical expression describing the combinations is provided, which can subsequently used to generate all combinations, including those which are not valid.

Morphosyntactic Specifications for Central & Eastern Lexicons

The categories listed below with the relevant attributes and values are based on the multext and eagles documents. The specifications constitute a “harmonized” set of features which properly describe lexical items of the different languages.

In keeping with the general aim of the project, these harmonized specifications — and the related resources — will contribute to the standardization of the corpus annotation work. This common set of features, besides the advantages described in the Introduction, will also be a common ground to perform comparisons of different annotation tool results, because the existence of many lexical description systems is causing nowadays a problem for comparing results.

Therefore the categories and features listed below are the common reference for the work done by the different groups.

2.1 Table of categories

=============== ====
Part-of-Speech  Code
=============== ====
Noun             N
Verb             V
Adjective        A
Pronoun          P
Determiner       D
Article          T
Adverb           R
Adposition       S
Conjunction      C
Numeral          M
Interjection     I
Residual         X
Abbreviation     Y
Particle         Q
=============== ====

Each category acts a type, i.e. it introduces attributes and their values. The equivalence of a particular attribute-value at a particular position in a morpho-syntactic description is thus defined only with respect to a given category. The attributes and the values pertinent to the above categories are presented in the next section.

After the section devoted to the tables, another section presents all the values listed in alphabetical order and provides the respective codes. A further section contains the list of the attributes, providing the information of the category/categories in which they are used, and for some of them, a very synthetic definition of their semantics.

2.2 Tables of attribute-values

The common multext-East tables of attribute values are given for all categories above, and have a rigid structure, which makes them suitable for automatically verifying the conformance of a particular morphosyntactic description with the tables, or for expanding a morphosyntactic description into its more verbose form.

At the beginning of each category the number of attributes introduced by the category is given. This is followed by a string summarizing the valid attributes of the category. The features constituting the minimal core, in common with multext (and eagles) are graphically contained between strings of stars (*).

This information is followed by the table proper. The table has the following columns:

1. Position gives the position of the attribute in the string of the morphosyntactic description;

2. Attribute gives the name of the attribute;

3. Value gives the name and one-letter code of the attribute-value;

4. Ro, Sl, Cs, Bg, Et, Hu i.e. a column for each of the multext-East languages. For easier comparison between them they have been grouped by language family: first Romance (Romanian), then Slavic (Bulgarian, Czech, Slovene), and finaly Finno-Ungric (Estonian, Hungarian).

The language columns define (by marking with ’x’), in the first line, whether the category is used by the language, and in subsequent lines, which attribute-values a particular language uses.

2.2.1 Noun (N)

11 Positions
**** **** **** **** **** ---- ---- ---- ---- ---- ----
PoS  Type Gend Numb Case Def  Cltc Anim OwnN OwnP OwdN
**** **** **** **** **** ---- ---- ---- ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           common         c  x   x   x   x   x   x
                 proper         p  x   x   x   x   x   x
- -------------- -------------- -
2 Gender         masculine      m  x   x   x   x
                 feminine       f  x   x   x   x
                 neuter         n  x   x   x   x
- -------------- -------------- -
3 Number         singular       s  x   x   x   x   x   x
                 plural         p  x   x   x   x   x   x
                 dual           d      x   x
          l.s.   count          t              x

COP project 106 multext-East  Deliverable D1.1 F— Common Tables  11

- -------------- -------------- -
4 Case           nominative     n      x   x   x   x   x
genitive       g      x   x       x   x
dative         d      x   x           x
accusative     a      x   x           x
vocative       v  x       x   x
locative       l      x   x
instrumental   i      x   x           x
l.s.  direct         r  x
l.s.  oblique        o  x
l.s.  partitive      1                  x
illative       x                  x   x
inessive       2                  x   x
elative        e                  x   x
allative       t                  x   x
adessive       3                  x   x
ablative       b                  x   x
l.s.  translative    4                  x
terminative    9                  x   x
essive         w                  x   x
l.s.  abessive       5                  x
l.s.  komitative     k                  x
l.s.  aditive        7                  x
l.s.  temporalis     m                      x
l.s.  causalis       c                      x
l.s.  sublative      s                      x
l.s.  delative       h                      x
l.s.  sociative      q                      x
l.s.  factive        y                      x
l.s.  superessive    p                      x
l.s.  distributive   u                      x
* ***************************** *
5 Definiteness   no             n  x           x
yes            y  x           x
l.s.  short_art      s              x
l.s.  full_art       f              x
- -------------- -------------- -
6 Clitic         no             n  x
yes            y  x
- -------------- -------------- -
7 Animate        no             n          x           x
yes            y          x           x
- -------------- -------------- -
8 Owner_Number   singular       s                      x
plural         p                      x
- -------------- -------------- -
9 Owner_Person   first          1                      x
second         2                      x
third          3                      x
---------------- -------------- -
10Owned_Number   singular       s                      x
plural         p                      x
=================================

2.2.2 Verb (V)

14 Positions
**** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ----
PoS  Type VFrm Tens Pers Numb Gend Voic Neg  Def  Cltc Case Anim Clt2
**** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           main           m  x   x   x   x   x   x
                 auxiliary      a  x   x   x   x   x   x
                 modal          o  x   x   x       x
                 copula         c  x   x   x
- -------------- -------------- -
2 VForm          indicative     i  x   x   x   x   x   x
                 subjunctive    s  x
                 imperative     m  x   x   x   x   x   x
                 conditional    c      x   x       x   x
                 infinitive     n  x   x   x       x   x
                 participle     p  x   x   x   x   x
                 gerund         g  x           x   x
                 supine         u      x           x
           l.s.  transgressive  t          x
           l.s.  quotative      q                  x
- -------------- -------------- -
3 Tense          present        p  x   x   x   x   x   x
                 imperfect      i  x           x   x
                 future         f      x   x
                 past           s  x   x   x   x   x   x
           l.s.  pluperfect     l  x
           l.s.  aorist         a              x
- -------------- -------------- -
4 Person         first          1  x   x   x   x   x   x
                 second         2  x   x   x   x   x   x
                 third          3  x   x   x   x   x   x
- -------------- -------------- -
5 Number         singular       s  x   x   x   x   x   x
                 plural         p  x   x   x   x   x   x
           l.s.  dual           d      x
- -------------- -------------- -
6 Gender         masculine      m  x   x   x   x
                 feminine       f  x   x   x   x
                 neuter         n  x   x   x   x
*********************************
7 Voice          active         a      x   x   x   x
                 passive        p      x   x   x   x
- -------------- -------------- -
8 Negative       no             n      x   x       x
                 yes            y      x   x       x
- -------------- -------------- -
9 Definiteness   no             n              x       x

yes            y              x       x
short_art      s              x
full_art       f              x
l.s.  1s2s           2                      x
- -------------- -------------- -
10Clitic         no             n  x
yes            y  x
- -------------- -------------- -
11Case           nominative     n
genitive       g
dative         d
accusative     a
locative       l
instrumental   i
illative       x                  x
inessive       2                  x
elative        e                  x
translative    4                  x
abessive       5                  x
- -------------- -------------- -
12Animate        no             n          x
yes            y          x
- -------------- -------------- -
13Clitic_s       no             n          x
yes            y          x
=================================

2.2.3 Adjective (A)

13 Positions
**** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ----
PoS  Type Degr Gend Numb Case Def  Cltc Anim Form OwnN OwnP OwdN
**** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           qualificative  f  x   x   x           x
indefinite     i
possessive     s      x   x
l.s.  ordinal        o      x
- -------------- -------------- -
2 Degree         positive       p  x   x   x       x   x
comparative    c  x   x   x       x   x
superlative    s  x   x   x       x   x
- -------------- -------------- -
3 Gender         masculine      m  x   x   x   x
feminine       f  x   x   x   x
neuter         n  x   x   x   x
- -------------- -------------- -
4 Number         singular       s  x   x   x   x   x   x
plural         p  x   x   x   x   x   x
dual           d      x   x
- -------------- -------------- -
5 Case           nominative     n      x   x       x   x
genitive       g      x   x       x   x
dative         d      x   x           x
accusative     a      x   x           x
vocative       v  x       x
locative       l      x   x
instrumental   i      x   x           x
l.s.  direct         r  x
l.s.  oblique        o  x
l.s.  partitive      1                  x
illative       x                  x   x
inessive       2                  x   x
elative        e                  x   x
allative       t                  x   x
adessive       3                  x   x
ablative       b                  x   x
l.s.  translative    4                  x
terminative    9                  x   x
essive         w                  x   x
l.s.  abessive       5                  x
l.s.  komitative     k                  x
l.s.  aditive        7                  x
l.s.  temporalis     m                      x
l.s.  causalis       c                      x
l.s.  sublative      s                      x

           l.s.  delative       h                      x
           l.s.  sociative      q                      x
           l.s.  factive        y                      x
           l.s.  superessive    p                      x
           l.s.  distributive   u                      x
           l.s.  essive_formal  f                      x
* ***************************** *
6 Definiteness   no             n  x           x
                 yes            y  x           x
           l.s.  short_art      s              x
           l.s.  full_art       f              x
- -------------- -------------- -
7 Clitic         no             n  x
                 yes            y  x
- -------------- -------------- -
8 Animate        no             n          x           x
                 yes            y          x           x
- -------------- -------------- -
9 Formation      nominal        n          x
                 compound       c          x
- -------------- -------------- -
10Owner_Number   singular       s                      x
                 plural         p                      x
- -------------- -------------- -
11Owner_Person   first          1                      x
                 second         2                      x
                 third          3                      x
---------------- -------------- -
12Owned_Number   singular       s                      x
                 plural         p                      x
=================================

2.2.4 Pronoun (P)

17 Positions
**** **** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ...
PoS  Type Pers Gend Numb Case OwnN OwnG Cltc RefT SynT Def  Anim Clt2
**** **** **** **** **** **** **** **** ---- ---- ---- ---- ---- ---- ...
... ---- ---- ----
PrFr OwnP OwdN
... ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           personal       p  x   x   x   x   x   x
demonstrative  d  x   x   x   x   x   x
indefinite     i  x   x   x   x   x   x
possessive     s  x   x   x   x   x   x
interrogative  q      x   x   x   x   x
relative       r      x   x   x   x   x
l.s.  exclamative    e
reflexive      x  x   x   x   x   x   x
reciprocal     y                  x   x
l.s.  int_rel        w  x
negative       z  x   x   x   x
l.s   emphatic       h  x
l.s.  determinal     m                  x
general        g      x   x   x
- -------------- -------------- -
2 Person         first          1  x   x   x   x   x   x
second         2  x   x   x   x   x   x
third          3  x   x   x   x   x   x
- -------------- -------------- -
3 Gender         masculine      m  x   x   x   x
feminine       f  x   x   x   x
neuter         n  x   x   x   x
- -------------- -------------- -
4 Number         singular       s  x   x   x   x   x   x
plural         p  x   x   x   x   x   x
dual           d      x   x
- -------------- -------------- -
5 Case           nominative  *  n  x1  x   x   x   x   x
genitive    *  g  x2  x   x       x   x
dative      *  d  x2  x   x   x       x
accusative  *  a  x1  x   x   x       x
vocative       v  x
locative       l      x   x
instrumental   i      x   x           x
l.s.  direct      *  r  xR1
l.s.  oblique     *  o  xR2
l.s.  partitive      1                  x
illative       x                  x   x
inessive       2                  x   x

elative        e                  x   x
allative       t                  x   x
adessive       3                  x   x
ablative       b                  x   x
l.s.  translative    4                  x
terminative    9                  x   x
essive         w                  x   x
l.s.  abessive       5                  x
l.s.  komitative     k                  x
l.s.  aditive        7                  x
l.s.  temporalis     m                      x
l.s.  causalis       c                      x
l.s.  sublative      s                      x
l.s.  delative       h                      x
l.s.  sociative      q                      x
l.s.  factive        y                      x
l.s.  superessive    p                      x
l.s.  distributive   u                      x
l.s.  essive_formal  f                      x
- -------------- -------------- -
6 Owner_Number   singular       s  x   x   x           x
plural         p  x   x   x           x
l.s.  dual           d      x
- -------------- -------------- -
7 Owner_Gender   masculine      m      x   x
feminine       f      x   x
neuter         n      x   x
*********************************
8 Clitic         no             n  x   x   x   x
yes            y  x   x   x   x
- -------------- -------------- -
9 Referent_Type  personal       p      x   x   x
possessive     s      x   x   x
attributive    a              x
quantitative   q              x
- -------------- -------------- -
10Syntactic_Type nominal        n      x   x
adjectival     a      x   x
- -------------- -------------- -
11Definiteness   no             n              x
yes            y              x
short_art      s              x
full_art       f              x
- -------------- -------------- -
12Animate        no             n          x
yes            y          x
- -------------- -------------- -
13Clitic_s       yes            y          x
no             n          x
- -------------- -------------- -
14Pronoun_Form   strong         s  x
weak           w  x
- -------------- -------------- -

15Owner_Person   first          1                      x
                 second         2                      x
                 third          3                      x
- -------------- -------------- -
16Owned_Number   singular       s                      x
                 plural         p                      x
=================================

Notes

* In the Romanian case system the value ’direct’ conflates ’nominative’ and ’accusative’, while the value ’oblique’ conflates ’genitive’ and ’dative’.

If we want to list these values as belonging to the same Case universe, i.e. if we want either ’direct’, ’nominative’ and ’accusative’ or ’oblique’, ’genitive’ and ’dative’ appear all together in the same system, their semantics should be explained and their internal relationships represented. We suggest here to follow the same strategy adopted within the eagles validation exercise, where different kinds of relationships between values have been formally represented by adding indexes to the values of interest:

- in this example, the kind of relation ’direct’ and ’oblique’ have with other values of the system is of ’replacement’. We mark this labelling ’direct’ and ’oblique’ with the index ’R’ and adding the numerical indexes 1 to ’direct’ and 2 to ’oblique’, which point to the values which are respectively replaced.

2.2.5 Determiner (D)

10 Positions
**** **** **** **** **** **** **** **** ---- ----
PoS  Type Pers Gend Numb Case OwnN OwnG Cltc Mod
**** **** **** **** **** **** **** **** ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   -   -   -   -   -
= ============== ============== =
1 Type           demonstrative  d  x
                 indefinite     i  x
                 possessive     s  x
                 interrogative  q
                 relative       r
                 exclamative    e
                 article        a
           l.s.  int_rel        w  x
           l.s.  negative       z  x
           l.s.  emphatic       h  x
- -------------- -------------- -
2 Person         first          1  x
                 second         2  x
                 third          3  x
- -------------- -------------- -
3 Gender         masculine      m  x
                 feminine       f  x
                 neuter         n  x
- -------------- -------------- -
4 Number         singular       s  x
                 plural         p  x
- -------------- -------------- -
5 Case           direct         r  x
                 oblique        o  x
- -------------- -------------- -
6 Owner_Number   singular       s  x
                 plural         p  x
- -------------- -------------- -
7 Owner_Gender   masculine      m
                 feminine       f
                 neuter         n
*********************************
8 Clitic         no             n  x
                 yes            y  x
- -------------- -------------- -
9 Modific_Type   prenomin       e  x
                 postnomin      o  x
=================================

2.2.6 Article (T)

6 Positions
**** **** **** **** **** ----
PoS  Type Gend Numb Case Cltc
**** **** **** **** **** ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   -   -   -   -   x
= ============== ============== =
1 Type           definite       f  x                   x
                 indefinite     i  x                   x
           l.s.  possessive     s  x
           l.s.  demonstrative  d  x
- -------------- -------------- -
2 Gender         masculine      m  x
                 feminine       f  x
                 neuter         n  x
- -------------- -------------- -
3 Number         singular       s  x
                 plural         p  x
- -------------- -------------- -
4 Case     l.s.  direct         r  x
           l.s.  oblique        o  x
*********************************
5 Clitic         no             n  x
                 yes            y  x
=================================

2.2.7 Adverb (R)

6 Positions
**** **** **** ---- ---- ----
PoS  Type Degr Cltc Numb Pers
**** **** **** ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           general        g  x   x   x   x       x
                 particle       p  x                   x
           l.s.  causal         o                      x
           l.s.  negative       z  x
           l.s.  adjectival     a              x
           l.s.  verbal         v                      x
                 modifier       m  x                   x
           l.s.  int_rel        w  x
           l.s.  interrogative  q                      x
- -------------- -------------- -
2 Degree         positive       p  x   x   x
                 comparative    c  x   x   x
                 superlative    s  x   x   x
*********************************
3 Clitic         no             n  x                   x
                 yes            y  x                   x
- -------------- -------------- -
4 Number         singular       s                      x
                 plural         p                      x
- -------------- -------------- -
5 Person         first          1                      x
                 second         2                      x
                 third          3                      x
=================================

2.2.8 Adposition (S)

5 Positions
**** **** **** ---- ----
PoS  Type Form Case Cltc
**** **** **** ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           preposition    p  x   x   x   x   x
                 postposition   t                  x   x
- -------------- -------------- -
2 Formation      simple         s  x   x   x
                 compound       c  x   x   x
*********************************
3 Case           nominative     n
  (req.by prep.) genitive       g  x   x   x
                 dative         d  x   x   x
                 accusative     a  x   x   x
                 locative       l      x   x
                 instrumental   i      x   x
- -------------- -------------- -
4 Clitic         no             n  x
                 yes            y  x
=================================

2.2.9 Conjunction (C)

8 Positions
**** **** ---- ---- ---- ---- ---- ----
PoS  Type Form CTyp SubT Cltc Numb Pers
**** **** ---- ---- ---- ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           coordinating   c  x   x   x   x   x   x
                 subordinating  s  x   x   x   x   x   x
*********************************
2 Formation      simple         s  x   x       x       x
                 compound       c  x   x       x       x
- -------------- -------------- -
3 Coord_Type     simple         s  x
                 repetit        r  x
                 correlat       c  x
                 sentence       p                      x
                 words          w                      x
- -------------- -------------- -
4 Sub_Type       negative       z  x
                 positive       p  x
- -------------- -------------- -
5 Clitic         no             n  x
                 yes            y  x
- -------------- -------------- -
6 Number         singular       s          x
                 plural         p          x
- -------------- -------------- -
7 Person         first          1          x
                 second         2          x
                 third          3          x
=================================

2.2.10 Numeral (M)

13 Positions
**** **** **** **** **** ---- ---- ---- ---- ---- ---- ---- ----
PoS  Type Gend Numb Case Frm  Def  Cltc Clas Anim OwnN OwnP OwdN
**** **** **** **** **** ---- ---- ---- ---- ---- ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           cardinal       c  x   x   x   x   x   x
ordinal        o  x   x   x   x   x   x
fractal        f  x                   x
multiple       m  x   x   x
collect        l  x                   x
special        s      x   x
- -------------- -------------- -
2 Gender         masculine      m  x   x   x   x
feminine       f  x   x   x   x
neuter         n  x   x   x   x
- -------------- -------------- -
3 Number         singular       s  x   x   x   x   x   x
plural         p  x   x   x   x   x   x
dual           d      x   x
- -------------- -------------- -
4 Case           nominative     n      x   x       x   x
genitive       g      x   x       x   x
dative         d      x   x           x
accusative     a      x   x           x
locative       l      x   x
instrumental   i      x   x           x
l.s.  direct         r  x
l.s.  oblique        o  x
l.s.  partitive      1                  x
illative       x                  x   x
inessive       2                  x   x
elative        e                  x   x
allative       t                  x   x
adessive       3                  x   x
ablative       b                  x   x
l.s.  translative    4                  x
terminative    9                  x   x
essive         w                  x   x
l.s.  abessive       5                  x
l.s.  komitative     k                  x
l.s.  aditive        7                  x
l.s.  temporalis     m                      x
l.s.  causalis       c                      x
l.s.  sublative      s                      x
l.s.  delative       h                      x
l.s.  sociative      q                      x
l.s.  factive        y                      x

l.s.  superessive    p                      x
l.s.  distributive   u                      x
l.s.  essive_formal  f                      x
l.s.  multiplicative 6                      x
*********************************
5 Form           digit          d  x   x   x   x   x   x
roman          r  x   x   x   x   x   x
letter         l  x   x   x   x   x   x
l.s.  both           b  x
l.s.  m-form         m              x
l.s.  approx         a              x
- -------------- -------------- -
6 Definiteness   no             n  x           x
yes            y  x           x
short_art      s              x
full_art       f              x
- -------------- -------------- -
7 Clitic         no             n  x
yes            y  x
- -------------- -------------- -
8 Class          definite1      1          x
definite2      2          x
definite34     3          x
definite       f          x
demonstrative  d          x
indefinite     i          x
interrogative  q          x
relative       r          x
- -------------- -------------- -
9 Animate        no             n          x
yes            y          x
- -------------- -------------- -
10Owner_Number   singular       s                      x
plural         p                      x
- -------------- -------------- -
11Owner_Person   first          1                      x
second         2                      x
third          3                      x
- -------------- -------------- -
12Owned_Number   singular       s                      x
plural         p                      x
=================================

2.2.11 Interjection (I)

3 Positions
**** ---- ----
PoS  Type Form
**** ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Type           mood           m                      x
                 other          o                      x
- -------------- -------------- -
2 Formation      simple         s              x
                 compound       c              x
=================================

2.2.12 Residual (X)

1 Position
****
PoS
****
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
=================================

2.2.13 Abbreviation (Y)

6 Positions
**** ---- ---- ---- ---- ----
PoS  SynT Gend Numb Case Def
**** ---- ---- ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   x   x
= ============== ============== =
1 Syntactic_Type nominal        n  x               x
                 verbal         v  x               x
                 adjectival     a  x               x
                 adverbial      r  x               x
- -------------------------------
2 Gender         masculine      m  x
                 feminine       f  x
                 neuter         n  x
- -------------- -------------- -
3 Number         singular       s  x               x
                 plural         p  x               x
- -------------- -------------- -
4 Case     l.s.  nominative     n                  x
           l.s.  genitive       g                  x
           l.s.  direct         r  x
           l.s.  oblique        o  x
           l.s.  partitive      1                  x
           l.s.  illative       x                  x
           l.s.  inessive       2                  x
           l.s.  elative        e                  x
           l.s.  allative       t                  x
           l.s.  adessive       3                  x
           l.s.  ablative       b                  x
           l.s.  translative    4                  x
           l.s.  terminative    9                  x
           l.s.  essive         w                  x
           l.s.  abessive       5                  x
           l.s.  komitative     k                  x
           l.s.  aditive        7                  x
- -------------- -------------- -
5 Definiteness   yes            y  x
                 no             n  x
=================================

2.2.14 Particle (Q)

5 Positions
**** ---- ---- ----
PoS  Type Form Cltc
**** ---- ---- ----
= ============== ============== =  RO  SL  CS  BG  ET  HU
P ATT            VAL            C  x   x   x   x   -   -
= ============== ============== =
1 Type           negative       z  x           x
                 infinitive     n  x
                 subjunctive    s  x
                 aspect         a  x
                 future         f  x
                 general        g              x
                 comparative    c              x
                 verbal         v              x
                 interrogative  q              x
                 modal          o              x
- -------------- -------------- -
2 Formation      simple         s              x
                 compound       c              x
- -------------- -------------- -
3 Clitic         no             n  x
                 yes            y  x
=================================

2.3 List of values with respective codes

The values presented within the tables are, in the following, listed in alphabetical order; the first column gives the name of the value, the second column its code and the third lists attributes for which the value is appropriate.

1s2s           2    Definiteness
abessive       5    Case
ablative       b    Case
accusative     a    Case
active         a    Voice
adessive       3    Case
aditive        7    Case
adjectival     a    Syntactic_Type, Type
allative       t    Case
aorist         a    Tense
approx         a    Form
article        a    Type
aspect         a    Type
attributive    a    Referent_Type
auxiliary      a    Type
both           b    Form
cardinal       c    Type
causal         o    Type
causalis       c    Case
collect        l    Type
common         c    Type
comparative    c    Degree, Type
compound       c    Formation
conditional    c    VForm
coordinating   c    Type
copula         c    Type
correlat       c    Coord_Type
count          t    Number
dative         d    Case
definite       f    Class, Type
definite1      1    Class
definite2      2    Class
definite34     3    Class
delative       h    Case
demonstrative  d    Class, Type
determinal     m    Type
digit          d    Form
direct         r    Case
distributive   u    Case
dual           d    Number, Owner_Number
elative        e    Case
emphatic       h    Type
essive         w    Case
essive_formal  f    Case
exclamative    e    Type
factive        y    Case
feminine       f    Gender, Owner_Gender
first          1    Owner_Person, Person
fractal        f    Type
full_art       f    Definiteness
future         f    Tense, Type
general        g    Type
genitive       g    Case
gerund         g    VForm
illative       x    Case
imperative     m    VForm
imperfect      i    Tense
indefinite     i    Class, Type

indicative     i    VForm
inessive       2    Case
infinitive     n    Type, VForm
instrumental   i    Case
int_rel        w    Type
interrogative  q    Class, Type
komitative     k    Case
letter         l    Form
locative       l    Case
main           m    Type
masculine      m    Gender, Owner_Gender
modal          o    Type
modifier       m    Type
mood           m    Type
multiple       m    Type
multiplicative 6    Case
negative       z    Sub_Type, Type
neuter         n    Gender, Owner_Gender
no             n    Animate, Clitic, Clitic_s, Definiteness, Negative
nominal        n    Formation, Syntactic_Type
nominative     n    Case
oblique        o    Case
ordinal        o    Type
other          o    Type
participle     p    VForm
particle       p    Type
partitive      1    Case
passive        p    Voice
past           s    Tense
personal       p    Referent_Type, Type
pluperfect     l    Tense
plural         p    Number, Owned_Number, Owner_Number
positive       p    Degree, Sub_Type
possessive     s    Referent_Type, Type
postnomin      o    Modific_Type
postposition   t    Type
prenomin       e    Modific_Type
preposition    p    Type
present        p    Tense
proper         p    Type
qualificative  f    Type
quantitative   q    Referent_Type
quotative      q    VForm
reciprocal     y    Type
reflexive      x    Type
relative       r    Class, Type
repetit        r    Coord_Type
roman          r    Form
second         2    Owner_Person, Person
sentence       p    Coord_Type
short_art      s    Definiteness
simple         s    Coord_Type, Formation
singular       s    Number, Owned_Number, Owner_Number
sociative      q    Case
special        s    Type
strong         s    Pronoun_Form
subjunctive    s    Type, VForm
sublative      s    Case
subordinating  s    Type
superessive    p    Case
superlative    s    Degree
supine         u    VForm
temporalis     m    Case
terminative    9    Case
third          3    Owner_Person, Person
transgressive  t    VForm

translative    4    Case
verbal         v    Type
vocative       v    Case
weak           w    Pronoun_Form
words          w    Coord_Type
yes            y    Animate, Clitic, Clitic_s, Definiteness, Negative

2.4 List of attributes

In this section, all the attributes presented in the tables are listed in alphabetical order. For some attributes which are not self-explanatory, a brief description of their semantics is provided.

Attribute  Relevant to Category
Animate  Adj Noun Num Pron Verb
Case  Abbr Adj Adpos Art Det Noun Num Pron Verb
Class  Num
Clitic  Adj Adpos Adv Art Conj Det Noun Num Part Pron Verb
Clitic s  Pron
Coord Type  Conj
Definiteness  Abbr Adj Noun Num Pron Verb
Degree  Adj Adv
Form  Num
Formation  Adj Adpos Conj Interj Part
Gender  Abbr Adj Art Det Noun Num Pron Verb
Modific Type  Det
Negative  Verb
Number  Abbr Adj Adv Art Conj Det Noun Num Pron Verb
Owned Number Adj Noun Num Pron
Owner Gender Det Pron
Owner Number Adj Det Noun Num Pron
Owner Person Adj Noun Num Pron
Person  Adv Conj Det Pron Verb
Pronoun Form Pron
Referent Type Pron
Sub Type  Conj
Syntactic Type Abbr Pron
Tense  Verb
Type  Adj Adpos Adv Art Conj Det Interj Noun Num Part Pron Verb
VForm  Verb
Voice  Verb

Class: distinguishes subtypes of Numerals in Czech which have a distinct syntactic distributions: e.g. subclasses for 1, 2, 3&4, etc. are distinguished.

Clitic s: the ’yes’ value of the Clitic s attribute denotes Czech pronouns having the clitic morpheme ’s’ appended as a suffix.

Definiteness: corresponds to the definite and indefinite article in English, which is expressed in Bulgarian by suffixes; these differ according to the gender and number of a word. For singular masculine there are two forms: full article and short article (full is used when a sing.masc. form is the syntactic subject of the clause, otherwise short article is used). The distinction full vs. short is not made for feminine, neuter and plural forms. Definiteness is also used in Romanian.

Form: used to distinguish different forms of numerals (Roman, digit, ’letter’) and in Bulgarian also for the semantic distinction between numerals that refer to male persons (but not children) or groups of male+female.

Formation: refers to the graphical components: simple, i.e. consisting of one word; compound, i.e. consisting of more than one word.

Modific Type: refers to the prenominal or postnominal positions of Determiners which distinguish different forms in Romanian.

Negative: the value ’yes’ encodes negative verbal word-forms in Czech, Slovene and in Estonian.

Owned Number: in the Hungarian system, different word-forms are distinguished for nominals on the basis of so called ’anaphoric possessive’ number, i.e. the number of the thing(s) possessed by the nominal in question.

Owner Gender: used to encode the Gender of the possessor in Pronouns and (in Romanian) Determiners.

Owner Number: used to specify the possessor number in Pronouns, as well as (in Romanian) in Determiners, and (in Hungarian) in Adjectives and Nouns.

Owner Person: used to specify the possessor person in in Hungarian in Adjectives and Nouns.

Pronoun Form: used to encode weak and strong pronouns in Romanian.

Referent Type: used to distinguish reflexive personal from reflexive possessive pronouns in the Slavic languages. In Bulgarian, it also describes a subdivision on the basis of semantic features which have effect on the morpho-syntactic paradigm, e.g. quantitative: the pronoun refers to quantity, etc.

Sub Type: used in Romanian to distinguish negative from positive Conjunctions.

Syntactic Type: used to distinguish the nominal and adjectival function of Pronouns in Slovene and Czech. Also used in Abbreviations to signal the Part of Speech of the abbreviation; currently used only by Romanian and Estonian.

Language Specific Applications

3.1 Application to Bulgarian

The application to Bulgarian has been carried out by Radoslav Pavlov and Ludmila Dimitrova (IM-BAS), Lydia Sinapova (IIT-BAS), and Kiril Iv. Simov (LML-BAS).

1. Noun (N)
= ============== ============== ============= =
P ATT            VAL              Example     C
= ============== ============== ============= =
1 Type           common           kniga       c
                 proper           Ivan        p
- -------------- -------------- ------------- -
2 Gender         masculine        stol        m
                 feminine         masa        f
                 neuter           vreteno     n
- -------------- -------------- ------------- -
3 Number         singular         momtche     s
                 plural           stolove     p
          l.s.   count            (dva) stola t
- -------------- -------------- ------------- -
4 Case           nominative       narod       n
                 vocative         narode      v
* ***************************** ************* *
5 Definiteness   no               utchitel    n
                 yes              zhenata     y
           l.s.  short_art        moliva      s
           l.s.  full_art         molivyt     f
- -------------- -------------- ------------- -
6 Clitic                            -         -
- -------------- -------------- ------------- -
7 Animate                           -         -
- -------------- -------------- ------------- -
8 Owner_Number                      -         -
- -------------- -------------- ------------- -
9 Owner_Person                      -         -

COP project 106 multext-East  Deliverable D1.1 F— Bulgarian  36

---------------- -------------- ------------- -
10Owned_Number                      -         -
=============================== ============= =

Combinations
=== ====== ====== ===== ==== === ==== ==== ==== ==== ==== =======
POS  Type   Gend   Numb Case Def Clit Anim OwnN OwnP OwdN Example
=== ====== ====== ===== ==== === ==== ==== ==== ==== ==== =======
N    [cp]  [mfn]   [sp]  -    n    -   -    -    -    -       1.
N    [cp]    m      s    -   [sf]  -   -    -    -    -       2.
N    [cp]  [mf]     -    v    -    -   -    -    -    -       3.
N    [cp]  [mfn]    p    -    y    -   -    -    -    -       4.
N    [cp]  [fn]     s    -    y    -   -    -    -    -       5.
N     c      m      t    -    -    -   -    -    -    -       6.
=== ====== ======= ==== ==== === ==== ==== ==== ==== ==== =======
Examples:
1. narod, zhena, selo, Ivan, Penka
2. naroda, narodyt
3. narode, zheno
4. narodite, zhenite, selata
5. zhenata, seloto
6. naroda

2. Verb (V)
= ============== ============== ============== =
P ATT            VAL             Example       C
= ============== ============== ============== =
1 Type           main            govorya       m
                 auxiliary       sym           a
- -------------- -------------- -------------- -
2 VForm          indicative      govorya       i
                 imperative      govorete      m
                 participle      govoril       p
                 gerund          govorejki     g
- -------------- -------------- -------------- -
3 Tense          present         govorya       p
                 imperfect       govoreh       i
                 past            govoreno      s
           l.s.  aorist          govorih       a
- -------------- -------------- -------------- -
4 Person         first           govorya       1
                 second          govorish      2
                 third           govori        3
- -------------- -------------- -------------- -
5 Number         singular        govorya       s
                 plural          govoryat      p
- -------------- -------------- -------------- -

6 Gender         masculine       govoril       m
                 feminine        govorila      f
                 neuter          govorilo      n
************************************************
7 Voice          active          govorest      a
                 passive         govoreno      p
- -------------- -------------- -------------- -
8 Negative                                     -
- -------------- -------------- -------------- -
9 Definiteness   no             govoril        n
                 yes            govorilite     y
                 short_art      govoriliya     s
                 full_art       govoriliyat    f
- -------------- -------------- -------------- -
10Clitic                                       -
- -------------- -------------- -------------- -
11Case                                         -
- -------------- -------------- -------------- -
12Animate                                      -
- -------------- -------------- -------------- -
13Clitic_s                                     -
============================================== =

Combinations
=== ===== ==== ==== ==== ===== ====  === === === === ==== == === =======
PoS  Type VFrm Tens Pers Numb  Gend Voic Neg Def Cl1 Case An Cl2 Example
=== ===== ==== ==== ==== ===== ====  === === === === ==== == === =======
V    [ma]  i   [pai][123] [sp]   -    -  -    -   -   -    -  -    1.
V    [ma]  m    -     2   [sp]   -    -  -    -   -   -    -  -    2.
V     m    p   [pai]  -    s    [mfn] a  -    n   -   -    -  -    3.
V     m    p   [pa]   -    s     m    a  -   [sf] -   -    -  -    4.
V     m    p   [pa]   -    s    [fn]  a  -    y   -   -    -  -    5.
V     m    p   [pai]  -    p     -    a  -    n   -   -    -  -    6.
V     m    p   [pa]   -    p     -    a  -    y   -   -    -  -    7.
V     m    p    s     -    s    [mfn] p  -    n   -   -    -  -    8.
V     m    p    s     -    s     m    p  -   [sf] -   -    -  -    9.
V     m    p    s     -    s    [fn]  p  -    y   -   -    -  -   10.
V     m    p    s     -    p     -    p  -   [ny] -   -    -  -   11.
V     a    p   [ai]   -    s    [mfn] a  -    n   -   -    -  -   12.
V     a    p   [a]    -    s     m    a  -   [sf] -   -    -  -   13.
V     a    p   [a]    -    s    [fn]  a  -    y   -   -    -  -   14.
V     a    p   [ai]   -    p     -    a  -    n   -   -    -  -   15.
V     a    p   [a]    -    p     -    a  -    y   -   -    -  -   16.
V    [ma]  g    -     -    -     -    -  -    -   -   -    -  -   17.
=== ===== ==== ==== ==== ===== ====  === === === === ==== == === =======

Examples:
 1. govorya, sym
 2. govori, bydete

 3. govorest, govorila
 4. govorestiya, govorestiyat
 5. govorestta, govoriloto
 6. govoresti, govoreli
 7. govoresti, govorilite
 8. govoren, govorena
 9. govoreniya, govoreniyat
10. govorenata, govorenoto
11. govoreni, govorenite
12. bil, bila
13. biliya, biliyat
14. bilata, biloto
15. bili
16. bilite
17. govorejki, bidejki

3. Adjective (A)
= ============== ============== =============== ==
P ATT            VAL             Example        CC
= ============== ============== =============== ==
1 Type                             -             -
- -------------- -------------- --------------- --
2 Degree                           -             -
- -------------- -------------- --------------- --
3 Gender         masculine       bedniyat        m
                 feminine        bedna           f
                 neuter          bednoto         n
- -------------- -------------- --------------- --
4 Number         singular        bedniya         s
                 plural          bednite         p
- -------------- -------------- --------------- --
5 Case                                           -
* ************************************************
6 Definiteness   no              bedna           n
                 yes             bednata         y
           l.s.  short_art       bedniya         s
           l.s.  full_art        bedniyat        f
- -------------- -------------- --------------- --
7 Clitic                                         -
- -------------- -------------- --------------- --
8 Animate                                        -
- -------------- -------------- --------------- --
9 Formation                                      -
- -------------- -------------- --------------- --
10Owner_Number                                   -
- -------------- -------------- --------------- --
11Owner_Person                                   -
---------------- -------------- --------------- --
12Owned_Number                                   -
==================================================

Combinations
=== ==== ==== ==== ==== ==== === ==== === === === === === =======
PoS Type Degr Gend Numb Case Def Clit  An Frm OnN OnP OdN Example
=== ==== ==== ==== ==== ==== === ==== === === === === === =======
A    -    -    s  [mfn]  -    n   -    -   -   -   -   -  beden, bedna
A    -    -    s    m    -   [sf] -    -   -   -   -   -  bedniya, bedniyat
A    -    -    s  [fn]   -    y   -    -   -   -   -   -  bednata, bednoto
A    -    -    p   -     -   [ny] -    -   -   -   -   -  bedni, bednite
=== ==== ==== ==== ==== ==== === ==== === === === === === =======

4. Pronoun (P)
= ============== ============== =========== =
P ATT            VAL            Example     C
= ============== ============== =========== =
1 Type           personal        az         p
                 demonstrative   tozi       d
                 indefinite      nyakoj     i
                 possessive      moj        s
                 interrogative   koi        q
                 relative        kojto      r
                 reflexive       sebe       x
                 negative        nikoj      z
                 general         vseki      g
- -------------- -------------- ----------- -
2 Person         first           az         1
                 second          ti         2
                 third           toj        3
- -------------- -------------- ----------- -
3 Gender         masculine       moj        m
                 feminine        moya       f
                 neuter          moe        n
- -------------- -------------- ----------- -
4 Number         singular        az         s
                 plural          nie        p
- -------------- -------------- ----------- -
5 Case           nominative      toj        n
                 dative          nemu       d
                 accusative      nego       a
- -------------- -------------- ----------- -
6 Owner_Number                              -
- -------------- -------------- ----------- -
7 Owner_Gender                              -
*********************************************
8 Clitic         no              sebe_si    n
                 yes             mi         y
- -------------- -------------- ----------- -
9 Referent_Type  personal        koj        p
                 possessive      tchij      s
                 attributive     kakva      a

                 quantitative    kolko      q
- -------------- -------------- ----------- -
10Syntactic_Type                            -
- -------------- -------------- ----------- -
11Definiteness   no              svoj       n
                 yes             svoyata    y
                 short_art       svoya      s
                 full_art        svoyat     f
- -------------- -------------- ----------- -
12Animate                                   -
- -------------- -------------- ----------- -
13Clitic_s                                  -
- -------------- -------------- ----------- -
14Pronoun_Form                              -
- -------------- -------------- ----------- -
15Owner_Person                              -
- -------------- -------------- ----------- -
16Owned_Number                              -
=============================================

Combinations
Personal pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    p  [12]   -    s   n    -   -   -   -    -     -   1.
P    p   3    [mfn] s   n    -   -   -   -    -     -   2.
P    p  [123]  -    p   n    -   -   -   -    -     -   3.
P    p  [12]   -    s  [ad]  -   -  [yn] -    -     -   4.
P    p   3    [mnf] s  [ad]  -   -  [yn] -    -     -   5.
P    p  [123]  -    p  [ad]  -   -  [yn] -    -     -   6.
Examples:
1. az, ti
2. toj, tya, to
3. nie, vie, te
4. mene, me, mi, mene_me
5. nego, go, nego_go
6. nas, ni, nas_ni, nam

Demonstrative pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    d    -  [mfn]  s   -    -   -   -   p    -    -   tozi,tova
P    d    -   -     p   -    -   -   -   p    -    -   tezi
P    d    -   -     -   -    -   -   -   q    -    -   tolkova

Indefinite, Interrogative, Relative, Negative pronouns

=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P [iqrz]  -    m    s  [nad] -   -   -   p    -   -     1.
P [iqrz]  -    m    s   -    -   -   -  [as]  -   -     2.
P [iqrz]  -   [fn]  s   -    -   -   -  [pas] -   -     3.
P [iqrz]  -    -    p   -    -   -   -  [pas] -   -     4.
P [iqrz]  -    -    -   -    -   -   -   q    -   -     5.
Examples:
1. nyakoj, nyakogo, nyakomu, nikoj
2. nyakakyv, netchij, nikakyv
3. nyakoya, nyakakva, nechiya
4. nyakoi, nyakavi, netchii
5. nyakolko, nikolko, kolko

Possessive pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    s   [123]  m   s   -   -    -   n   -    -  [nsf] 1.
P    s   [123] [fn] s   -   -    -   n   -    -  [ny]  2.
P    s   [123]  -   p   -   -    -   n   -    -  [ny]  3.
P    s   [123]  -  [sp] -   -    -   y   -    -   -    4.
Examples:
1. moj, moya, moyat
2. moya, moyata, tvoya
3. moi, moite, tvoi
4. mi, ti, mu, j, ni, vi, im

Reflexive pronouns
=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    x    -    -    -  [ad]  -   -  [ny] p    -    -    1.
P    x    -    m    s   -    -   -    n  s    -  [nsf]  2.
P    x    -   [fn]  s   -    -   -    n  s    -  [ny]   3.
P    x    -    -    p   -    -   -    n  s    -  [ny]   4.
P    x    -    -    -   -    -   -    y  s    -    -    5.
Examples:
1. sebe, se, si, sebe_si
2. svoj, svoyat, svoya
3. svoya, svoyata
4. svoi, svoite
5. si

General pronouns

=== ==== ==== ==== === ==== === === === ==== ==== === =======
PoS Type Pers Gnd  Nmb Case OnN OnG Clt RefT SynT Def Example
=== ==== ==== ==== === ==== === === === ==== ==== === =======
P    g    -    m    s  [nad] -   -   -   p    -    -    1.
P    g    -    m    s    -   -   -   -   a    -    -    2.
P    g    -   [fn]  s    -   -   -   -  [pa]  -    -    3.
P    g    -    -    p    -   -   -   -  [pa]  -    -    4.
P    g    -    m    s    -   -   -   -   q    -   [sf]  5.
P    g    -   [fn]  s    -   -   -   -   q    -   [ny]  6.
P    g    -    -    p    -   -   -   -   q    -   [ny]  7.
Examples:
1. vseki, vsekigo, vsekimu
2. vsyakakyv
3. vsyaka, vsyakakva
4. vsyakoi, vsyakakvi
5. vsitchkiya, vsitchikiyat
6. vsitchka, vsitchkata
7. vsitchki, vsitchkite

5. Determiner (D)
Not applicable.

6. Article (T)
Not applicable.

7. Adverb (R)
= ============== ============== ============ =
P ATT            VAL             Example     C
= ============== ============== ============ =
1 Type           general          tuk        g
           l.s.  adjectival       umno       a
- -------------- --------------- ----------- -
2 Degree                                     -
**********************************************
3 Clitic                                     -
- -------------- -------------- ------------ -
4 Number                                     -
- -------------- -------------- ------------ -
5 Person                                     -
==============================================
Combinations
===== ====== ==== ==== ==== ===== ======================
POS    Type  Degr Clct Numb Pers  Example
===== ====== ==== ==== ==== ===== ======================

R        g    -    -    -    -     tuk, mnogo
R        a    -    -    -    -     umno, veselo, studeno
========================================================
Note: Adverbs of type adjectival have the same form as adjectives in
Gender = neuter, Person = 3, Number = singular.

8. Adposition (S)
= ============== ============== ============= ==
P ATT            VAL             Example       C
= ============== ============== ============= ==
1 Type           preposition     na, v         p
- -------------- -------------- ------------- --
2 Formation                                    -
************************************************
3 Case                                         -
- -------------- -------------- ------------- --
4 Clitic                                       -
================================================
Combinations
=== ===== ==== ==== ==== ============================
POS Type  Form Case Cltc  Examples
=== ===== ==== ==== ==== ============================
S    p     -    -    -    na, za, v, okolo, spored
=== ===== ===========================================

9. Conjunction (C)
= ============== ============== ============== ==
P ATT            VAL            Example         C
= ============== ============== ============== ==
1 Type           coordinating   a, ala, ili     c
                 subordinating  tche, da        s
*************************************************
2 Formation      simple         a, i, tche      s
                 compound       i_da, za_da     c
- -------------- -------------- -------------- --
3 Coord_Type                                    -
- -------------- -------------- -------------- --
4 Sub_Type                                      -
- -------------- -------------- -------------- --
5 Clitic                                        -
- -------------- -------------- -------------- --
6 Number                                        -
- -------------- -------------- -------------- --
7 Person                                        -
=================================================

Combinations
=== ==== ==== ==== ==== ==== ==== ===== ==========================
PoS Type Form CTyp SubT Clit Numb Pers  Examples
=== ==== ==== ==== ==== ==== ==== ===== ==========================
C    c   [sc]  -    -     -   -    -    i, a, ami, tche, tyj_tche
C    s   [sc]  -    -     -   -    -    tche, da, kato, za_da
=== ==== ==== ==== ==== ==== ==== ===== ==========================

10. Numeral (M)
= ============== ============== ============ =
P ATT            VAL             Example     C
= ============== ============== ============ =
1 Type           cardinal         edin       c
                 ordinal          vtori      o
- -------------- -------------- ------------ -
2 Gender         masculine        edin       m
                 feminine         edna       f
                 neuter           edno       n
- -------------- -------------- ------------ -
3 Number         singular         edin       s
                 plural           edni       p
- -------------- -------------- ------------ -
4 Case                            -          -
**********************************************
5 Form           digit            1984       d
                 roman            IX         r
                 letter           dva        l
           l.s.  m-form           dvama      m
           l.s.  approx           stotina    a
- -------------- -------------- ------------ -
6 Definiteness   no               edin       n
                 yes              ednata     y
                 short_art        ediniya    s
                 full_art         ediniyat   f
- -------------- -------------- ------------ -
7 Clitic                                     -
- -------------- -------------- ------------ -
8 Class                                      -
- -------------- -------------- ------------ -
9 Animate                                    -
- -------------- -------------- ------------ -
10Owner_Number                               -
- -------------- -------------- ------------ -
11Owner_Person                               -
- -------------- -------------- ------------ -
12Owned_Number                               -
===============================================

Combinations
=== ===== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ===========
PoS Type  Gend Numb Case Frm  Def  Clit Clas Anim OwnN OwnP OwdN   Example
=== ===== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ===========
Cardinal
M     c    -    -    -    d    -    -    -    -    -    -    -       1.
M     c    -    -    -    r    -    -    -    -    -    -    -       2.
M     c    m    s    -    l   [nsf] -    -    -    -    -    -       3.
M     c   [fn]  s    -    l   [ny]  -    -    -    -    -    -       4.
M     c    -    p    -    l   [ny]  -    -    -    -    -    -       5.
M     c   [mfn] p    -    l   [ny]  -    -    -    -    -    -       6.
M     c    m    p    -    m   [ny]  -    -    -    -    -    -       7.
M     c    -    p    -    l   [ny]  -    -    -    -    -    -       8.
M     c    -    p    -    a   [ny]  -    -    -    -    -    -       9.
Ordinal
M     o    m    s    -    -   [nsf] -    -    -    -    -    -      10.
M     o   [fn]  s    -    -   [ny]  -    -    -    -    -    -      11.
M     o    -    p    -    -   [ny]  -    -    -    -    -    -      12.
=== ===== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ==== ===========
Examples:
1. 1984
2. IX
3. edin, ediniya, ediniyat
4. edna, ednata, edno, ednoto
5. edni, ednite
6. dva, dvata, dve, dvete
7. dvama, dvamata, trima
8. tri, trite, chetiri
9. stotina, stotinata
10. pyrvi, pyrviya, pyrviyat
11. pyrva, pyrvata
12. pyrvi, pyrvite

11. Interjection (I)
= ============== ============== ============ ==
P ATT            VAL             Example      C
= ============== ============== ============ ==
1 Type                                        -
- -------------- -------------- ------------ --
2 Formation      simple          ah           s
                 compound        bozhe moj!   c
============================================ ==

Combinations

=== ==== ========== =================================================
POS Type Formation  Examples
=== ==== ========== =================================================
I    -    s         o, hm-m, ha-ha, ah-lele, uvi
I    -    c         diavol_da_go_vzeme, ama_rabota, bozhe_moj
=====================================================================

12. Residual (X)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
P ATT                           -
=================================

13. Abbreviation (Y)
= ============== ============== ============ =
P ATT            VAL             Example     C
= ============== ============== ============ =
1 Syntactic_Type                             -
- -------------- -------------- ------------ -
2 Gender
- -------------- -------------- ------------ -
3 Number
- -------------- -------------- ------------ -
4 Case
- -------------- -------------- ------------ -
5 Definiteness
==============================================
Note: As there are no Attributes for Abbreviations, this subsection
      is has no examples
Combinations
==== ==== ==== ==== === ===========================
PoS  SynT Gend Numb Case Def  Examples
==== ==== ==== ==== === ===========================
Y     -    -   -     -   -    t.n., vzh., lat., razg...
===================================================

14. Particle (Q)
= ============== ============== =========== ==
P ATT            VAL             Example     C
= ============== ============== =========== ==
1 Type           negative        ne, ni      z
                 general         a, be       g
                 comparative     po, naj     c
                 verbal          da, ste     v

                 interrogative   li, dali    q
                 modal           da, dano    o
- -------------- -------------- ----------- --
2 Formation      simple          a, ne       s
                 compound        hajde_de    c
- -------------- -------------- ----------- --
3 Clitic                                     -
==============================================

Combinations
=== ============ ======= =======================
POS  Type         Form   Cltc Examples
=== ============ ======= =======================
Q    [go]         [sc]    -    a, be, neka,hajde_de
Q    [cvzq]        s      -    po, naj, ne, li, nali
=== ============ ======= =======================
Observations
In Bulgarian superlative particles  "po" and "naj" precede
the superlative Forms of adjectives and adverbs:
"po-hubav", "naj-hubav", "po-bavno", "naj-bavno".

COP project 106 multext-East  Deliverable D1.1 F— Czech  48

3.2 Application to Czech

The application to Czech has been elaborated by Vladimír Petkevič, Faculty of Philosophy, Charles University, Prague.

Acknowledgements: The most appreciated base used for the elaboration of the application was constituted by morphosyntactic tables for Czech elaborated by Hana Skoumalová. For all possible errors solely the author is to blame.

All Czech diacritical characters used have been encoded in the following way:

1. The Czech ’hachek’ diacritic (ˇ ) over the following characters is marked by the corresponding nondiacritic counterpart followed by ’<’. This concerns the characters c, d, e, n, r, s, t, z followed by <. For instance: t<

2. In Czech, each vowel (a, e, i, o, u, y) can be diacritically marked as to its quantity by ’ over the vowel. We shall mark it by ' following the vowel, e.g. e'. In addition to ' there exists also the ring sign (^{o}) over the letter u. Here we shall use this notation: u0. Thus, there are the following vocal diacritical alternatives:

     a' e' i' o' u' u0 y'
1. Noun (N)
1.1. Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Czech term
= ============== ============== = ================= =========================
1 Type           common         c kniha             obecne' jme'no
                 proper         p Petr              vlastni' jme'no
- -------------- -------------- - ----------------- -------------------------
2 Gender         masculine      m otec              masculinum
                 feminine       f kniha             femininum
                 neuter         n slunce            neutrum
- -------------- -------------- - ----------------- -------------------------
3 Number         singular       s kniha             singula'r
                 plural         p knihy             plura'l
                 dual           d rukama            dua'l
- -------------- -------------- - ----------------- -------------------------
4 Case           nominative     n kniha             nominativ
                 genitive       g knihy             genitiv
                 dative         d knize             dativ
                 accusative     a knihu             akuzativ
                 vocative       v kniho!            vokativ
                 locative       l knize             loka'l
                 instrumental   i knihou            instrumenta'l
* ************** ************** * ----------------- -------------------------
5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -

- -------------- -------------- - ----------------- -------------------------
7 Animate   l.s. no             n hrad              nez<ivotne'
           l.s. yes            y otec              z<ivotne'
- -------------- -------------- - ----------------- -------------------------
8 Owner_Number                  -
- -------------- -------------- -
9 Owner_Person                  -
- -------------- -------------- -
10Owned_Number                  -
=================================
Notes:
  1. Czech  distinguishes masculine animate  (Animate=yes above)
  and masculine inanimate  (Animate=no) Gender. Masculine inani-
  mate Nouns always have the same form in the nominative and ac-
  cusative Case,  whereas masculine animate  Nouns have predomi-
  nantly the same form in the genitive and accusative Case. Mas-
  culine animate  Nouns and masculine inanimate  Nouns differ in
  accusative singular and nominative plural only.
2. The dual Number manifests itself only in the instrumental Case
  of several Nouns denoting dual parts of the human body.
3. Czech distinguishes 7 cases, the locative Case being obligato-
  rily prepositional. With most paradigms, the form of the voca-
  tive differs from that of  the nominative in singular; in plu-
  ral the form of the vocative is identical to that of the nomi-
  native for all the paradigms.
4. Verbal nouns are classified as Nouns.
1.2 Combinations
*** **** **** **** **** **** =============================================
PoS Type Gend Numb Case Anim Examples
*** **** **** **** **** **** =============================================
N   p    m   [sp] any   y   Pavel, Pavlove'
N   p    m   [sp] any   n   Hradec, Hradce
N   p    f   [sp] any   -   Lenka, Lenky
N   p    n   [sp] any   -   Bradlo, U'sti'
N   c    m    s    n    -   kluk,      vor
N   c    m    s    g    -   kluka,     voru
N   c    m    s    d    -   klukovi,   voru
N   c    m    s    a    y   kluka
N   c    m    s    a    n              vor
N   c    m    s    v    -   kluku,     vore!
N   c    m    s    l    -   klukovi,   voru
N   c    m    s    i    -   klukem,    vorem
N   c    m    p    n    y   kluci
N   c    m    p    n    n              vory
N   c    m    p    g    -   kluku0     voru0
N   c    m    p    d    -   kluku0m    voru0m
N   c    m    p    a    -   kluky      vory
N   c    m    p    v    -   kluci!     vory!
N   c    m    p    l    -   kluci'ch   vorech

 N   c    m    p    i    -   kluky      vory
 N   c   [fn]  s    n    -   sova,      kolo
 N   c   [fn]  s    g    -   sovy,      kola
 N   c   [fn]  s    d    -   sove<      kolu
 N   c   [fn]  s    a    -   sovu       kolo
 N   c   [fn]  s    v    -   sovo!      kolo!
 N   c   [fn]  s    l    -   sove<      kole / kolu
 N   c   [fn]  s    i    -   sovou      kolem
 N   c   [fn]  p    n    -   sovy       kola
 N   c   [fn]  p    g    -   sov        kol
 N   c   [fn]  p    d    -   sova'm     kolu0m
 N   c   [fn]  p    a    -   sovy       kola
 N   c   [fn]  p    v    -   sovy!      kola!
 N   c   [fn]  p    l    -   sova'ch    kolech
 N   c   [fn]  p    i    -   sovami     koly
 N   c   [fn]  d    i    -   rukama     oc<ima    (dual number)
*** **** **** **** **** **** =============================================
Note: In the Combinations above, 'any' is a variable standing for
      any admissible value.
2. Verb (V)
2.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C  Example            Czech term
= ============== ============== = =================== ========================
1 Type           main           m de<lat              vy'znamove' sloveso
                 auxiliary      a mi't                pomocne' sloveso
                 modal          o muset               moda'lni' sloveso
                 copula         c by't                spona (kopula)
- -------------- -------------- - ------------------- ------------------------
2 VForm          indicative     i de<la'm             indikativ
                 imperative     m de<lej!             imperativ
                 conditional    c by                  podmin<ovaci' c<a'stice
                 infinitive     n de<lat              infinitiv
                 participle     p de<lal, de<la'n     pr<i'c<esti'
            l.s. transgressive  t de<laje, ude<lavs<i pr<echodni'k
- -------------- -------------- - ------------------- ------------------------
3 Tense          present        p de<la'm             pre'zens
                 future         f budu, pobe<z<i'm    futurum
                 past           s de<lal              c<as minuly'
- -------------- -------------- - ------------------- ------------------------
4 Person         first          1 de<la'm             prvni' osoba
                 second         2 de<la's<            druha' osoba
                 third          3 de<la'              tr<eti' osoba
- -------------- -------------- - ------------------- ------------------------
5 Number         singular       s de<la'm             singula'r
                 plural         p de<la'me            plura'l
- -------------- -------------- - ------------------- ------------------------
6 Gender         masculine      m de<lal              masculinum /rod muz<sky'
                 feminine       f de<lala             femininum /rod z<ensky'

                 neuter         n de<lalo             neutrum /rod str<edni'
* ************** ************** * ------------------- ------------------------
7 Voice          active         a de<lal              aktivum
                 passive        p de<la'n             pasi'vum
- -------------- -------------- - ------------------- ------------------------
8 Negative       no             n de<la'm             kladne' sloveso
                 yes            y nede<la'm           za'porne' sloveso
- -------------- -------------- - ------------------- ------------------------
9 Definiteness                  -
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- - ------------------- ------------------------
12Animate   l.s. no             n de<laly             nez<ivotne'
            l.s. yes            y de<lali             z<ivotne'
- -------------- -------------- - ------------------- ------------------------
13Clitic_s  l.s. no             n ude<lal             tvar bez klit. -s
            l.s. yes            y ude<lals            tvar s klit. -s
==============================================================================
Notes
1.  The verb "by't" (E. "to be") in all its functions is characte-
    rized as Type=c  (i.e. the copula), which clearly  is an over-
    simplification because  the verb has  more meanings (auxiliary
    etc.).
2.  Auxiliary verbs (Type=a) include  neither the verb "by't" (see
    above), nor the modal verbs.
3.  The 'past participle' in Czech is used for expressing compound
    active past Tense and is encoded as:
    Type=p(articiple), Tense=p(ast), Voice=a(ctive).
4.  Passive participles are encoded as:
    Type=participle, Tense='-', Voice=Passive.
    This encoding is  used only for the passive  participle in the
    predicative  position,   e.g.  "byl  chycen"   (E.  "(he)  was
    caught").  The case  marked passive  participles, both  in the
    predicative and in the  attributive position, e.g. "pade<lany'
    rukopis" (E. "forged manuscript"),  are be classified as (qua-
    lificative) adjectives.
5.  Adjectival  active and  passive participles,  e.g. "stoji'ci'"
    (E. "standing") or "ude<lany'"  (E. "performed" or "done", cf.
    Note 4 above) are classified as adjectives.
6.  Negative verbs are marked  as Negative=y, whereas non-negative
    verbs are marked as Negative=n.
7.  The term transgressive roughly corresponds to the term 'verbal
    participle'. The transgressives distinguish:
     Number (singular, plural),
     Gender (masculine, non-masculine)
     Tense (past, present)
    The Gender opposition of  masculine animate vs. masculine ina-
    nimate is  neutralized in transgressives;  thus, there is  one
    masculine form and one non-masculine (the form of feminine and

   neuter is neutralized).
8.  Gender and Animate values  correspond to those associated with
   Nouns and are necessary to account properly for agreement.
9.  Gender  manifests itself  in past  participles, passive parti-
   ciples and transgressives only.
10. Normally,  Verbs  form  the  future  Tense periphrastically by
   auxiliary  "by't" (E.  "to be")  plus infinitive  of the  main
   Verb.  In addition  to the  copula, there  are, however,  some
   Verbs  which  form  future  Tense  non-periphrastically,  i.e.
   synthetically (Verbs of motion).  Such verbal forms are marked
   as Tense=f.
11. The  'yes' value  of the  Clitic_s attribute  denotes a verbal
   form having the clitic morpheme 's' appended as a suffix. This
   's' morpheme  expresses 2nd Person  singular present Tense  of
   the auxiliary Verb  "by't" (i.e. the form "jsi").  There is no
   intermediate  hyphen  between  the  verbal  form  and  the 's'
   morpheme.
12. The  Clitic_s  attribute  is  specified  for  VForm=infinitive
   (VForm=n) and Vform=p(articiple) only.
13. Some  modal and  auxiliary Verbs  do not  form imperative  and
   transgressive.
14. The   voice  value   ('a'  or   'p')  is   not  specified  for
   VForm=c(onditional)  and VForm=t(ransgressive),  in which case
   VForm='-'.

2.2 Combinations
*** **** **** **** **** **** ***** ---- ---- ---- ---- ========================
PoS Type VFrm Tens Pers Numb Gend  Voic Neg  Anim Cl_s Examples
*** **** **** **** **** **** ***** ---- ---- ---- ---- ========================
V   m     n    -    -    -    -   [na] [ny]   -  [ny] pra't, neprat
V   a     n    -    -    -    -   [na] [ny]   -  [ny] mi't, nemi't
V   o     n    -    -    -    -   [na] [ny]   -  [ny] muset, nemuset
V   c     n    -    -    -    -   [na] [ny]   -  [ny] by't, neby't
V   c     c    -    3    s    -     -    -    -    -  by
V   c     c    -  [12]   s    -     -    -    -    -  bych, bys
V   c     c    -  [12]   p    -     -    -    -    -  bychom, byste
V   c     i    f  [123]  s    -     a    n    -    -  budu, budes<, bude
V   c     i    f  [123]  s    -     a    y    -    -  nebudu, nebudes<, nebude
V   c     i    f  [123]  p    -     a    n    -    -  budeme, budete, budou
V   c     i    f  [123]  p    -     a    y    -    -  nebudeme, nebudete, nebudou
V   m     i    f  [123]  s    -     a    n    -    -  poleti'm, poleti's<, poleti'
V   m     i    f  [123]  s    -     a    y    -    -  nepoleti'm, nepoleti's<,
                                                      nepoleti'
V   m     i    f  [123]  p    -     a    n    -    -  poleti'me, poleti'te, poleti'
V   m     i    f  [123]  p    -     a    y    -    -  nepoleti'me, nepoleti'te,
                                                      nepoleti'
V   c     i    p  [123]  s    -     a    n    -    -  jsem, jsi, je
V   c     i    p  [123]  s    -     a    y    -    -  nejsem, nejsi, neni'
V   c     i    p  [123]  p    -     a    n    -    -  jsme, jste, jsou
V   c     i    p  [123]  p    -     a    y    -    -  nejsme, nejste, nejsou
V   m     i    p  [123]  s    -     a    n    -    -  tr<u, tr<es<, tr<e

V   m     i    p  [123]  s    -     a    y    -    -  netr<u, netr<es<, netr<e
V   m     i    p  [123]  p    -     a    n    -    -  tr<eme, tr<ete, tr<ou
V   m     i    p  [123]  p    -     a    y    -    -  netr<eme, netr<ete, netr<ou
V   a     i    p  [123]  s    -     a    n    -    -  ma'm, ma's<, ma'
V   a     i    p  [123]  s    -     a    y    -    -  nema'm, nema's<, nema'
V   a     i    p  [123]  p    -     a    n    -    -  ma'me, ma'te, maji'
V   a     i    p  [123]  p    -     a    y    -    -  nema'me, nema'te, nemaji'
V   o     i    p  [123]  s    -     a    n    -    -  musi'm, musi's<, musi'
V   o     i    p  [123]  s    -     a    y    -    -  nemusi'm, nemusi's<,
                                                      nemusi'
V   o     i    p  [123]  p    -     a    n    -    -  musi'me, musi'te,
                                                      museji'/musi'
V   o     i    p  [123]  p    -     a    y    -    -  nemusi'me, nemusi'te,
                                                      nemuseji'
V   c     m    p    1    p    -     a   [ny]  -    -  bud<me!, nebud<me!
V   c     m    p    2   [sp]  -     a    n    -    -  bud<!, bud<te!
V   c     m    p    2   [sp]  -     a    y    -    -  nebud<!, nebud<te!
V   m     m    p    1    p    -     a   [ny]  -    -  pracujme!, nepracujme!
V   m     m    p    2   [sp]  -     a    n    -    -  pracuj!, pracujte!
V   m     m    p    2   [sp]  -     a    y    -    -  nepracuj!, nepracujte!
V   c     p    s    -    s  [mfn]   a    n    -    n  byl, byla, bylo
V   c     p    s    -    s  [mfn]   a    y    -    n  nebyl, nebyla, nebylo
V   c     p    s    -    s  [mfn]   a    n    -    y  byls, bylas, bylos
V   c     p    s    -    s  [mfn]   a    y    -    y  nebyls, nebylas,
                                                      nebylos
V   c     p    s    -    p   m      a    n  [yn]   n  byli, byly
V   c     p    s    -    p   m      a    y  [yn]   n  nebyli, nebyly
V   c     p    s    -    p  [fn]    a    n    -    n  byly, byla
V   c     p    s    -    p  [fn]    a    y    -    n  nebyly, nebyla
V   a     p    s    -    s  [mfn]   a    n    -    n  me<l, me<la, me<lo
V   a     p    s    -    s  [mfn]   a    y    -    n  neme<l, neme<la,
                                                      neme<lo
V   a     p    s    -    s  [mfn]   a    n    -    y  me<ls, me<las, me<los
V   a     p    s    -    s  [mfn]   a    y    -    y  neme<ls, neme<las,
                                                      neme<los
V   a     p    s    -    p   m      a    n  [yn]   n  me<li, me<ly
V   a     p    s    -    p   m      a    y  [yn]   n  neme<li, neme<ly
V   a     p    s    -    p  [fn]    a    n    -    n  me<ly, me<la
V   a     p    s    -    p  [fn]    a    y    -    n  neme<ly, neme<la
V   o     p    s    -    s  [mfn]   a    n    -    n  musel, musela, muselo
V   o     p    s    -    s  [mfn]   a    y    -    n  nemusel, nemusela,
                                                      nemuselo
V   o     p    s    -    s  [mfn]   a    n    -    y  musels, muselas,
                                                      muselos
V   o     p    s    -    s  [mfn]   a    y    -    y  nemusels, nemuselas,
                                                      nemuselos
V   o     p    s    -    p   m      a    n  [yn]   n  museli, musely
V   o     p    s    -    p   m      a    y  [yn]   n  nemuseli, nemusely
V   o     p    s    -    p  [fn]    a    n    -    n  musely, musela
V   o     p    s    -    p  [fn]    a    y    -    n  nemusely, nemusela
V   m     p    s    -    s  [mfn]   a    n    -    n  de<lal, de<lala,
                                                      de<lalo
V   m     p    s    -    s  [mfn]   a    y    -    n  nede<lal, nede<lala,

                                                      nede<lalo
V   m     p    s    -    s  [mfn]   a    n    -    y  de<lals, de<lalas,
                                                      de<lalos
V   m     p    s    -    s  [mfn]   a    y    -    y  nede<lals, nede<lalas,
                                                      nede<lalos
V   m     p    s    -    p   m      a    n  [yn]   n  de<lali, de<laly
V   m     p    s    -    p   m      a    y  [yn]   n  nede<lali, nede<laly
V   m     p    s    -    p  [fn]    a    n    -    n  de<laly, de<lala
V   m     p    s    -    p  [fn]    a    y    -    n  nede<laly, nede<lala
V   m     p    -    -    s  [mfn]   p    n    -    n  ude<la'n, ude<la'na
                                                      ude<la'no
V   m     p    -    -    s  [mfn]   p    y    -    n  neude<la'n, neude<la'na
                                                      neude<la'no
V   m     p    -    -    p   m      p    n  [yn]   n  ude<la'ni, ude<la'ny
V   m     p    -    -    p   m      p    y  [yn]   n  neude<la'ni, neude<la'ny
V   m     p    -    -    p  [fn]    p    n    -    n  ude<la'ny, ude<la'na
V   m     p    -    -    p  [fn]    p    y    -    n  neude<la'ny, neude<la'na
V   c     t    p    -    s   m      -  [ny]   -    -  jsa, nejsa
V   c     t    p    -    s  [fn]    -  [ny]   -    -  jsouc, nejsouc
V   c     t    p    -    p   -      -  [ny]   -    -  jsouce, nejsouce
V   m     t    p    -    s   m      -  [ny]   -    -  de<laje, nede<laje
V   m     t    p    -    s  [fn]    -  [ny]   -    -  de<laji'c, nede<laji'c
V   m     t    p    -    p   -      -  [ny]   -    -  de<laji'ce nede<laji'ce
V   c     t    s    -    s   m      -  [ny]   -    -  byv, nebyv
V   c     t    s    -    s  [fn]    -  [ny]   -    -  byvs<i, nebyvs<i
V   c     t    s    -    p   -      -  [ny]   -    -  byvs<e, nebyvs<e
V   m     t    s    -    s   m      -  [ny]   -    -  ude<lav, neude<lav
V   m     t    s    -    s  [fn]    -  [ny]   -    -  ude<lavs<i, neude<lavs<i'
V   m     t    s    -    p   -      -  [ny]   -    -  ude<lavs<e, neude<lavs<e
*** **** **** **** **** **** ***** ---- ---- ---- ---- =======================

3. Adjective (A)
3.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           qualificative  f dobry'              kvalitativni' adjektivum
                possessive     s matc<in             posesi'vni' adjektivum
- -------------- -------------- --------------------- ------------------------
2 Degree         positive       p dobry'              pozitiv / 1. stupen<
                comparative    c leps<i'             komparativ / 2. stupen<
                superlative    s nejleps<i'          superlativ / 3. stupen<
- -------------- -------------- -  ------------------ ------------------------
3 Gender         masculine      m dobry'              masculinum / rod muz<sky'
                feminine       f dobra'              femininum / rod z<ensky'
                neuter         n dobre'              neutrum  / rod str<edni'
- -------------- -------------- --------------------- ------------------------
4 Number         singular       dobry'                singula'r
                plural         dobr<i'               plura'l
                dual           dobry'ma              dua'l

- -------------- -------------- --------------------- ------------------------
5 Case           nominative     n dobry'              nominativ
                 genitive       g dobre'ho            genitiv
                 dative         d dobre'mu            dativ
                 accusative     a dobre'ho            akuzativ
                 vocative       v dobry'!             vokativ
                 locative       l dobre'm             loka'l
                 instrumental   i dobry'm             instrumenta'l
* ***************************** * ------------------- ------------------------
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- - ------------------- ------------------------
8 Animate        no             n dobre'              nez<ivotne'
                 yes            y dobr<i'             z<ivotne'
- -------------- -------------- - ------------------- ------------------------
9 Formation l.s. nominal        n sta'r               jmenny' tvar
            l.s. compound       c stary'              sloz<eny' tvar
- -------------- -------------- --------------------- ------------------------
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================
Notes
1. Three  deverbative  adjectival  participles,  i.e. past active
   participle, passive  participle and present  active participle
   are not  distinguished. They are conflated  in the 'qualifica-
   tive' value of the Type attribute (Type=f).
2. Only qualificative (and passive  participle) Adjectives can be
   specified for Degree and for  nominal (short) Form. 3. The at-
   tributes Gender,  Number, Case and  Animate correspond to  the
   same categories  within the Nouns. They  are necessary for the
   proper account of agreement of Adjectives with Nouns.
4. The Formation  attribute distinguishes a  nominal (short) form
   from a so-called compound (long) form of an Adjective. The no-
   minal form can be used in the predicative function only. It is
   specified for nominative and accusative Case only.
5. The Adjectives  which have no  degrees of comparison  have the
   Degree value equal to p(ositive).
3.2 Combinations
*** **** ***** **** **** **** **** **** ======================================
PoS Type Degr  Gend Numb Case Anim Form Examples
*** **** ***** **** **** **** **** **** ======================================
 A   s     -    m     s    n    -   -   otcu0v
 A   s     -    m     s    g    -   -   otcova
 A   s     -    m     s    d    -   -   otcovu
 A   s     -    m     s    a    y   -   otcova

A   s     -    m     s    a    n   -            otcu0v
A   s     -    m     s    v    -   -   otcu0v!
A   s     -    m     s    l    -   -   otcove< / otcovu
A   s     -    m     s    i    -   -   otcovy'm
A   s     -    m     p    n    y   -   otcovi
A   s     -    m     p    n    n   -            otcovy
A   s     -    m     p    g    -   -   otcovy'ch
A   s     -    m     p    d    -   -   otcovy'm
A   s     -    m     p    a    -   -   otcovy
A   s     -    m     p    v    y   -   otcovi!
A   s     -    m     p    v    n   -             otcovy!
A   s     -    m     p    l    -   -   otcovy'ch
A   s     -    m     p    i    -   -   otcovy'mi
A   s     -   [fn]   s    n    -   -   otcova,  otcovo
A   s     -   [fn]   s    g    -   -   otcovy,  otcova
A   s     -   [fn]   s    d    -   -   otcove<, otcovu
A   s     -   [fn]   s    a    -   -   otcovu,  otcovo
A   s     -   [fn]   s    v    -   -   otcova!, otcovo
A   s     -   [fn]   s    l    -   -   otcove<,  otcove< / otcovu
A   s     -   [fn]   s    i    -   -   otcovou, otcovy'm
A   s     -   [fn]   p    n    -   -   otcovy,  otcova
A   s     -   [fn]   p    g    -   -   otcovy'ch
A   s     -   [fn]   p    d    -   -   otcovy'm
A   s     -   [fn]   p    a    -   -   otcovy,  otcova
A   s     -   [fn]   p    v    -   -   otcovy!  otcova!
A   s     -   [fn]   p    l    -   -   otcovy'ch
A   s     -   [fn]   p    i    -   -   otcovy'mi
A   s     -   [fn]   d    i    -   -   otcovy'ma
A   f   [pcs]  m     s    n    -   c   dobry'
A   f   [pcs]  m     s    g    -   c   dobre'ho
A   f   [pcs]  m     s    d    -   c   dobre'mu
A   f   [pcs]  m     s    a    y   c   dobre'ho
A   f   [pcs]  m     s    a    n   c   dobry'
A   f   [pcs]  m     s    v    -   c   dobry'!
A   f   [pcs]  m     s    l    -   c   dobre'm
A   f   [pcs]  m     s    i    -   c   dobry'm
A   f   [pcs]  m     p    n    y   c   dobr<i'
A   f   [pcs]  m     p    n    n   c            dobre'
A   f   [pcs]  m     p    g    -   c   dobry'ch
A   f   [pcs]  m     p    d    -   c   dobry'm
A   f   [pcs]  m     p    a    -   c   dobre'
A   f   [pcs]  m     p    v    y   c   dobr<i'
A   f   [pcs]  m     p    v    n   c            dobre'
A   f   [pcs]  m     p    l    -   c   dobry'ch
A   f   [pcs]  m     p    i    -   c   dobry'mi
A   f   [pcs] [fn]   s    n    -   c   dobra',  dobre'
A   f   [pcs] [fn]   s    g    -   c   dobre',  dobre'ho
A   f   [pcs] [fn]   s    d    -   c   dobre',  dobre'mu
A   f   [pcs] [fn]   s    a    -   c   dobrou,  dobre'
A   f   [pcs] [fn]   s    v    -   c   dobra'!, dobre'
A   f   [pcs] [fn]   s    l    -   c   dobre',  dobre'm
A   f   [pcs] [fn]   s    i    -   c   dobrou,  dobr'ym
A   f   [pcs] [fn]   p    n    -   c   dobre',  dobra'

A   f   [pcs] [fn]   p    g    -   c   dobry'ch
A   f   [pcs] [fn]   p    d    -   c   dobry'm
A   f   [pcs] [fn]   p    a    -   c   dobre',  dobra'
A   f   [pcs] [fn]   p    v    -   c   dobre',  dobra'
A   f   [pcs] [fn]   p    l    -   c   dobry'ch
A   f   [pcs] [fn]   p    i    -   c   dobry'mi
A   f   [pcs] [fn]   d    i    -   c   dobry'ma
A   f     -   [mfn]  s    n    -   n   sta'r, sta'ra, sta'ro
A   f     -   [mfn]  s    a    -   n   sta'ra, sta'ru, sta'ro
A   f     -     m    p    n   [ny] n   sta'r<i, sta'ry
A   f     -   [fn]   p    n    -   n   sta'ry, sta'ra
A   f     -   [mfn]  p    a    -   n   sta'ry, sta'ry, sta'ra
*** **** ***** **** **** **** **** **** ======================================

4. Pronoun (P)
4.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           personal       p ja'                 osobni'
demonstrative  d ten                 ukazovaci'
indefinite     i ne<kdo              neurc<ite'
possessive     s jeji'               posesi'vni'
interrogative  q kdo                 tazaci'
relative       r jenz<               vztaz<ne' / relativni'
reflexive      x se                  zvratne'  / reflexi'vni'
negative       z nikdo               za'porna' / negativni'
general        g kaz<dy'             tota'lni'
- -------------- -------------- --------------------- ------------------------
2 Person         first          1 ja'                 prvni' osoba
second         2 ty                  druha' osoba
third          3 on                  tr<eti' osoba
- -------------- -------------- --------------------- ------------------------
3 Gender         masculine      m on                  masculinum
feminine       f ona                 femininum
neuter         n to                  neutrum
- -------------- -------------- --------------------- ------------------------
4 Number         singular       s ty                  singula'r
plural         p oni                 plura'l
dual           d te<ma               dua'l
- -------------- -------------- --------------------- ------------------------
5 Case           nominative     n ty                  nominativ
genitive       g tebe                genitiv
dative         d tobe<,ti            dativ
accusative     a tebe, te<           akuzativ
locative       l tobe<               loka'l
instrumental   i tebou               instrumenta'l
- -------------- -------------- - ------------------- ------------------------
6 Owner_Number   singular       s mu0j                singula'r
plural         p jejich              plura'l

- -------------- -------------- - ------------------- ------------------------
7 Owner_Gender   masculine      m jehoz<              masculinum
                 feminine       f jeji'hoz<           femininum
                 neuter         n jehoz<              neutrum
********************************* ------------------- ------------------------
8 Clitic         no             n tobe<               nepr<i'klonne'
                 yes            y ti                  pr<i'klonka
- -------------- -------------- - ------------------- ------------------------
9 Referent_Type  personal       p se, sebe, si        osobni' zvratne'
                 possessive     s svu0j               zvratne' posesi'vni'
- -------------- -------------- - ------------------- ------------------------
10Syntactic_Type nominal        n on                  syntakticke' substantivum
                 adjectival     a ktery'              syntakticke' adjektivum
- -------------- -------------- - ------------------- ------------------------
11Definiteness                  -
- -------------- -------------- - ------------------- ------------------------
12Animate        no             n ktery'              nez<ivotne'
                 yes            y kter<i', ktere'ho   z<ivotne'
- -------------- -------------- - ------------------- ------------------------
13Clitic_s  l.s. yes            y tys,ses             za'jmeno s kliticky'm -s
            l.s. no             n ty                  za'jmeno bez klitic. -s
- -------------- -------------- - ------------------- ------------------------
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person                  -
- -------------- -------------- -
16Owned_Number                  -
================ ================
Notes
1. Gender,  Number,  Case  and  Animate  correspond  to  the same
   categories  within Nouns.  They are  necessary for  the proper
   account of agreement of adjectival Pronouns with Nouns.
2. Type=reflexive ecompasses all reflexive pronouns (se, sebe, si
   svu0j) as well as "se" in  its role as the obligatory particle
   of  reflexive verbs.  Personal and  possessive reflexives  are
   further distinguished via the Referent_Type attribute.
   "se" in all its roles will be marked as the reflexive personal
   clitic pronoun.
3. Pronouns  are distinguished  between having  a (syntactically)
   nominal and  (syntactically) adjectival function.  All pronoun
   types except demonstrative and  possessive can be nominal, and
   all except personal can be adjectival.
5. Referent_Type  is  used  to  distinguish  personal  reflexives
   (which include  "se" in all its  functions) from the possesive
   reflexives ("svu0j").
6. Negative and general Pronouns ("general" Pronouns concern the
   Pronouns like  "vs<ichni" [E.. "all"],  "kaz<dy'" [E. "every"]
   etc.)  are important  from  the  viewpoint of  their syntactic
   distribution.
7. The  Clitic attribute  distinguishes clitical  vs. nonclitical
   pronominal forms, e.g. "ti" vs. "tobe<".

8. The 'yes' value of the Clitic_s attribute denotes a pronominal
  form having the clitic morpheme 's' appended as a suffix. This
  morpheme expresses  2nd Person singular  present Tense of  the
  auxiliary Verb  "by't" (E. "to be").  There is no intermediate
  hyphen between the Pronoun and the 's' morpheme.
9. Owner_Number and  Owner_Gender concern the  possessor's number
  and gender, respectively.

4.2 Combinations
*** **** ***** ***** **** **** **** **** ---- --- --- ----- ---- =============
PoS Type Pers  Gend  Numb Case Ow_N Ow_G Clit Ref Syn Anim  Cl_s  Examples
*** **** ***** ***** **** **** **** **** ---- --- --- ----- ---- =============
P    x    -     -     -  [ga]   -    -    y   p   n    -    n   se
P    x    -     -     -  [ga]   -    -    y   p   n    -    y   ses
P    x    -     -     -  [ga]   -    -    n   p   n             sebe
P    x    -     -     -   d     -    -    y   p   n    -    n   si
P    x    -     -     -   d     -    -    y   p   n    -    y   sis
P    x    -     -     -  [dl]   -    -    n   p   n    -   [ny] sobe<
P    x    -     -     -   i     -    -    n   p   n    -   [ny] sebou
P    x    -   [mfn]  all any    -    -    n   s   a  [ny-]  n   svu0j
P    p    1     -   [sp] all    -    -  [ny]  -   n    -   [ny] ja', my
P    p    2     -   [sp] all    -    -  [ny]  -   n    -    n   ty, vy
P    p    2     -     s   n     -    -    n   -   n    -    y   tys
P    p    3   [mfn] [sp] all    -    -  [ny]  -   n    -   [ny] on, oni,
                                                                ony, ona
P    s    1   [mfn]  all all    s    -    n   -   a  [ny-] [ny] mu0j
P    s    2   [mfn]  all [na]   s    -    n   -   a  [ny-] [ny] tvu0j
P    s    3     -     -   -     s  [mn]   n   -   a    -   [ny] jeho
P    s    3   [mfn]  all all    s    f    n   -   a    -   [ny] jeji'
P    s    3     -     -   -     p  [mfn]  n   -   a    -    -   jejich
P    d    -   [mfn]  all all    -    -    n   -   a  [ny-] [ny] 1)
P    i    -   [mn]    s  all    -    -    n   -   n    -    -   2)
P    i    -   [mfn] [sp] all    -    -    n   -   a  [ny-]  -   3)
P   [qr]  -   [mn]    s  all    -    -    n   -   n    -    -   4)
P   [qr]  -   [mfn] [sp] all    -    -    n   -   a  [ny-]  -   5)
P    z    -   [mn]    s  all    -    -    n   -   n    -    -   6)
P    z    -   [mfn] [sp] all    -    -    n   -   a  [ny-]  -   7)
P    g    -   [mfn] [sp] all    -    -    n   -   n  [ny-]  -   8)
*** **** ***** ***** **** **** **** **** ---- --- --- ----- ---- =============
1) ten, tento, tamten, tenhle, tenhleten, takovy', onaky'
2) ne<kdo, ne<jaky', kdovi'kdo, leckdo, leccos, ne<co
3) jaky', ktery', kdovi'jaky', lecjaky'
4) kdo, co
5) jaky', ktery', jenz<, coz<
6) nikdo, nic
7) z<a'dny', nijaky'
8) vs<echen, kaz<dy'

5. Determiner (D)

Not applicable.

6. Article (T)
Not applicable.

7. Adverb (R)
7.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           general        g dobr<e              pr<i'slovce
- -------------- -------------- --------------------- ------------------------
2 Degree         positive       p ma'lo               pozitiv / 1. stupen<
                 comparative    c me'ne<              komparativ / 2. stupen<
                 superlative    s nejme'ne<           superlativ / 3. stupen<
* ************** ************** * ------------------- ------------------------
3 Clitic                        -
- -------------- -------------- -
4 Number                        -
- -------------- -------------- -
5 Person                        -
= ============== ============== =
Notes
1. The  classification  given  above  is  greatly  simplified  as
   compared  to traditional  Czech grammars.  There can  be other
   classifications as well, eg.:
   (a) syntactic: Adverb modifying a Verb
                  Adverb modifying an Adjective
                  Adverb modifying another Adverb
                  Adverb modifying a Noun from the left
                  Adverb modifying a Noun from the right
   (b) semantic   temporal
                  spatial
                  qualitative
                   ...
   We refrained from this typology here.
2. Particles form a separate part  of speech category (see below)
   as is customary in Czech grammars.
3. The Adjectives  which have no  degrees of comparison  have the
   Degree value equal to p(ositive) similarly as Adjectives.

7.2 Combinations
*** **** **** ================================================================

Pos Type Deg  Examples
*** **** **** ================================================================
 R   g    p   dobr<e
 R   g    c   le'pe
 R   g    s   nejle'pe
 R   g    p   tady
*** **** **** ================================================================

8. Adposition (S)
8.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           preposition    p nad                 pr<edloz<ka
- -------------- -------------- - ------------------- ------------------------
2 Formation      simple         s na, v               predloz<ka (prava')
                 compound       c nan<, oc<           kontrakce pr<edl. + za'jm.
* ************** ************** * ------------------- ------------------------
3 Case
  (req.by prep.) genitive       g bez                 genitiv
                 dative         d proti               dativ
                 accusative     a pro                 akuzativ
                 locative       l v                   loka'l
                 instrumental   i s                   instrumenta'l
- -------------- -------------- - ---- -------------- ------------------------
4 Clitic                        -
=================================
Notes
1. Czech has only Prepositions, no postpositions.
2. For the  disambiguation of word forms  belonging to declension
   parts of speech it seems  necessary to include the information
   about the case which each Preposition requires.
   3.  A preposition  can be  contracted with  a pronoun;  such a
   preposition has Formation=c(ompound).

8.2 Combinations
*** **** **** ---- ===========================================================
PoS Type Form Case Examples
*** **** **** ---- ===========================================================
 S    p    s    g  bez, krome<
 S    p    s    d  k, proti
 S    p    s    a  pro, mimo
 S    p    s    l  o, v, pr<i
 S    p    s    i  s, pod
 S    p    c    -  nan<, zac<
*** **** **** ---- ===========================================================

9. Conjunction (C)
9.1 Lexicon
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           coordinating   c a, ani              spojka sour<adici'
                 subordinating  s aby, protoz<e       spojka podr<adici'
* ************** ************** * ------------------- ------------------------
2 Formation                     -
- -------------- -------------- -
3 Coord_Type                    -
- -------------- -------------- -
4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- - ------------------- ------------------------
6 Number         singular       s kdybych             spojka 'kdyby'+sg. od 'by'
                 plural         p kdybyste            spojka 'kdyby'+pl. od 'by'
- -------------- -------------- - ------------------- ------------------------
7 Person         first          1 abych               spojka 'aby'+1.os. od 'by'
                 second         2 abyste              spojka 'aby'+2.os. od 'by'
                 third          3 aby                 spojka 'aby'+3.os. od 'by'
================================================================================
Notes
1. Conjunctions "abych", "abys", ..., "kdybych", ... are composed
   of  the Conjunction  "aby" or  "kdyby" +  conditional particle
   ("bych", "bys",  "by", "bychom", "byste").  These Conjunctions
   are therefore specified for Person  and Number for the sake of
   agreement relation.
2. The class of two-part Conjunctions has not been introduced.

9.2 Combinations
*** **** ---- ---- ===========================================================
PoS Type Numb Pers Examples
*** **** ---- ---- ===========================================================
 C    c    -   -   a, i, ba, nebot<, ani
 C    s    -   -   z<e, protoz<e, zda
 C    s    -   3   aby, kdyby
 C    s    s  [12] abych, abys, kdybych, kdybys
 C    s    p  [12] abychom, abyste, kdybychom, kdybyste
*** **** =====================================================================

10. Numeral (M)

= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type           cardinal       c dva                 c<i'slovka za'kladni'
                 ordinal        o pa'ty'              c<i'slovka r<adova'
                 multiple       m dvakra't            c<i'slovka na'sobna'
                 special        s dvoji'              c<i'slovka druhova'
- -------------- -------------- - ------------------- ------------------------
2 Gender         masculine      m dva                 masculinum
                 feminine       f dve<                femininum
                 neuter         n dve<                neutrum
- -------------- -------------- - ------------------- ------------------------
3 Number         singular       s druhy'              singula'r
                 plural         p druzi'              plura'l
                 dual           d druhy'ma            dua'l
- -------------- -------------- --------------------- ------------------------
4 Case           nominative     n dva, dve<           nominativ
                 genitive       g dvou                genitiv
                 dative         d dve<ma              dativ
                 accusative     a dva, dve<           akuzativ
                 locative       l dvou                loka'l
                 instrumental   i dve<ma              instrumenta'l
- -------------- -------------- - ------------------- ------------------------
5 Form           digit          d 5                   c<i'slo (arabske')
                 roman          r MCMVIII             c<i'slo (r<i'mske')
                 letter         l s<edesa't           c<i'slovka slovem
* ************** ************** * ------------------- ------------------------
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- - ------------------- ------------------------
8 Class          definite1      1 jeden
                 definite2      2 dva
                 definite34     3 tr<i
                 definite       f pe<t
                 demonstrative  d tolik
                 indefinite     i ne<kolik
                 interrogative  q kolik
                 relative       r kolik
---------------- -------------- - ------------------- ------------------------
9 Animate        no             n pa'te'              nez<ivotne'
                 yes            y pa'ti'              z<ivotne'
- -------------- -------------- - ------------------- ------------------------
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
=================================
Notes
1. Numerals have been specified as a separate category because of

   their specific  syntactic distribution. We  have specified two
   syntactic classifications by means  of the attributes Type and
   Class;  they  concern  different  syntactic distributions. For
   instance "ne<kolik" (E. "several") will be characterized as:
   Type:  cardinal
   Class: indefinite
2. Among   the  definite   numbers  there   are  four  subclasses
   (definite1, definite2,  definite34, definite) which  differ in
   the syntactic distribution and contain the following Numerals:
   {1}, {2}, {3,4}, {5,6,...}
3. Gender,  Number and  Case  correspond  to the  same categories
   within  Nouns. They  are necessary  for the  proper account of
   agreement of Numerals with Nouns.
4. In direct cases (nominative, accusative) the cardinal numerals
   belonging to class with one  of the values {fdiqr} (see above)
   behave like neuter Nouns in singular, whereas in oblique cases
   they behave like adjectival attributes in plural.

Combinations
============
*** **** ****  **** **** ---- ------ ----- ===================================
PoS Type Gend  Numb Case Form Class  Anim  Examples
*** **** ****  **** **** ---- ------ ----- ===================================
 M    -    -     -    -    d  [123f]   -   56
 M    -    -     -    -    r  [123f]   -   MVIII
 M    c  [mfn]   s   any   l     1     -   jeden, jedna, jedno
 M    c  [mfn]   p   any   l     2     -   dva, dve<
 M    c    -     p   any   l     3     -   tr<i, c<tyr<i
 M    c    n    [sp] any   l     f     -   pe<t, deset, padesa't
 M    c    n    [sp] any   l  [diqr]   -   tolik, ne<kolik, kolik
 M    o  [mfn] [spd] any   l    any  [ny-] prvni', druhy', ne<kolika'ty'
 M    m    -     -    -    l    any    -   dvakra't, pe<tkra't, ne<kolikra't
 M    s  [mfn] [spd]  n    l    any  [ny-] dvoji', troje, patery'
*** **** ****  **** **** ---- ------ ----- ===================================
Note: In the Combinations above, 'any' is a variable standing for
      any admissible value.

11. Interjection (I)
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type                          - au!, och!           citoslovce
- -------------- -------------- - ------------------- ------------------------
2 Formation                     -
=================================

12. Residual (X)

= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
                                - sic, a+b            Zby'vaji'ci' elementy
= ============== ============== =

13. Abbreviation (Y)
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Syntactic_Type                - atd.                zkratka
- -------------- -------------- - ------------------- ------------------------
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
- -------------- -------------- -
5 Definiteness                  -
=================================

14. Particle (Q)
= ============== ============== = =================== ========================
P ATT            VAL            C Example             Czech term
= ============== ============== = =================== ========================
1 Type                          - ano, ne             partikule / c<a'stice
- -------------- -------------- - ------------------- ------------------------
2 Formation                     -
- -------------- -------------- -
3 Clitic                        -
=================================

COP project 106 multext-East  Deliverable D1.1 F— Estonian  66

3.3 Application to Estonian

The application to Estonian has been elaborated by H.J. Kaalep.

Notes:

1. Two additional tables are provided for every POS: one for definitions and one for examples. For nouns, adjectives, numerals and pronouns the table of examples contains more rows than the table of definitions. This is because in Estonian there are two equal ways for forming 11–12 plural cases for these POS and both are given.

2. The ordering of attributes is the same as in the synoptic tables of the proposal.

1. Nouns (N)
1.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           common         c jalg
                 proper         p London
- -------------- -------------- - ------------
2 Gender                        -
- -------------- -------------- - ------------
3 Number         singular       s jalg
                 plural         p jalad
- -------------- -------------- - ------------
4 Case           nominative     n jalg
                 genitive       g jala
           l.s.  partitive      1 jalga
                 illative       x jalasse
                 inessive       2 jalas
                 elative        e jalast
                 allative       t jalale
                 adessive       3 jalal
                 ablative       b jalalt
           l.s.  translative    4 jalaks
                 terminative    9 jalani
                 essive         w jalana
           l.s.  abessive       5 jalata
           l.s.  komitative     k jalaga
           l.s.  aditive        7 jalga
* ***************************** * ************
5 Definiteness                  -
- -------------- -------------- - ------------
6 Clitic                        -
- -------------- -------------- - ------------
7 Animate                       -
- -------------- -------------- - ------------
8 Owner_Number                  -

- -------------- -------------- - ------------
9 Owner_Person                  -
---------------- -------------- - ------------
10Owned_Number                  -
================================= ============

1.2. Definitions
Lexicon      Definitions
------------------------------------
Nc-sn        Noun common singular nominative
Nc-sg        Noun common singular genitive
Nc-s1        Noun common singular partitive
Nc-sx        Noun common singular illative
Nc-s2        Noun common singular inessive
Nc-se        Noun common singular elative
Nc-st        Noun common singular allative
Nc-s3        Noun common singular adessive
Nc-sb        Noun common singular ablative
Nc-s4        Noun common singular translative
Nc-s9        Noun common singular terminative
Nc-sw        Noun common singular essive
Nc-s5        Noun common singular abessive
Nc-sk        Noun common singular komitative
Nc-s7        Noun common singular aditive
Nc-pn        Noun common plural nominative
Nc-pg        Noun common plural genitive
Nc-p1        Noun common plural partitive
Nc-px        Noun common plural illative
Nc-p2        Noun common plural inessive
Nc-pe        Noun common plural elative
Nc-pt        Noun common plural allative
Nc-p3        Noun common plural adessive
Nc-pb        Noun common plural ablative
Nc-p4        Noun common plural translative
Nc-p9        Noun common plural terminative
Nc-pw        Noun common plural essive
Nc-p5        Noun common plural abessive
Nc-pk        Noun common plural komitative
Np-sn        Noun proper singular nominative
Np-sg        Noun proper singular genitive
Np-s1        Noun proper singular partitive
Np-sx        Noun proper singular illative
Np-s2        Noun proper singular inessive
Np-se        Noun proper singular elative
Np-st        Noun proper singular allative
Np-s3        Noun proper singular adessive
Np-sb        Noun proper singular ablative
Np-s4        Noun proper singular translative
Np-s9        Noun proper singular terminative
Np-sw        Noun proper singular essive

Np-s5        Noun proper singular abessive
Np-sk        Noun proper singular komitative
Np-s7        Noun proper singular aditive
Np-pn        Noun proper plural nominative
Np-pg        Noun proper plural genitive
Np-p1        Noun proper plural partitive
Np-px        Noun proper plural illative
Np-p2        Noun proper plural inessive
Np-pe        Noun proper plural elative
Np-pt        Noun proper plural allative
Np-p3        Noun proper plural adessive
Np-pb        Noun proper plural ablative
Np-p4        Noun proper plural translative
Np-p9        Noun proper plural terminative
Np-pw        Noun proper plural essive
Np-p5        Noun proper plural abessive
Np-pk        Noun proper plural komitative

1.3. Examples
Lexicon      Examples
----------------------
Nc-sn        jalg
Nc-sg        jala
Nc-s1        jalga
Nc-sx        jalasse
Nc-s2        jalas
Nc-se        jalast
Nc-st        jalale
Nc-s3        jalal
Nc-sb        jalalt
Nc-s4        jalaks
Nc-s9        jalani
Nc-sw        jalana
Nc-s5        jalata
Nc-sk        jalaga
Nc-s7        jalga
Nc-pn        jalad
Nc-pg        jalgade
Nc-p1        jalgasid
Nc-px        jalgadesse
Nc-p2        jalgades
Nc-pe        jalgadest
Nc-pt        jalgadele
Nc-p3        jalgadel
Nc-pb        jalgadelt
Nc-p4        jalgadeks
Nc-p9        jalgadeni
Nc-pw        jalgadena
Nc-p5        jalgadeta
Nc-pk        jalgadega
Nc-pg        jalge

Nc-p1        jalgu
Nc-px        jalusse
Nc-p2        jalus
Nc-pe        jalust
Nc-pt        jalule
Nc-p3        jalul
Nc-pb        jalult
Nc-p4        jaluks
Np-sn        London
Np-sg        Londoni
Np-s1        Londonit
Np-sx        Londonisse
Np-s2        Londonis
Np-se        Londonist
Np-st        Londonile
Np-s3        Londonil
Np-sb        Londonilt
Np-s4        Londoniks
Np-s9        Londonini
Np-sw        Londonina
Np-s5        Londonita
Np-sk        Londoniga
Np-s7        Londoni
Np-pn        Londonid
Np-pg        Londonite
Np-p1        Londoneid
Np-px        Londonitesse
Np-p2        Londonites
Np-pe        Londonitest
Np-pt        Londonitele
Np-p3        Londonitel
Np-pb        Londonitelt
Np-p4        Londoniteks
Np-p9        Londoniteni
Np-pw        Londonitena
Np-p5        Londoniteta
Np-pk        Londonitega
Np-px        Londoneisse
Np-p2        Londoneis
Np-pe        Londoneist
Np-pt        Londoneile
Np-p3        Londoneil
Np-pb        Londoneilt
Np-p4        Londoneiks
Np-p9        Londoneini
Np-pw        Londoneina
Np-p5        Londoneita
Np-pk        Londoneiga

2. Verbs (V)

2.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           main           m lugema
                 auxiliary      a olema
                 modal          o saama
- -------------- -------------- - ------------
2 VForm          indicative     i loen
                 imperative     m loe
                 conditional    c loeksin
                 infinitive     n lugeda
                 participle     p lugev
                 gerund         g lugedes
                 supine         u lugema
           l.s.  quotative      q lugevat
- -------------- -------------- - ------------
3 Tense          present        p loen
                 imperfect      i lugesin
                 past           s lugenuvat
- -------------- -------------- - ------------
4 Person         first          1 loen
                 second         2 loed
                 third          3 loeb
- -------------- -------------- - ------------
5 Number         singular       s loen
                 plural         p loeme
- -------------- -------------- - ------------
6 Gender                        -
********************************* ************
7 Voice          active         a loen
                 passive        p loetakse
- -------------- -------------- - ------------
8 Negative       no             n loen
                 yes            y (ei) + loeta
- -------------- -------------- - ------------
9 Definiteness                  -
- -------------- -------------- - ------------
10Clitic                        -
- -------------- -------------- - ------------
11Case
                 illative       x lugema
                 inessive       2 lugemas
                 elative        e lugemast
                 translative    4 lugemaks
                 abessive       5 lugemata
- -------------- -------------- - ------------
12Animate                       -
- -------------- -------------- - ------------
13Clitic_s                      -
================================= ============

2.2. Definitions
Lexicon        Definition
-------------------------
Vmip1s-an---   Verb main indic present first singular active afirmative
Vmip2s-an---   Verb main indic present second singular active afirmative
Vmip3s-an---   Verb main indic present third singular active afirmative
Vmip1p-an---   Verb main indic present first plural active afirmative
Vmip2p-an---   Verb main indic present second plural active afirmative
Vmip3p-an---   Verb main indic present third plural active afirmative
Vmip1s-ay---   Verb main indic present first singular active negative
Vmip2s-ay---   Verb main indic present second singular active negative
Vmip3s-ay---   Verb main indic present third singular active negative
Vmip1p-ay---   Verb main indic present first plural active negative
Vmip2p-ay---   Verb main indic present second plural active negative
Vmip3p-ay---   Verb main indic present third plural active negative
Vmip---pn---   Verb main indic present passive afirmative
Vmip---py---   Verb main indic present passive negative
Vmii1s-an---   Verb main indic imperfect first singular active afirmative
Vmii2s-an---   Verb main indic imperfect second singular active afirmative
Vmii3s-an---   Verb main indic imperfect third singular active afirmative
Vmii1p-an---   Verb main indic imperfect first plural active afirmative
Vmii2p-an---   Verb main indic imperfect second plural active afirmative
Vmii3p-an---   Verb main indic imperfect third plural active afirmative
Vmii---pn---   Verb main indic imperfect passive afirmative
Vmmp2s-an---   Verb main imper present second singular active afirmative
Vmmp3s-an---   Verb main imper present third singular active afirmative
Vmmp1p-an---   Verb main imper present first plural active afirmative
Vmmp2p-an---   Verb main imper present second plural active afirmative
Vmmp3p-an---   Verb main imper present third plural active afirmative
Vmmp---pn---   Verb main imper present passive afirmative
Vmcp1s-an---   Verb main condit present first singular active afirmative
Vmcp2s-an---   Verb main condit present second singular active afirmative
Vmcp3s-an---   Verb main condit present third singular active afirmative
Vmcp1p-an---   Verb main condit present first plural active afirmative
Vmcp2p-an---   Verb main condit present second plural active afirmative
Vmcp3p-an---   Verb main condit present third plural active afirmative
Vmcp---pn---   Verb main condit present passive afirmative
Vmcs1s-an---   Verb main condit past first singular active afirmative
Vmcs2s-an---   Verb main condit past second singular active afirmative
Vmcs3s-an---   Verb main condit past third singular active afirmative
Vmcs1p-an---   Verb main condit past first plural active afirmative
Vmcs2p-an---   Verb main condit past second plural active afirmative
Vmcs3p-an---   Verb main condit past third plural active afirmative
Vmcs---pn---   Verb main condit past passive afirmative
Vmn-----n---   Verb main infinit
Vmpp---an---   Verb main partic present active afirmative
Vmpp---pn---   Verb main partic present passive afirmative

Vmps---an---   Verb main partic past active afirmative
Vmps---pn---   Verb main partic past passive afirmative
Vmg-----n---   Verb main gerund
Vms----an--x   Verb main supine active afirmative illative
Vms----an--2   Verb main supine active afirmative inessive
Vms----an--e   Verb main supine active afirmative elative
Vms----an--4   Verb main supine active afirmative translative
Vms----an--5   Verb main supine active afirmative abessive
Vms----pn---   Verb main supine passive afirmative
Vmqp1s-an---   Verb main quotat present first singular active afirmative
Vmqp2s-an---   Verb main quotat present second singular active afirmative
Vmqp3s-an---   Verb main quotat present third singular active afirmative
Vmqp1p-an---   Verb main quotat present first plural active afirmative
Vmqp2p-an---   Verb main quotat present second plural active afirmative
Vmqp3p-an---   Verb main quotat present third plural active afirmative
Vmqp---pn---   Verb main quotat present passive afirmative
Vmqs1s-an---   Verb main quotat past first singular active afirmative
Vmqs2s-an---   Verb main quotat past second singular active afirmative
Vmqs3s-an---   Verb main quotat past third singular active afirmative
Vmqs1p-an---   Verb main quotat past first plural active afirmative
Vmqs2p-an---   Verb main quotat past second plural active afirmative
Vmqs3p-an---   Verb main quotat past third plural active afirmative
Vmqs---pn---   Verb main quotat past passive afirmative
Vm------n---   Verb main afirmative
Vai-----y---   Verb aux indic negative
Vaip1s-an---   Verb aux indic present first singular active afirmative
Vaip2s-an---   Verb aux indic present second singular active afirmative
Vaip3s-an---   Verb aux indic present third singular active afirmative
Vaip1p-an---   Verb aux indic present first plural active afirmative
Vaip2p-an---   Verb aux indic present second plural active afirmative
Vaip3p-an---   Verb aux indic present third plural active afirmative
Vaip1s-ay---   Verb aux indic present first singular active negative
Vaip2s-ay---   Verb aux indic present second singular active negative
Vaip3s-ay---   Verb aux indic present third singular active negative
Vaip1p-ay---   Verb aux indic present first plural active negative
Vaip2p-ay---   Verb aux indic present second plural active negative
Vaip3p-ay---   Verb aux indic present third plural active negative
Vaip---pn---   Verb aux indic present passive afirmative
Vaip---py---   Verb aux indic present passive negative
Vaii1s-an---   Verb aux indic imperfect first singular active afirmative
Vaii2s-an---   Verb aux indic imperfect second singular active afirmative
Vaii3s-an---   Verb aux indic imperfect third singular active afirmative
Vaii1p-an---   Verb aux indic imperfect first plural active afirmative
Vaii2p-an---   Verb aux indic imperfect second plural active afirmative
Vaii3p-an---   Verb aux indic imperfect third plural active afirmative
Vaii1s-ay---   Verb aux indic imperfect first singular active negative
Vaii2s-ay---   Verb aux indic imperfect second singular active negative
Vaii3s-ay---   Verb aux indic imperfect third singular active negative

Vaii1p-ay---   Verb aux indic imperfect first plural active negative
Vaii2p-ay---   Verb aux indic imperfect second plural active negative
Vaii3p-ay---   Verb aux indic imperfect third plural active negative
Vaii---pn---   Verb aux indic imperfect passive afirmative
Vamp2s-an---   Verb aux imper present second singular active afirmative
Vamp3s-an---   Verb aux imper present third singular active afirmative
Vamp1p-an---   Verb aux imper present first plural active afirmative
Vamp2p-an---   Verb aux imper present second plural active afirmative
Vamp3p-an---   Verb aux imper present third plural active afirmative
Vamp---pn---   Verb aux imper present passive afirmative
Vamp2s-ay---   Verb aux imper present second singular active negative
Vamp3s-ay---   Verb aux imper present third singular active negative
Vamp1p-ay---   Verb aux imper present first plural active negative
Vamp2p-ay---   Verb aux imper present second plural active negative
Vamp3p-ay---   Verb aux imper present third plural active negative
Vamp---py---   Verb aux imper present passive negative
Vac-----y---   Verb aux condit negative
Vacp1s-an---   Verb aux condit present first singular active afirmative
Vacp2s-an---   Verb aux condit present second singular active afirmative
Vacp3s-an---   Verb aux condit present third singular active afirmative
Vacp1p-an---   Verb aux condit present first plural active afirmative
Vacp2p-an---   Verb aux condit present second plural active afirmative
Vacp3p-an---   Verb aux condit present third plural active afirmative
Vacp1s-ay---   Verb aux condit present first singular active negative
Vacp2s-ay---   Verb aux condit present second singular active negative
Vacp3s-ay---   Verb aux condit present third singular active negative
Vacp1p-ay---   Verb aux condit present first plural active negative
Vacp2p-ay---   Verb aux condit present second plural active negative
Vacp3p-ay---   Verb aux condit present third plural active negative
Vacp---pn---   Verb aux condit present passive afirmative
Vacs1s-an---   Verb aux condit past first singular active afirmative
Vacs2s-an---   Verb aux condit past second singular active afirmative
Vacs3s-an---   Verb aux condit past third singular active afirmative
Vacs1p-an---   Verb aux condit past first plural active afirmative
Vacs2p-an---   Verb aux condit past second plural active afirmative
Vacs3p-an---   Verb aux condit past third plural active afirmative
Vacs1s-ay---   Verb aux condit past first singular active negative
Vacs2s-ay---   Verb aux condit past second singular active negative
Vacs3s-ay---   Verb aux condit past third singular active negative
Vacs1p-ay---   Verb aux condit past first plural active negative
Vacs2p-ay---   Verb aux condit past second plural active negative
Vacs3p-ay---   Verb aux condit past third plural active negative
Vacs---pn---   Verb aux condit past passive afirmative
Van-----n---   Verb aux infinit
Vapp---an---   Verb aux partic present active afirmative
Vapp---pn---   Verb aux partic present passive afirmative
Vaps---an---   Verb aux partic past active afirmative
Vaps---pn---   Verb aux partic past passive afirmative

Vag-----n---   Verb aux gerund
Vas----an--x   Verb aux supine active afirmative illative
Vas----an--2   Verb aux supine active afirmative inessive
Vas----an--e   Verb aux supine active afirmative elative
Vas----an--4   Verb aux supine active afirmative translative
Vas----an--5   Verb aux supine active afirmative abessive
Vas----pn---   Verb aux supine passive afirmative
Vaq-----y---   Verb aux quotat negative
Vaqp1s-an---   Verb aux quotat present first singular active afirmative
Vaqp2s-an---   Verb aux quotat present second singular active afirmative
Vaqp3s-an---   Verb aux quotat present third singular active afirmative
Vaqp1p-an---   Verb aux quotat present first plural active afirmative
Vaqp2p-an---   Verb aux quotat present second plural active afirmative
Vaqp3p-an---   Verb aux quotat present third plural active afirmative
Vaqp1s-ay---   Verb aux quotat present first singular active negative
Vaqp2s-ay---   Verb aux quotat present second singular active negative
Vaqp3s-ay---   Verb aux quotat present third singular active negative
Vaqp1p-ay---   Verb aux quotat present first plural active negative
Vaqp2p-ay---   Verb aux quotat present second plural active negative
Vaqp3p-ay---   Verb aux quotat present third plural active negative
Vaqp---pn---   Verb aux quotat present passive afirmative
Vaqs1s-an---   Verb aux quotat past first singular active afirmative
Vaqs2s-an---   Verb aux quotat past second singular active afirmative
Vaqs3s-an---   Verb aux quotat past third singular active afirmative
Vaqs1p-an---   Verb aux quotat past first plural active afirmative
Vaqs2p-an---   Verb aux quotat past second plural active afirmative
Vaqs3p-an---   Verb aux quotat past third plural active afirmative
Vaqs---pn---   Verb aux quotat past passive afirmative
Voip1s-an---   Verb modal indic present first singular active afirmative
Voip2s-an---   Verb modal indic present second singular active afirmative
Voip3s-an---   Verb modal indic present third singular active afirmative
Voip1p-an---   Verb modal indic present first plural active afirmative
Voip2p-an---   Verb modal indic present second plural active afirmative
Voip3p-an---   Verb modal indic present third plural active afirmative
Voip---pn---   Verb modal indic present passive afirmative
Voip---py---   Verb modal indic present passive negative
Voii1s-an---   Verb modal indic imperfect first singular active afirmative
Voii2s-an---   Verb modal indic imperfect second singular active afirmative
Voii3s-an---   Verb modal indic imperfect third singular active afirmative
Voii1p-an---   Verb modal indic imperfect first plural active afirmative
Voii2p-an---   Verb modal indic imperfect second plural active afirmative
Voii3p-an---   Verb modal indic imperfect third plural active afirmative
Voii---pn---   Verb modal indic imperfect passive afirmative
Vomp2s-an---   Verb modal imper present second singular active afirmative
Vomp3s-an---   Verb modal imper present third singular active afirmative
Vomp1p-an---   Verb modal imper present first plural active afirmative
Vomp2p-an---   Verb modal imper present second plural active afirmative
Vomp3p-an---   Verb modal imper present third plural active afirmative
Vomp---pn---   Verb modal imper present passive afirmative

Vocp1s-an---   Verb modal condit present first singular active afirmative
Vocp2s-an---   Verb modal condit present second singular active afirmative
Vocp3s-an---   Verb modal condit present third singular active afirmative
Vocp1p-an---   Verb modal condit present first plural active afirmative
Vocp2p-an---   Verb modal condit present second plural active afirmative
Vocp3p-an---   Verb modal condit present third plural active afirmative
Vocp---pn---   Verb modal condit present passive afirmative
Vocs1s-an---   Verb modal condit past first singular active afirmative
Vocs2s-an---   Verb modal condit past second singular active afirmative
Vocs3s-an---   Verb modal condit past third singular active afirmative
Vocs1p-an---   Verb modal condit past first plural active afirmative
Vocs2p-an---   Verb modal condit past second plural active afirmative
Vocs3p-an---   Verb modal condit past third plural active afirmative
Vocs---pn---   Verb modal condit past passive afirmative
Von-----n---   Verb modal infinit
Vopp---an---   Verb modal partic present active afirmative
Vopp---pn---   Verb modal partic present passive afirmative
Vops---an---   Verb modal partic past active afirmative
Vops---pn---   Verb modal partic past passive afirmative
Vog-----n---   Verb modal gerund
Vos----an--x   Verb modal supine active afirmative illative
Vos----an--2   Verb modal supine active afirmative inessive
Vos----an--e   Verb modal supine active afirmative elative
Vos----an--4   Verb modal supine active afirmative translative
Vos----an--5   Verb modal supine active afirmative abessive
Vos----pn---   Verb modal supine passive afirmative
Voqp1s-an---   Verb modal quotat present first singular active afirmative
Voqp2s-an---   Verb modal quotat present second singular active afirmative
Voqp3s-an---   Verb modal quotat present third singular active afirmative
Voqp1p-an---   Verb modal quotat present first plural active afirmative
Voqp2p-an---   Verb modal quotat present second plural active afirmative
Voqp3p-an---   Verb modal quotat present third plural active afirmative
Voqp---pn---   Verb modal quotat present passive afirmative
Voqs1s-an---   Verb modal quotat past first singular active afirmative
Voqs2s-an---   Verb modal quotat past second singular active afirmative
Voqs3s-an---   Verb modal quotat past third singular active afirmative
Voqs1p-an---   Verb modal quotat past first plural active afirmative
Voqs2p-an---   Verb modal quotat past second plural active afirmative
Voqs3p-an---   Verb modal quotat past third plural active afirmative
Voqs---pn---   Verb modal quotat past passive afirmative

2.3. Examples
Lexicon         Example
-----------------------

Vmip1s-an---    loen
Vmip2s-an---    loed
Vmip3s-an---    loeb
Vmip1p-an---    loeme
Vmip2p-an---    loete
Vmip3p-an---    loevad
Vmip1s-ay---    (ei)+l&auml;he
Vmip2s-ay---    (ei)+l&auml;he
Vmip3s-ay---    (ei)+l&auml;he
Vmip1p-ay---    (ei)+l&auml;he
Vmip2p-ay---    (ei)+l&auml;he
Vmip3p-ay---    (ei)+l&auml;he
Vmip---pn---    loetakse
Vmip---py---    (ei)+loeta
Vmii1s-an---    lugesin
Vmii2s-an---    lugesid
Vmii3s-an---    luges
Vmii1p-an---    lugesime
Vmii2p-an---    lugesite
Vmii3p-an---    lugesid
Vmii---pn---    loeti
Vmmp2s-an---    loe
Vmmp3s-an---    lugegu
Vmmp1p-an---    lugegem
Vmmp2p-an---    lugege
Vmmp3p-an---    lugegu
Vmmp---pn---    loetagu
Vmcp1s-an---    loeksin
Vmcp2s-an---    loeksid
Vmcp1p-an---    loeksime
Vmcp2p-an---    loeksite
Vmcp3p-an---    loeksid
Vmcp1s-an---    loeks
Vmcp2s-an---    loeks
Vmcp3s-an---    loeks
Vmcp1p-an---    loeks
Vmcp2p-an---    loeks
Vmcp3p-an---    loeks
Vmcp---pn---    loetaks
Vmcs1s-an---    lugenuksin
Vmcs2s-an---    lugenuksid
Vmcs1p-an---    lugenuksime
Vmcs2p-an---    lugenuksite
Vmcs3p-an---    lugenuksid
Vmcs1s-an---    lugenuks
Vmcs2s-an---    lugenuks
Vmcs3s-an---    lugenuks
Vmcs1p-an---    lugenuks
Vmcs2p-an---    lugenuks
Vmcs3p-an---    lugenuks
Vmcs---pn---    loetuks

Vmn-----n---    lugeda
Vmpp---an---    lugev
Vmpp---pn---    loetav
Vmps---an---    lugenud
Vmps---pn---    loetud
Vmg-----n---    lugedes
Vms----an--x    lugema
Vms----an--2    lugemas
Vms----an--e    lugemast
Vms----an--4    lugemaks
Vms----an--5    lugemata
Vms----pn---    loetama
Vmqp1s-an---    lugevat
Vmqp2s-an---    lugevat
Vmqp3s-an---    lugevat
Vmqp1p-an---    lugevat
Vmqp2p-an---    lugevat
Vmqp3p-an---    lugevat
Vmqp---pn---    loetavat
Vmqs1s-an---    lugenuvat
Vmqs2s-an---    lugenuvat
Vmqs3s-an---    lugenuvat
Vmqs1p-an---    lugenuvat
Vmqs2p-an---    lugenuvat
Vmqs3p-an---    lugenuvat
Vmqs---pn---    loetuvat
Vm------n---    kuulukse
Vai-----y---    ei
Vaip1s-an---    olen
Vaip2s-an---    oled
Vaip3s-an---    on
Vaip1p-an---    oleme
Vaip2p-an---    olete
Vaip3p-an---    on
Vaip1s-ay---    pole
Vaip2s-ay---    pole
Vaip3s-ay---    pole
Vaip1p-ay---    pole
Vaip2p-ay---    pole
Vaip3p-ay---    pole
Vaip---pn---    ollakse
Vaip---py---    (ei)+olda
Vaii1s-an---    olin
Vaii2s-an---    olid
Vaii3s-an---    oli
Vaii1p-an---    olime

Vaii2p-an---    olite
Vaii3p-an---    olid
Vaii1s-ay---    polnud
Vaii2s-ay---    polnud
Vaii3s-ay---    polnud
Vaii1p-ay---    polnud
Vaii2p-ay---    polnud
Vaii3p-ay---    polnud
Vaii---pn---    oldi
Vamp2s-an---    ole
Vamp3s-an---    olgu
Vamp1p-an---    olgem
Vamp2p-an---    olge
Vamp3p-an---    olgu
Vamp---pn---    oldagu
Vamp2s-ay---    &auml;ra
Vamp3s-ay---    &auml;rgu
Vamp1p-ay---    &auml;rgem
Vamp2p-ay---    &auml;rge
Vamp3p-ay---    &auml;rgu
Vamp---py---    &auml;rgu
Vac-----y---    ei
Vacp1s-an---    oleksin
Vacp2s-an---    oleksid
Vacp1o-an---    oleksime
Vacp2p-an---    oleksite
Vacp3p-an---    oleksid
Vacp1s-an---    oleks
Vacp2s-an---    oleks
Vacp3s-an---    oleks
Vacp1p-an---    oleks
Vacp2p-an---    oleks
Vacp3p-an---    oleks
Vacp1s-ay---    poleks
Vacp2s-ay---    poleks
Vacp3s-ay---    poleks
Vacp1p-ay---    poleks
Vacp2p-ay---    poleks
Vacp3p-ay---    poleks
Vacp---pn---    oldaks
Vacs1s-an---    olnuksin
Vacs2s-an---    olnuksid
Vacs1p-an---    olnuksime
Vacs2p-an---    olnuksite
Vacs3p-an---    olnuksid
Vacs1s-an---    olnuks
Vacs2s-an---    olnuks
Vacs3s-an---    olnuks
Vacs1p-an---    olnuks
Vacs2p-an---    olnuks
Vacs3p-an---    olnuks

Vacs1s-ay---    polnuks
Vacs2s-ay---    polnuks
Vacs3s-ay---    polnuks
Vacs1p-ay---    polnuks
Vacs2p-ay---    polnuks
Vacs3p-ay---    polnuks
Vacs---pn---    olduks
Van-----n---    olla
Vapp---an---    olev
Vapp---pn---    oldav
Vaps---an---    olnud
Vaps---pn---    oldud
Vag-----n---    olles
Vas----an--x    olema
Vas----an--2    olemas
Vas----an--e    olemast
Vas----an--4    olemaks
Vas----an--5    olemata
Vas----pn---    oldama
Vaq-----y---    ei
Vaqp1s-an---    olevat
Vaqp2s-an---    olevat
Vaqp3s-an---    olevat
Vaqp1p-an---    olevat
Vaqp2p-an---    olevat
Vaqp3p-an---    olevat
Vaqp1s-ay---    polevat
Vaqp2s-ay---    polevat
Vaqp3s-ay---    polevat
Vaqp1p-ay---    polevat
Vaqp2p-ay---    polevat
Vaqp3p-ay---    polevat
Vaqp---pn---    oldavat
Vaqs1s-an---    olnuvat
Vaqs2s-an---    olnuvat
Vaqs3s-an---    olnuvat
Vaqs1p-an---    olnuvat
Vaqs2p-an---    olnuvat
Vaqs3p-an---    olnuvat
Vaqs---pn---    olduvat
Voip1s-an---    saan
Voip2s-an---    saad
Voip3s-an---    saab
Voip1p-an---    saame
Voip2p-an---    saate
Voip3p-an---    saavad
Voip---pn---    saadakse

Voip---py---    (ei)+saada
Voii1s-an---    sain
Voii2s-an---    said
Voii3s-an---    sai
Voii1p-an---    saime
Voii2p-an---    saite
Voii3p-an---    said
Voii---pn---    saadi
Vomp2s-an---    saa
Vomp3s-an---    saagu
Vomp1p-an---    saagem
Vomp2p-an---    saage
Vomp3p-an---    saagu
Vomp---pn---    saadagu
Vocp1s-an---    saaksin
Vocp2s-an---    saaksid
Vocp1o-an---    saaksime
Vocp2p-an---    saaksite
Vocp3p-an---    saaksid
Vocp1s-an---    saaks
Vocp2s-an---    saaks
Vocp3s-an---    saaks
Vocp1p-an---    saaks
Vocp2p-an---    saaks
Vocp3p-an---    saaks
Vocp---pn---    saadaks
Vocs1s-an---    saanuksin
Vocs2s-an---    saanuksid
Vocs1p-an---    saanuksime
Vocs2p-an---    saanuksite
Vocs3p-an---    saanuksid
Vocs1s-an---    saanuks
Vocs2s-an---    saanuks
Vocs3s-an---    saanuks
Vocs1p-an---    saanuks
Vocs2p-an---    saanuks
Vocs3p-an---    saanuks
Vocs---pn---    saaduks
Von-----n---    saada
Vopp---an---    saav
Vopp---pn---    saadav
Vops---an---    saanud
Vops---pn---    saadud
Vog-----n---    saades
Vos----an--x    saama
Vos----an--2    saamas
Vos----an--e    saamast

Vos----an--4    saamaks
Vos----an--5    saamata
Vos----pn---    saadama
Voqp1s-an---    saavat
Voqp2s-an---    saavat
Voqp3s-an---    saavat
Voqp1p-an---    saavat
Voqp2p-an---    saavat
Voqp3p-an---    saavat
Voqp---pn---    saadavat
Voqs1s-an---    saanuvat
Voqs2s-an---    saanuvat
Voqs3s-an---    saanuvat
Voqs1p-an---    saanuvat
Voqs2p-an---    saanuvat
Voqs3p-an---    saanuvat
Voqs---pn---    saaduvat

3. Adjectives (A)
3.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type                          -
- -------------- -------------- - ------------
2 Degree         positive       p oluline
                 comparative    c olulisem
                 superlative    s olulisim
- -------------- -------------- - ------------
3 Gender                        -
- -------------- -------------- - ------------
4 Number         singular       s oluline
                 plural         p olulised
- -------------- -------------- - ------------
5 Case           nominative     n oluline
                 genitive       g olulise
           l.s.  partitive      1 olulist
                 illative       x olulisesse
                 inessive       2 olulises
                 elative        e olulisest
                 allative       t olulisele
                 adessive       3 olulisel
                 ablative       b oluliselt
           l.s.  translative    4 oluliseks
                 terminative    9 oluliseni
                 essive         w olulisena
           l.s.  abessive       5 oluliseta
           l.s.  komitative     k olulisega
           l.s.  aditive        7 olulisse

* ***************************** * ************
6 Definiteness                  -
- -------------- -------------- - ------------
7 Clitic                        -
- -------------- -------------- - ------------
8 Animate                       -
- -------------- -------------- - ------------
9 Formation                     -
- -------------- -------------- - ------------
10Owner_Number                  -
- -------------- -------------- - ------------
11Owner_Person                  -
---------------- -------------- - ------------
12Owned_Number                  -
================================= ============

3.2. Definitions
Lexicon    Definitions
----------------------
A-p---     Adjective positive
A-p-sn     Adjective positive singular nominative
A-p-sg     Adjective positive singular genitive
A-p-s1     Adjective positive singular partitive
A-p-sx     Adjective positive singular illative
A-p-s2     Adjective positive singular inessive
A-p-se     Adjective positive singular elative
A-p-st     Adjective positive singular allative
A-p-s3     Adjective positive singular adessive
A-p-sb     Adjective positive singular ablative
A-p-s4     Adjective positive singular translative
A-p-s9     Adjective positive singular terminative
A-p-sw     Adjective positive singular essive
A-p-s5     Adjective positive singular abessive
A-p-sk     Adjective positive singular komitative
A-p-s7     Adjective positive singular aditive
A-p-pn     Adjective positive plural nominative
A-p-pg     Adjective positive plural genitive
A-p-p1     Adjective positive plural partitive
A-p-px     Adjective positive plural illative
A-p-p2     Adjective positive plural inessive
A-p-pe     Adjective positive plural elative
A-p-pt     Adjective positive plural allative
A-p-p3     Adjective positive plural adessive
A-p-pb     Adjective positive plural ablative
A-p-p4     Adjective positive plural translative
A-p-p9     Adjective positive plural terminative
A-p-pw     Adjective positive plural essive
A-p-p5     Adjective positive plural abessive
A-p-pk     Adjective positive plural komitative
A-c-sn     Adjective comparative singular nominative

A-c-sg     Adjective comparative singular genitive
A-c-s1     Adjective comparative singular partitive
A-c-sx     Adjective comparative singular illative
A-c-s2     Adjective comparative singular inessive
A-c-se     Adjective comparative singular elative
A-c-st     Adjective comparative singular allative
A-c-s3     Adjective comparative singular adessive
A-c-sb     Adjective comparative singular ablative
A-c-s4     Adjective comparative singular translative
A-c-s9     Adjective comparative singular terminative
A-c-sw     Adjective comparative singular essive
A-c-s5     Adjective comparative singular abessive
A-c-sk     Adjective comparative singular komitative
A-c-pn     Adjective comparative plural nominative
A-c-pg     Adjective comparative plural genitive
A-c-p1     Adjective comparative plural partitive
A-c-px     Adjective comparative plural illative
A-c-p2     Adjective comparative plural inessive
A-c-pe     Adjective comparative plural elative
A-c-pt     Adjective comparative plural allative
A-c-p3     Adjective comparative plural adessive
A-c-pb     Adjective comparative plural ablative
A-c-p4     Adjective comparative plural translative
A-c-p9     Adjective comparative plural terminative
A-c-pw     Adjective comparative plural essive
A-c-p5     Adjective comparative plural abessive
A-c-pk     Adjective comparative plural komitative
A-s-sn     Adjective superlative singular nominative
A-s-sg     Adjective superlative singular genitive
A-s-s1     Adjective superlative singular partitive
A-s-sx     Adjective superlative singular illative
A-s-s2     Adjective superlative singular inessive
A-s-se     Adjective superlative singular elative
A-s-st     Adjective superlative singular allative
A-s-s3     Adjective superlative singular adessive
A-s-sb     Adjective superlative singular ablative
A-s-s4     Adjective superlative singular translative
A-s-s9     Adjective superlative singular terminative
A-s-sw     Adjective superlative singular essive
A-s-s5     Adjective superlative singular abessive
A-s-sk     Adjective superlative singular komitative
A-s-pn     Adjective superlative plural nominative
A-s-pg     Adjective superlative plural genitive
A-s-p1     Adjective superlative plural partitive
A-s-px     Adjective superlative plural illative
A-s-p2     Adjective superlative plural inessive
A-s-pe     Adjective superlative plural elative
A-s-pt     Adjective superlative plural allative
A-s-p3     Adjective superlative plural adessive
A-s-pb     Adjective superlative plural ablative
A-s-p4     Adjective superlative plural translative
A-s-p9     Adjective superlative plural terminative

A-s-pw     Adjective superlative plural essive
A-s-p5     Adjective superlative plural abessive
A-s-pk     Adjective superlative plural komitative

3.3. Examples
Lexicon   Examples
-----------------------
A-p---    ammuli
A-p-sn    oluline
A-p-sg    olulise
A-p-s1    olulist
A-p-sx    olulisesse
A-p-s2    olulises
A-p-se    olulisest
A-p-st    olulisele
A-p-s3    olulisel
A-p-sb    oluliselt
A-p-s4    oluliseks
A-p-s9    oluliseni
A-p-sw    olulisena
A-p-s5    oluliseta
A-p-sk    olulisega
A-p-s7    olulisse
A-p-pn    olulised
A-p-pg    oluliste
A-p-p1    olulisi
A-p-px    olulistesse
A-p-p2    olulistes
A-p-pe    olulistest
A-p-pt    olulistele
A-p-p3    olulistel
A-p-pb    olulistelt
A-p-p4    olulisteks
A-p-p9    olulisteni
A-p-pw    olulistena
A-p-p5    olulisteta
A-p-pk    olulistega
A-p-px    olulisisse
A-p-p2    olulisis
A-p-pe    olulisist
A-p-pt    olulisile
A-p-p3    olulisil
A-p-pb    olulisilt
A-p-p4    olulisiks
A-p-p9    olulisini
A-p-pw    olulisina
A-p-p5    olulisita
A-p-pk    olulisiga
A-c-sn    olulisem
A-c-sg    olulisema

A-c-s1    olulisemat
A-c-sx    olulisemasse
A-c-s2    olulisemas
A-c-se    olulisemast
A-c-st    olulisemale
A-c-s3    olulisemal
A-c-sb    olulisemalt
A-c-s4    olulisemaks
A-c-s9    olulisemani
A-c-sw    olulisemana
A-c-s5    olulisemata
A-c-sk    olulisemaga
A-c-pn    olulisemad
A-c-pg    olulisemate
A-c-p1    olulisemaid
A-c-px    olulisematesse
A-c-p2    olulisemates
A-c-pe    olulisematest
A-c-pt    olulisematele
A-c-p3    olulisematel
A-c-pb    olulisematelt
A-c-p4    olulisemateks
A-c-p9    olulisemateni
A-c-pw    olulisematena
A-c-p5    olulisemateta
A-c-pk    olulisematega
A-c-px    olulisemaisse
A-c-p2    olulisemais
A-c-pe    olulisemaist
A-c-pt    olulisemaile
A-c-p3    olulisemail
A-c-pb    olulisemailt
A-c-p4    olulisemaiks
A-c-p9    olulisemaini
A-c-pw    olulisemaina
A-c-p5    olulisemaita
A-c-pk    olulisemaiga
A-s-sn    olulisim
A-s-sg    olulisima
A-s-s1    olulisimat
A-s-sx    olulisimasse
A-s-s2    olulisimas
A-s-se    olulisimast
A-s-st    olulisimale
A-s-s3    olulisimal
A-s-sb    olulisimalt
A-s-s4    olulisimaks
A-s-s9    olulisimani
A-s-sw    olulisimana
A-s-s5    olulisimata
A-s-sk    olulisimaga
A-s-pn    olulisimad

A-s-pg    olulisimate
A-s-p1    olulisimaid
A-s-px    olulisimatesse
A-s-p2    olulisimates
A-s-pe    olulisimatest
A-s-pt    olulisimatele
A-s-p3    olulisimatel
A-s-pb    olulisimatelt
A-s-p4    olulisimateks
A-s-p9    olulisimateni
A-s-pw    olulisimatena
A-s-p5    olulisimateta
A-s-pk    olulisimatega
A-s-px    olulisimaisse
A-s-p2    olulisimais
A-s-pe    olulisimaist
A-s-pt    olulisimaile
A-s-p3    olulisimail
A-s-pb    olulisimailt
A-s-p4    olulisimaiks
A-s-p9    olulisimaini
A-s-pw    olulisimaina
A-s-p5    olulisimaita
A-s-pk    olulisimaiga

4. Pronouns (P)
4.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           personal       p mina
                 demonstrative  d selline
                 indefinite     i mingi
                 possessive     s iseenese
                 interrogative  q milline
                 relative       r milline
                 reflexive      x iseenda
                 reciprocal     y teineteise
           l.s.  determinal     m ise
- -------------- -------------- - ------------
2 Person         first          1 mina
                 second         2 sina
                 third          3 tema
- -------------- -------------- - ------------
3 Gender                        -
- -------------- -------------- - ------------
4 Number         singular       s mina
                 plural         p meie
- -------------- -------------- - ------------

5 Case           nominative     n mina
                 genitive       g minu
           l.s.  partitive      1 mind
                 illative       x minusse
                 inessive       2 minus
                 elative        e minust
                 allative       t minule
                 adessive       3 minul
                 ablative       b minult
           l.s.  translative    4 minuks
                 terminative    9 minuni
                 essive         w minuna
           l.s.  abessive       5 minuta
           l.s.  komitative     k minuga
           l.s.  aditive        7 sellisse
- -------------- -------------- - ------------
6 Owner_Number                  -
- -------------- -------------- - ------------
7 Owner_Gender                  -
********************************* ************
8 Clitic                        -
- -------------- -------------- - ------------
9 Referent_Type                 -
- -------------- -------------- - ------------
10Syntactic_Type                -
- -------------- -------------- - ------------
11Definiteness                  -
- -------------- -------------- - ------------
12Animate                       -
- -------------- -------------- - ------------
13Clitic_s                      -
- -------------- -------------- - ------------
14Pronoun_Form                  -
- -------------- -------------- - ------------
15Owner_Person                  -
- -------------- -------------- - ------------
16Owned_Number                  -
================================= ============

4.2. Definitions
Lexicon    Definitions
------------------------------------
Pp1-sn     Pronoun personal first singular nominative
Pp1-sg     Pronoun personal first singular genitive
Pp1-s1     Pronoun personal first singular partitive
Pp1-sx     Pronoun personal first singular illative
Pp1-s2     Pronoun personal first singular inessive
Pp1-se     Pronoun personal first singular elative
Pp1-st     Pronoun personal first singular allative
Pp1-s3     Pronoun personal first singular adessive
Pp1-sb     Pronoun personal first singular ablative

Pp1-s4     Pronoun personal first singular translative
Pp1-s9     Pronoun personal first singular terminative
Pp1-sw     Pronoun personal first singular essive
Pp1-s5     Pronoun personal first singular abessive
Pp1-sk     Pronoun personal first singular komitative
Pp1-pn     Pronoun personal first plural nominative
Pp1-pg     Pronoun personal first plural genitive
Pp1-p1     Pronoun personal first plural partitive
Pp1-px     Pronoun personal first plural illative
Pp1-p2     Pronoun personal first plural inessive
Pp1-pe     Pronoun personal first plural elative
Pp1-pt     Pronoun personal first plural allative
Pp1-p3     Pronoun personal first plural adessive
Pp1-pb     Pronoun personal first plural ablative
Pp1-p4     Pronoun personal first plural translative
Pp1-p9     Pronoun personal first plural terminative
Pp1-pw     Pronoun personal first plural essive
Pp1-p5     Pronoun personal first plural abessive
Pp1-pk     Pronoun personal first plural komitative
Pp2-sn     Pronoun personal second singular nominative
Pp2-sg     Pronoun personal second singular genitive
Pp2-s1     Pronoun personal second singular partitive
Pp2-sx     Pronoun personal second singular illative
Pp2-s2     Pronoun personal second singular inessive
Pp2-se     Pronoun personal second singular elative
Pp2-st     Pronoun personal second singular allative
Pp2-s3     Pronoun personal second singular adessive
Pp2-sb     Pronoun personal second singular ablative
Pp2-s4     Pronoun personal second singular translative
Pp2-s9     Pronoun personal second singular terminative
Pp2-sw     Pronoun personal second singular essive
Pp2-s5     Pronoun personal second singular abessive
Pp2-sk     Pronoun personal second singular komitative
Pp2-pn     Pronoun personal second plural nominative
Pp2-pg     Pronoun personal second plural genitive
Pp2-p1     Pronoun personal second plural partitive
Pp2-px     Pronoun personal second plural illative
Pp2-p2     Pronoun personal second plural inessive
Pp2-pe     Pronoun personal second plural elative
Pp2-pt     Pronoun personal second plural allative
Pp2-p3     Pronoun personal second plural adessive
Pp2-pb     Pronoun personal second plural ablative
Pp2-p4     Pronoun personal second plural translative
Pp2-p9     Pronoun personal second plural terminative
Pp2-pw     Pronoun personal second plural essive
Pp2-p5     Pronoun personal second plural abessive
Pp2-pk     Pronoun personal second plural komitative
Pp3-sn     Pronoun personal third singular nominative
Pp3-sg     Pronoun personal third singular genitive
Pp3-s1     Pronoun personal third singular partitive
Pp3-sx     Pronoun personal third singular illative

Pp3-s2     Pronoun personal third singular inessive
Pp3-se     Pronoun personal third singular elative
Pp3-st     Pronoun personal third singular allative
Pp3-s3     Pronoun personal third singular adessive
Pp3-sb     Pronoun personal third singular ablative
Pp3-s4     Pronoun personal third singular translative
Pp3-s9     Pronoun personal third singular terminative
Pp3-sw     Pronoun personal third singular essive
Pp3-s5     Pronoun personal third singular abessive
Pp3-sk     Pronoun personal third singular komitative
Pp3-pn     Pronoun personal third plural nominative
Pp3-pg     Pronoun personal third plural genitive
Pp3-p1     Pronoun personal third plural partitive
Pp3-px     Pronoun personal third plural illative
Pp3-p2     Pronoun personal third plural inessive
Pp3-pe     Pronoun personal third plural elative
Pp3-pt     Pronoun personal third plural allative
Pp3-p3     Pronoun personal third plural adessive
Pp3-pb     Pronoun personal third plural ablative
Pp3-p4     Pronoun personal third plural translative
Pp3-p9     Pronoun personal third plural terminative
Pp3-pw     Pronoun personal third plural essive
Pp3-p5     Pronoun personal third plural abessive
Pp3-pk     Pronoun personal third plural komitative
Pd--sn     Pronoun demonstrative singular nominative
Pd--sg     Pronoun demonstrative singular genitive
Pd--s1     Pronoun demonstrative singular partitive
Pd--sx     Pronoun demonstrative singular illative
Pd--s2     Pronoun demonstrative singular inessive
Pd--se     Pronoun demonstrative singular elative
Pd--st     Pronoun demonstrative singular allative
Pd--s3     Pronoun demonstrative singular adessive
Pd--sb     Pronoun demonstrative singular ablative
Pd--s4     Pronoun demonstrative singular translative
Pd--s9     Pronoun demonstrative singular terminative
Pd--sw     Pronoun demonstrative singular essive
Pd--s5     Pronoun demonstrative singular abessive
Pd--sk     Pronoun demonstrative singular komitative
Pd--s7     Pronoun demonstrative singular aditive
Pd--pn     Pronoun demonstrative plural nominative
Pd--pg     Pronoun demonstrative plural genitive
Pd--p1     Pronoun demonstrative plural partitive
Pd--px     Pronoun demonstrative plural illative
Pd--p2     Pronoun demonstrative plural inessive
Pd--pe     Pronoun demonstrative plural elative
Pd--pt     Pronoun demonstrative plural allative
Pd--p3     Pronoun demonstrative plural adessive
Pd--pb     Pronoun demonstrative plural ablative
Pd--p4     Pronoun demonstrative plural translative
Pd--p9     Pronoun demonstrative plural terminative
Pd--pw     Pronoun demonstrative plural essive
Pd--p5     Pronoun demonstrative plural abessive

Pd--pk     Pronoun demonstrative plural komitative
Pi----     Pronoun indefinite
Pi--sn     Pronoun indefinite singular nominative
Pi--sg     Pronoun indefinite singular genitive
Pi--s1     Pronoun indefinite singular partitive
Pi--sx     Pronoun indefinite singular illative
Pi--s2     Pronoun indefinite singular inessive
Pi--se     Pronoun indefinite singular elative
Pi--st     Pronoun indefinite singular allative
Pi--s3     Pronoun indefinite singular adessive
Pi--sb     Pronoun indefinite singular ablative
Pi--s4     Pronoun indefinite singular translative
Pi--s9     Pronoun indefinite singular terminative
Pi--sw     Pronoun indefinite singular essive
Pi--s5     Pronoun indefinite singular abessive
Pi--sk     Pronoun indefinite singular komitative
Pi--pn     Pronoun indefinite plural nominative
Pi--pg     Pronoun indefinite plural genitive
Pi--p1     Pronoun indefinite plural partitive
Pi--px     Pronoun indefinite plural illative
Pi--p2     Pronoun indefinite plural inessive
Pi--pe     Pronoun indefinite plural elative
Pi--pt     Pronoun indefinite plural allative
Pi--p3     Pronoun indefinite plural adessive
Pi--pb     Pronoun indefinite plural ablative
Pi--p4     Pronoun indefinite plural translative
Pi--p9     Pronoun indefinite plural terminative
Pi--pw     Pronoun indefinite plural essive
Pi--p5     Pronoun indefinite plural abessive
Pi--pk     Pronoun indefinite plural komitative
Ps----     Pronoun possessive
Ps--sn     Pronoun possessive singular nominative
Ps--sg     Pronoun possessive singular genitive
Ps--s1     Pronoun possessive singular partitive
Ps--sx     Pronoun possessive singular illative
Ps--s2     Pronoun possessive singular inessive
Ps--se     Pronoun possessive singular elative
Ps--st     Pronoun possessive singular allative
Ps--s3     Pronoun possessive singular adessive
Ps--sb     Pronoun possessive singular ablative
Ps--s4     Pronoun possessive singular translative
Ps--s9     Pronoun possessive singular terminative
Ps--sw     Pronoun possessive singular essive
Ps--s5     Pronoun possessive singular abessive
Ps--sk     Pronoun possessive singular komitative
Ps--pn     Pronoun possessive plural nominative
Ps--pg     Pronoun possessive plural genitive
Ps--p1     Pronoun possessive plural partitive
Ps--px     Pronoun possessive plural illative
Ps--p2     Pronoun possessive plural inessive
Ps--pe     Pronoun possessive plural elative

Ps--pt     Pronoun possessive plural allative
Ps--p3     Pronoun possessive plural adessive
Ps--pb     Pronoun possessive plural ablative
Ps--p4     Pronoun possessive plural translative
Ps--p9     Pronoun possessive plural terminative
Ps--pw     Pronoun possessive plural essive
Ps--p5     Pronoun possessive plural abessive
Ps--pk     Pronoun possessive plural komitative
Pq--sn     Pronoun interrogative singular nominative
Pq--sg     Pronoun interrogative singular genitive
Pq--s1     Pronoun interrogative singular partitive
Pq--sx     Pronoun interrogative singular illative
Pq--s2     Pronoun interrogative singular inessive
Pq--se     Pronoun interrogative singular elative
Pq--st     Pronoun interrogative singular allative
Pq--s3     Pronoun interrogative singular adessive
Pq--sb     Pronoun interrogative singular ablative
Pq--s4     Pronoun interrogative singular translative
Pq--s9     Pronoun interrogative singular terminative
Pq--sw     Pronoun interrogative singular essive
Pq--s5     Pronoun interrogative singular abessive
Pq--sk     Pronoun interrogative singular komitative
Pq--pn     Pronoun interrogative plural nominative
Pq--pg     Pronoun interrogative plural genitive
Pq--p1     Pronoun interrogative plural partitive
Pq--px     Pronoun interrogative plural illative
Pq--p2     Pronoun interrogative plural inessive
Pq--pe     Pronoun interrogative plural elative
Pq--pt     Pronoun interrogative plural allative
Pq--p3     Pronoun interrogative plural adessive
Pq--pb     Pronoun interrogative plural ablative
Pq--p4     Pronoun interrogative plural translative
Pq--p9     Pronoun interrogative plural terminative
Pq--pw     Pronoun interrogative plural essive
Pq--p5     Pronoun interrogative plural abessive
Pq--pk     Pronoun interrogative plural komitative
Pr--sn     Pronoun relative singular nominative
Pr--sg     Pronoun relative singular genitive
Pr--s1     Pronoun relative singular partitive
Pr--sx     Pronoun relative singular illative
Pr--s2     Pronoun relative singular inessive
Pr--se     Pronoun relative singular elative
Pr--st     Pronoun relative singular allative
Pr--s3     Pronoun relative singular adessive
Pr--sb     Pronoun relative singular ablative
Pr--s4     Pronoun relative singular translative
Pr--s9     Pronoun relative singular terminative
Pr--sw     Pronoun relative singular essive
Pr--s5     Pronoun relative singular abessive
Pr--sk     Pronoun relative singular komitative
Pr--pn     Pronoun relative plural nominative

Pr--pg     Pronoun relative plural genitive
Pr--p1     Pronoun relative plural partitive
Pr--px     Pronoun relative plural illative
Pr--p2     Pronoun relative plural inessive
Pr--pe     Pronoun relative plural elative
Pr--pt     Pronoun relative plural allative
Pr--p3     Pronoun relative plural adessive
Pr--pb     Pronoun relative plural ablative
Pr--p4     Pronoun relative plural translative
Pr--p9     Pronoun relative plural terminative
Pr--pw     Pronoun relative plural essive
Pr--p5     Pronoun relative plural abessive
Pr--pk     Pronoun relative plural komitative
Px--sg     Pronoun reflexive singular genitive
Px--s1     Pronoun reflexive singular partitive
Px--sx     Pronoun reflexive singular illative
Px--s2     Pronoun reflexive singular inessive
Px--se     Pronoun reflexive singular elative
Px--st     Pronoun reflexive singular allative
Px--s3     Pronoun reflexive singular adessive
Px--sb     Pronoun reflexive singular ablative
Px--s4     Pronoun reflexive singular translative
Px--s9     Pronoun reflexive singular terminative
Px--sw     Pronoun reflexive singular essive
Px--s5     Pronoun reflexive singular abessive
Px--sk     Pronoun reflexive singular komitative
Px--pg     Pronoun reflexive plural genitive
Px--p1     Pronoun reflexive plural partitive
Px--px     Pronoun reflexive plural illative
Px--p2     Pronoun reflexive plural inessive
Px--pe     Pronoun reflexive plural elative
Px--pt     Pronoun reflexive plural allative
Px--p3     Pronoun reflexive plural adessive
Px--pb     Pronoun reflexive plural ablative
Px--p4     Pronoun reflexive plural translative
Px--p9     Pronoun reflexive plural terminative
Px--pw     Pronoun reflexive plural essive
Px--p5     Pronoun reflexive plural abessive
Px--pk     Pronoun reflexive plural komitative
Py--sg     Pronoun reciprocal singular genitive
Py--s1     Pronoun reciprocal singular partitive
Py--sx     Pronoun reciprocal singular illative
Py--s2     Pronoun reciprocal singular inessive
Py--se     Pronoun reciprocal singular elative
Py--st     Pronoun reciprocal singular allative
Py--s3     Pronoun reciprocal singular adessive
Py--sb     Pronoun reciprocal singular ablative
Py--s4     Pronoun reciprocal singular translative
Py--s9     Pronoun reciprocal singular terminative
Py--sw     Pronoun reciprocal singular essive
Py--s5     Pronoun reciprocal singular abessive

Py--sk     Pronoun reciprocal singular komitative
Py--pg     Pronoun reciprocal plural genitive
Py--p1     Pronoun reciprocal plural partitive
Py--px     Pronoun reciprocal plural illative
Py--p2     Pronoun reciprocal plural inessive
Py--pe     Pronoun reciprocal plural elative
Py--pt     Pronoun reciprocal plural allative
Py--p3     Pronoun reciprocal plural adessive
Py--pb     Pronoun reciprocal plural ablative
Py--p4     Pronoun reciprocal plural translative
Py--p9     Pronoun reciprocal plural terminative
Py--pw     Pronoun reciprocal plural essive
Py--p5     Pronoun reciprocal plural abessive
Py--pk     Pronoun reciprocal plural komitative
Pm--sn     Pronoun determinal singular nominative
Pm--sg     Pronoun determinal singular genitive
Pm--s1     Pronoun determinal singular partitive
Pm--sx     Pronoun determinal singular illative
Pm--s2     Pronoun determinal singular inessive
Pm--se     Pronoun determinal singular elative
Pm--st     Pronoun determinal singular allative
Pm--s3     Pronoun determinal singular adessive
Pm--sb     Pronoun determinal singular ablative
Pm--s4     Pronoun determinal singular translative
Pm--s9     Pronoun determinal singular terminative
Pm--sw     Pronoun determinal singular essive
Pm--s5     Pronoun determinal singular abessive
Pm--sk     Pronoun determinal singular komitative
Pm--pn     Pronoun determinal plural nominative
Pm--pg     Pronoun determinal plural genitive
Pm--p1     Pronoun determinal plural partitive
Pm--px     Pronoun determinal plural illative
Pm--p2     Pronoun determinal plural inessive
Pm--pe     Pronoun determinal plural elative
Pm--pt     Pronoun determinal plural allative
Pm--p3     Pronoun determinal plural adessive
Pm--pb     Pronoun determinal plural ablative
Pm--p4     Pronoun determinal plural translative
Pm--p9     Pronoun determinal plural terminative
Pm--pw     Pronoun determinal plural essive
Pm--p5     Pronoun determinal plural abessive
Pm--pk     Pronoun determinal plural komitative

4.3. Examples
Lexicon  Examples
------------------------------------
Pp1-sn   mina
Pp1-sg   minu
Pp1-s1   mind
Pp1-sx   minusse

Pp1-s2   minus
Pp1-se   minust
Pp1-st   minule
Pp1-s3   minul
Pp1-sb   minult
Pp1-s4   minuks
Pp1-s9   minuni
Pp1-sw   minuna
Pp1-s5   minuta
Pp1-sk   minuga
Pp1-pn   meie
Pp1-pg   meie
Pp1-p1   meid
Pp1-px   meisse
Pp1-p2   meis
Pp1-pe   meist
Pp1-pt   meile
Pp1-p3   meil
Pp1-pb   meilt
Pp1-p4   meieks
Pp1-p9   meieni
Pp1-pw   meiena
Pp1-p5   meieta
Pp1-pk   meiega
Pp2-sn   sina
Pp2-sg   sinu
Pp2-s1   sind
Pp2-sx   sinusse
Pp2-s2   sinus
Pp2-se   sinust
Pp2-st   sinule
Pp2-s3   sinul
Pp2-sb   sinult
Pp2-s4   sinuks
Pp2-s9   sinuni
Pp2-sw   sinuna
Pp2-s5   sinuta
Pp2-sk   sinuga
Pp2-pn   teie
Pp2-pg   teie
Pp2-p1   teid
Pp2-px   teisse
Pp2-p2   teis
Pp2-pe   teist
Pp2-pt   teile
Pp2-p3   teil
Pp2-pb   teilt
Pp2-p4   teieks
Pp2-p9   teieni
Pp2-pw   teiena
Pp2-p5   teieta
Pp2-pk   teiega

Pp3-sn   tema
Pp3-sg   tema
Pp3-s1   teda
Pp3-sx   temasse
Pp3-s2   temas
Pp3-se   temast
Pp3-st   temale
Pp3-s3   temal
Pp3-sb   temalt
Pp3-s4   temaks
Pp3-s9   temani
Pp3-sw   temana
Pp3-s5   temata
Pp3-sk   temaga
Pp3-pn   nemad
Pp3-pg   nende
Pp3-p1   neid
Pp3-px   nendesse
Pp3-p2   nendes
Pp3-pe   nendest
Pp3-pt   nendele
Pp3-p3   nendel
Pp3-pb   nendelt
Pp3-p4   nendeks
Pp3-p9   nendeni
Pp3-pw   nendena
Pp3-p5   nendeta
Pp3-pk   nendega
Pp3-px   neisse
Pp3-p2   neis
Pp3-pe   neist
Pp3-pt   neile
Pp3-p3   neil
Pp3-pb   neilt
Pp3-p4   neiks
Pd--sn   selline
Pd--sg   sellise
Pd--s1   sellist
Pd--sx   sellisesse
Pd--s2   sellises
Pd--se   sellisest
Pd--st   sellisele
Pd--s3   sellisel
Pd--sb   selliselt
Pd--s4   selliseks
Pd--s9   selliseni
Pd--sw   sellisena
Pd--s5   selliseta
Pd--sk   sellisega
Pd--s7   sellisse
Pd--pn   sellised

Pd--pg   selliste
Pd--p1   selliseid
Pd--px   sellistesse
Pd--p2   sellistes
Pd--pe   sellistest
Pd--pt   sellistele
Pd--p3   sellistel
Pd--pb   sellistelt
Pd--p4   sellisteks
Pd--p9   sellisteni
Pd--pw   sellistena
Pd--p5   sellisteta
Pd--pk   sellistega
Pd--px   selliseisse
Pd--p2   selliseis
Pd--pe   selliseist
Pd--pt   selliseile
Pd--p3   selliseil
Pd--pb   selliseilt
Pd--p4   selliseiks
Pd--p9   selliseini
Pd--pw   selliseina
Pd--p5   selliseita
Pd--pk   selliseiga
Pi----   muist
Pi--sn   mingi
Pi--sg   mingi
Pi--s1   mingit
Pi--sx   mingisse
Pi--s2   mingis
Pi--se   mingist
Pi--st   mingile
Pi--s3   mingil
Pi--sb   mingilt
Pi--s4   mingiks
Pi--s9   mingini
Pi--sw   mingina
Pi--s5   mingita
Pi--sk   mingiga
Pi--pn   mingid
Pi--pg   mingite
Pi--p1   mingeid
Pi--px   mingitesse
Pi--p2   mingites
Pi--pe   mingitest
Pi--pt   mingitele
Pi--p3   mingitel
Pi--pb   mingitelt
Pi--p4   mingiteks
Pi--p9   mingiteni
Pi--pw   mingitena
Pi--p5   mingiteta

Pi--pk   mingitega
Pi--px   mingeisse
Pi--p2   mingeis
Pi--pe   mingeist
Pi--pt   mingeile
Pi--p3   mingeil
Pi--pb   mingeilt
Pi--p4   mingeiks
Pi--p9   mingeini
Pi--pw   mingeina
Pi--p5   mingeita
Pi--pk   mingeiga
Ps----   oma
Ps--sg   iseenese
Ps--s1   iseennast
Ps--sx   iseenesesse
Ps--s2   iseeneses
Ps--se   iseenesest
Ps--st   iseenesele
Ps--s3   iseenesel
Ps--sb   iseeneselt
Ps--s4   iseeneseks
Ps--s9   iseeneseni
Ps--sw   iseenesena
Ps--s5   iseeneseta
Ps--sk   iseenesega
Ps--pg   iseeneste
Ps--p1   iseendid
Ps--px   iseenestesse
Ps--p2   iseenestes
Ps--pe   iseenestest
Ps--pt   iseenestele
Ps--p3   iseenestel
Ps--pb   iseenestelt
Ps--p4   iseenesteks
Ps--p9   iseenesteni
Ps--pw   iseenestena
Ps--p5   iseenesteta
Ps--pk   iseenestega
Pq--sn   milline
Pq--sg   millise
Pq--s1   millist
Pq--sx   millisesse
Pq--s2   millises
Pq--se   millisest
Pq--st   millisele
Pq--s3   millisel
Pq--sb   milliselt
Pq--s4   milliseks
Pq--s9   milliseni
Pq--sw   millisena

Pq--s5   milliseta
Pq--sk   millisega
Pq--pn   millised
Pq--pg   milliste
Pq--p1   milliseid
Pq--px   millistesse
Pq--p2   millistes
Pq--pe   millistest
Pq--pt   millistele
Pq--p3   millistel
Pq--pb   millistelt
Pq--p4   millisteks
Pq--p9   millisteni
Pq--pw   millistena
Pq--p5   millisteta
Pq--pk   millistega
Pq--px   milliseisse
Pq--p2   milliseis
Pq--pe   milliseist
Pq--pt   milliseile
Pq--p3   milliseil
Pq--pb   milliseilt
Pq--p4   milliseiks
Pq--p9   milliseini
Pq--pw   milliseina
Pq--p5   milliseita
Pq--pk   milliseiga
Pr--sn   milline
Pr--sg   millise
Pr--s1   millist
Pr--sx   millisesse
Pr--s2   millises
Pr--se   millisest
Pr--st   millisele
Pr--s3   millisel
Pr--sb   milliselt
Pr--s4   milliseks
Pr--s9   milliseni
Pr--sw   millisena
Pr--s5   milliseta
Pr--sk   millisega
Pr--pn   millised
Pr--pg   milliste
Pr--p1   milliseid
Pr--px   millistesse
Pr--p2   millistes
Pr--pe   millistest
Pr--pt   millistele
Pr--p3   millistel
Pr--pb   millistelt
Pr--p4   millisteks
Pr--p9   millisteni

Pr--pw   millistena
Pr--p5   millisteta
Pr--pk   millistega
Pr--px   milliseisse
Pr--p2   milliseis
Pr--pe   milliseist
Pr--pt   milliseile
Pr--p3   milliseil
Pr--pb   milliseilt
Pr--p4   milliseiks
Pr--p9   milliseini
Pr--pw   milliseina
Pr--p5   milliseita
Pr--pk   milliseiga
Px--sg   iseenda
Px--s1   iseend
Px--sx   iseendasse
Px--s2   iseendas
Px--se   iseendast
Px--st   iseendale
Px--s3   iseendal
Px--sb   iseendalt
Px--s4   iseendaks
Px--s9   iseendani
Px--sw   iseendana
Px--s5   iseendata
Px--sk   iseendaga
Px--pg   iseendi
Px--p1   iseendid
Px--px   iseendisse
Px--p2   iseendis
Px--pe   iseendist
Px--pt   iseendile
Px--p3   iseendil
Px--pb   iseendilt
Px--p4   iseendiks
Px--p9   iseendini
Px--pw   iseendina
Px--p5   iseendita
Px--pk   iseendiga
Py--sg   teineteise
Py--s1   teineteist
Py--sx   teineteisesse
Py--s2   teineteises
Py--se   teineteisest
Py--st   teineteisele
Py--s3   teineteisel
Py--sb   teineteiselt
Py--s4   teineteiseks
Py--s9   teineteiseni
Py--sw   teineteisena

Py--s5   teineteiseta
Py--sk   teineteisega
Py--pg   teineteiste
Py--p1   teineteisi
Py--px   teineteistesse
Py--p2   teineteistes
Py--pe   teineteistest
Py--pt   teineteistele
Py--p3   teineteistel
Py--pb   teineteistelt
Py--p4   teineteisteks
Py--p9   teineteisteni
Py--pw   teineteistena
Py--p5   teineteisteta
Py--pk   teineteistega
Pm--sn   ise
Pm--sg   enda
Pm--s1   end
Pm--sx   endasse
Pm--s2   endas
Pm--se   endast
Pm--st   endale
Pm--s3   endal
Pm--sb   endalt
Pm--s4   endaks
Pm--s9   endani
Pm--sw   endana
Pm--s5   endata
Pm--sk   endaga
Pm--pn   ise
Pm--pg   endi
Pm--p1   endid
Pm--px   endisse
Pm--p2   endis
Pm--pe   endist
Pm--pt   endile
Pm--p3   endil
Pm--pb   endilt
Pm--p4   endiks
Pm--p9   endini
Pm--pw   endina
Pm--p5   endita
Pm--pk   endiga

5. Determiners (D)
Not applicable.

6. Atricles (T)

Not applicable.

7. Adverbs (R)
7.1 Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type                          - asjatult
- -------------- -------------- - ------------
2 Degree                        -
**********************************************
3 Clitic                        -
- -------------- -------------- - ------------
4 Number                        -
- -------------- -------------- - ------------
5 Person                        -
================================= ============

7.2. Definitions
Lexicon  Definitions
--------------------
R        Adverb

7.3. Examples
Lexicon  Examples
-----------------
R        asjatult

8. Adpositions (S)
8.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           preposition    p enne
                 postposition   t kaudu
- -------------- -------------- - ------------
2 Formation                     -
**********************************************
3 Case                          -
- -------------- -------------- - ------------
4 Clitic                        -
================================= ============

8.2. Definitions
Lexicon  Definitions
-----------------------------------
Sp       Adposition preposition
St       Adposition postposition

8.3. Examples
Lexicon  Examples
-----------------------------------
Sp       enne
St       kaudu

9. Conjunctions (C)
9.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           coordinating   c ja
                 subordinating  s et
**********************************************
2 Formation                     -
- -------------- -------------- - ------------
3 Coord_Type                    -
- -------------- -------------- - ------------
4 Sub_Type                      -
- -------------- -------------- - ------------
5 Clitic                        -
- -------------- -------------- - ------------
6 Number                        -
- -------------- -------------- - ------------
7 Person                        -
================================= ============

9.2. Definitions
Lexicon  Definitions
-----------------------------------
Cc       Conjuction coordinating
Cs       Conjuction subordinating

9.3. Examples
Lexicon  Examples

-----------------------------------
Cc       ja
Cs       et

10. Numerals (M)
10.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type           cardinal       c kaheksa
                 ordinal        o kaheksas
- -------------- -------------- - ------------
2 Gender                        -
- -------------- -------------- - ------------
3 Number         singular       s kaheksa
                 plural         p kaheksad
- -------------- -------------- - ------------
4 Case           nominative     n kaheksa
                 genitive       g kaheksa
           l.s.  partitive      1 kaheksat
                 illative       x kaheksasse
                 inessive       2 kaheksas
                 elative        e kaheksast
                 allative       t kaheksale
                 adessive       3 kaheksal
                 ablative       b kaheksalt
           l.s.  translative    4 kaheksaks
                 terminative    9 kaheksani
                 essive         w kaheksana
           l.s.  abessive       5 kaheksata
           l.s.  komitative     k kaheksaga
           l.s.  aditive        7 viite
********************************* ------------
5 Form           digit          d 1984
                 roman          r IX
                 letter         l kaheksa
- -------------- -------------- - ------------
6 Definiteness                  -
- -------------- -------------- - ------------
7 Clitic                        -
- -------------- -------------- - ------------
8 Class                         -
- -------------- -------------- - ------------
9 Animate                       -
- -------------- -------------- - ------------
10Owner_Number                  -
- -------------- -------------- - ------------
11Owner_Person                  -
- -------------- -------------- - ------------
12Owned_Number                  -

================================= ============

10.2. Definitions
Lexicon    Definitions
------------------------------------
Mc-snl     Numeral cardinal singular nominative letter
Mc-sgl     Numeral cardinal singular genitive letter
Mc-s1l     Numeral cardinal singular partitive letter
Mc-sxl     Numeral cardinal singular illative letter
Mc-s2l     Numeral cardinal singular inessive letter
Mc-sel     Numeral cardinal singular elative letter
Mc-stl     Numeral cardinal singular allative letter
Mc-s3l     Numeral cardinal singular adessive letter
Mc-sbl     Numeral cardinal singular ablative letter
Mc-s4l     Numeral cardinal singular translative letter
Mc-s9l     Numeral cardinal singular terminative letter
Mc-swl     Numeral cardinal singular essive letter
Mc-s5l     Numeral cardinal singular abessive letter
Mc-skl     Numeral cardinal singular komitative letter
Mc-s7l     Numeral cardinal singular aditive letter
Mc-pnl     Numeral cardinal plural nominative letter
Mc-pgl     Numeral cardinal plural genitive letter
Mc-p1l     Numeral cardinal plural partitive letter
Mc-pxl     Numeral cardinal plural illative letter
Mc-p2l     Numeral cardinal plural inessive letter
Mc-pel     Numeral cardinal plural elative letter
Mc-ptl     Numeral cardinal plural allative letter
Mc-p3l     Numeral cardinal plural adessive letter
Mc-pbl     Numeral cardinal plural ablative letter
Mc-p4l     Numeral cardinal plural translative letter
Mc-p9l     Numeral cardinal plural terminative letter
Mc-pwl     Numeral cardinal plural essive letter
Mc-p5l     Numeral cardinal plural abessive letter
Mc-pkl     Numeral cardinal plural komitative letter
Mo-snl     Numeral ordinal singular nominative letter
Mo-sgl     Numeral ordinal singular genitive letter
Mo-s1l     Numeral ordinal singular partitive letter
Mo-sxl     Numeral ordinal singular illative letter
Mo-s2l     Numeral ordinal singular inessive letter
Mo-sel     Numeral ordinal singular elative letter
Mo-stl     Numeral ordinal singular allative letter
Mo-s3l     Numeral ordinal singular adessive letter
Mo-sbl     Numeral ordinal singular ablative letter
Mo-s4l     Numeral ordinal singular translative letter
Mo-s9l     Numeral ordinal singular terminative letter
Mo-swl     Numeral ordinal singular essive letter
Mo-s5l     Numeral ordinal singular abessive letter
Mo-skl     Numeral ordinal singular komitative letter
Mo-pnl     Numeral ordinal plural nominative letter
Mo-pgl     Numeral ordinal plural genitive letter

Mo-p1l     Numeral ordinal plural partitive letter
Mo-pxl     Numeral ordinal plural illative letter
Mo-p2l     Numeral ordinal plural inessive letter
Mo-pel     Numeral ordinal plural elative letter
Mo-ptl     Numeral ordinal plural allative letter
Mo-p3l     Numeral ordinal plural adessive letter
Mo-pbl     Numeral ordinal plural ablative letter
Mo-p4l     Numeral ordinal plural translative letter
Mo-p9l     Numeral ordinal plural terminative letter
Mo-pwl     Numeral ordinal plural essive letter
Mo-p5l     Numeral ordinal plural abessive letter
Mo-pkl     Numeral ordinal plural komitative letter
M----d     Numeral digit
M----r     Numeral roman

10.3. Examples
Lexicon   Examples
-----------------------
Mc-snl    kaheksa
Mc-sgl    kaheksa
Mc-s1l    kaheksat
Mc-sxl    kaheksasse
Mc-s2l    kaheksas
Mc-sel    kaheksast
Mc-stl    kaheksale
Mc-s3l    kaheksal
Mc-sbl    kaheksalt
Mc-s4l    kaheksaks
Mc-s9l    kaheksani
Mc-swl    kaheksana
Mc-s5l    kaheksata
Mc-skl    kaheksaga
Mc-s7l    viite
Mc-pnl    kaheksad
Mc-pgl    kaheksate
Mc-p1l    kaheksaid
Mc-pxl    kaheksatesse
Mc-p2l    kaheksates
Mc-pel    kaheksatest
Mc-ptl    kaheksatele
Mc-p3l    kaheksatel
Mc-pbl    kaheksatelt
Mc-p4l    kaheksateks
Mc-p9l    kaheksateni
Mc-pwl    kaheksatena
Mc-p5l    kaheksateta
Mc-pkl    kaheksatega
Mc-pxl    kaheksaisse
Mc-p2l    kaheksais
Mc-pel    kaheksaist
Mc-ptl    kaheksaile

Mc-p3l    kaheksail
Mc-pbl    kaheksailt
Mc-p4l    kaheksaiks
Mc-p9l    kaheksaini
Mc-pwl    kaheksaina
Mc-p5l    kaheksaita
Mc-pkl    kaheksaiga
Mo-snl    kaheksas
Mo-sgl    kaheksanda
Mo-s1l    kaheksandat
Mo-sxl    kaheksandasse
Mo-s2l    kaheksandas
Mo-sel    kaheksandast
Mo-stl    kaheksandale
Mo-s3l    kaheksandal
Mo-sbl    kaheksandalt
Mo-s4l    kaheksandaks
Mo-s9l    kaheksandani
Mo-swl    kaheksandana
Mo-s5l    kaheksandata
Mo-skl    kaheksandaga
Mo-pnl    kaheksandad
Mo-pgl    kaheksandate
Mo-p1l    kaheksandaid
Mo-pxl    kaheksandatesse
Mo-p2l    kaheksandates
Mo-pel    kaheksandatest
Mo-ptl    kaheksandatele
Mo-p3l    kaheksandatel
Mo-pbl    kaheksandatelt
Mo-p4l    kaheksandateks
Mo-p9l    kaheksandateni
Mo-pwl    kaheksandatena
Mo-p5l    kaheksandateta
Mo-pkl    kaheksandatega
Mo-pxl    kaheksandaisse
Mo-p2l    kaheksandais
Mo-pel    kaheksandaist
Mo-ptl    kaheksandaile
Mo-p3l    kaheksandail
Mo-pbl    kaheksandailt
Mo-p4l    kaheksandaiks
Mo-p9l    kaheksandaini
Mo-pwl    kaheksandaina
Mo-p5l    kaheksandaita
Mo-pkl    kaheksandaiga
M----d    1984
M----r    IX

11. Interjections (I)

11.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C EXAMPLE
= ============== ============== = ============
1 Type                          - oi
- -------------- -------------- - ------------
2 Formation                     -
================================= ============

11.2. Definitions
Lexicon  Definitions
---------------------
I        Interjection

11.3. Examples
Lexicon  Examples
-----------------
I        oi

12. Residual (X)
Not applicable.

13. Abreviations (Y)
13.1. Lexicon
= ============== ============== = ============
P ATT            VAL            C
= ============== ============== = ============
1 Syntactic_Type nominal        n CD
                 verbal         v vt
                 adjectival     a hisp
                 adverbial      r jne
- ------------------------------- ------------
2 Gender                        -
- -------------- -------------- - ------------
3 Number         singular       s CD
                 plural         p CDd
- -------------- -------------- - ------------
4 Case     l.s.  nominative     n CD
           l.s.  genitive       g CD
           l.s.  partitive      1 CDd
           l.s.  illative       x CDsse
           l.s.  inessive       2 CDs
           l.s.  elative        e CDst

           l.s.  allative       t CDle
           l.s.  adessive       3 CDl
           l.s.  ablative       b CDlt
           l.s.  translative    4 CDks
           l.s.  terminative    9 CDni
           l.s.  essive         w CDna
           l.s.  abessive       5 CDta
           l.s.  komitative     k CDga
           l.s.  aditive        7
- -------------- -------------- - ------------
5 Definiteness                  -
================================= ============

13.2. Definitions
Lexicon   Definitions
------------------------------------
Yn-sn     Abbreviation nominal singular nominative
Yn-sg     Abbreviation nominal singular genitive
Yn-s1     Abbreviation nominal singular partitive
Yn-sx     Abbreviation nominal singular illative
Yn-s2     Abbreviation nominal singular inessive
Yn-se     Abbreviation nominal singular elative
Yn-st     Abbreviation nominal singular allative
Yn-s3     Abbreviation nominal singular adessive
Yn-sb     Abbreviation nominal singular ablative
Yn-s4     Abbreviation nominal singular translative
Yn-s9     Abbreviation nominal singular terminative
Yn-sw     Abbreviation nominal singular essive
Yn-s5     Abbreviation nominal singular abessive
Yn-sk     Abbreviation nominal singular komitative
Yn-pn     Abbreviation nominal plural nominative
Yn-pg     Abbreviation nominal plural genitive
Yn-p1     Abbreviation nominal plural partitive
Yn-px     Abbreviation nominal plural illative
Yn-p2     Abbreviation nominal plural inessive
Yn-pe     Abbreviation nominal plural elative
Yn-pt     Abbreviation nominal plural allative
Yn-p3     Abbreviation nominal plural adessive
Yn-pb     Abbreviation nominal plural ablative
Yn-p4     Abbreviation nominal plural translative
Yn-p9     Abbreviation nominal plural terminative
Yn-pw     Abbreviation nominal plural essive
Yn-p5     Abbreviation nominal plural abessive
Yn-pk     Abbreviation nominal plural komitative
Yv---     Abbreviation verbal
Ya---     Abbreviation adjectival
Yr---     Abbreviation adverbial

13.3. Examples
Lexicon    Examples
-------------------
Yn-sn      CD
Yn-sg      CD
Yn-s1      CDd
Yn-sx      CDsse
Yn-s2      CDs
Yn-se      CDst
Yn-st      CDle
Yn-s3      CDl
Yn-sb      CDlt
Yn-s4      CDks
Yn-s9      CDni
Yn-sw      CDna
Yn-s5      CDta
Yn-sk      CDga
Yn-pn      CDd
Yn-pg      CDde
Yn-p1      CDsid
Yn-px      CDdesse
Yn-p2      CDdes
Yn-pe      CDdest
Yn-pt      CDdele
Yn-p3      CDdel
Yn-pb      CDdelt
Yn-p4      CDdeks
Yn-p9      CDdeni
Yn-pw      CDdena
Yn-p5      CDdeta
Yn-pk      CDdega
Yv---      vt
Ya---      hisp
Yr---      jne

14. Particles (Q)
Not applicable.

COP project 106 multext-East  Deliverable D1.1 F— Hungarian  110

3.4 Application to Hungarian

The application to Hungarian has been elaborated by Laszlo Tihanyi.

1. Noun (N)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           common         c
                 proper         p
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 instrumental   i
                 illative       x
                 inessive       2
                 elative        e
                 allative       t
                 adessive       3
                 ablative       b
                 terminative    9
                 essive         w
           l.s.  temporalis     m
           l.s.  causalis       c
           l.s.  sublative      s
           l.s.  delative       h
           l.s.  sociative      q
           l.s.  factive        y
           l.s.  superessive    p
           l.s.  distributive   u
* ***************************** *
5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -
- -------------- -------------- -
7 Animate        no             n
                 yes            y
- -------------- -------------- -
8 Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
9 Owner_Person   first          1
                 second         2
                 third          3

---------------- -------------- -
10Owned_Number   singular       s
                 plural         p
=================================

Notes:
1. Gender: Not applicable to Hungarian.
2. Numbers:
   Hungarian has three types of number in the nominal inflection:
   1. The number of the noun.
   2. The number of owners that own the noun.
   3. The number of the context given referent, which is some possession
      of the noun, i.e. belongs to the noun (anaphoric possessive).
3. Genitive is rarely marked in Hungarian. If marked then with the same
   suffix as that of dative case.
   Nouns with zero suffix can be nominative or genitive, so they are
   ambigious.
4. Common Hungarian nouns can have
   2(Number)*7(OwnerPerson,OwnerNumber)*3(OwnedNumber)*22(Case)=1324
   different forms.
5. The usage of -szerte[SZERTE] meaning 'many places around' is restricted
   so we do not take it as a regular case ending. The same stands for
   '-lag/-leg' marked with [LAG], e.g. 'ha'zilag' means (made) at home.

Combinations:
0   1    3    4                        7     8        9         10
--- ---- ---- ------------------------ ----  -------- --------- --------
Pos Type Num  Case                     Anim  OwnerNum OwnerPers OwnedNum
--- ---- ---- ------------------------ ----  -------- --------- --------
 N  [cp] [sp] [ngdaix2et3b9wmcshqypu] [yn]  [sp]     [123]     [sp]

Examples on combinations:
a) Combinations on Number, OwnerNumber, OwnerPerson and OwnedNumber:
01234567890
Nc-sn--n---  asztal[N]                          table
Nc-sn--n--s  asztal[N]+e'[POS]                  thing of the table
Nc-sn--n--p  asztal[N]+e'i[POSi]                things of the table
Nc-sn--ns1-  asztal[N]+om[PSe1]                 my table
Nc-sn--ns1s  asztal[N]+om[PSe1]+e'[POS]         thing of my table
Nc-sn--ns1p  asztal[N]+om[PSe1]+e'i[POSi]       things of my table
Nc-sn--ns2-  asztal[N]+od[PSe2]                 your table
Nc-sn--ns2s  asztal[N]+od[PSe2]+e'[POS]         thing of your table

Nc-sn--ns2p  asztal[N]+od[PSe2]+e'i[POSi]       things of your table
Nc-sn--ns3-  asztal[N]+a[PSe3]                  his/her table
Nc-sn--ns3s  asztal[N]+a'[PSe3]+e'[POS]         thing of his/her table
Nc-sn--ns3p  asztal[N]+a'[PSe3]+e'i[POSi]       things of his/her table
Nc-sn--np1-  asztal[N]+unk[PSt1]                our table
Nc-sn--np1s  asztal[N]+unk[PSt1]+e'[POS]        thing of our table
Nc-sn--np1p  asztal[N]+unk[PSt1]+e'i[POSi]      things of our table
Nc-sn--np2-  asztal[N]+otok[PSt2]               your table
Nc-sn--np2s  asztal[N]+otok[PSt2]+e'[POS]       thing of your table
Nc-sn--np2p  asztal[N]+otok[PSt2]+e'i[POSi]     things of your table
Nc-sn--np3-  asztal[N]+uk[PSt2]                 their table
Nc-sn--np3s  asztal[N]+uk[PSt2]+e'[POS]         thing of their table
Nc-sn--np3p  asztal[N]+uk[PSt2]+e'i[POSi]       things of their table
Nc-pn--n---  asztal[N]+ok[PL]                   tables
Nc-pn--n--s  asztal[N]+ok[PL]+e'[POS]           thing of the tables
Nc-pn--n--p  asztal[N]+ok[PL]+e'i[POSi]         things of the tables
Nc-pn--ns1-  asztal[N]+aim[PSe1i]               my tables
Nc-pn--ns1s  asztal[N]+aim[PSe1i]+e'[POS]       thing of my tables
Nc-pn--ns1p  asztal[N]+aim[PSe1i]+e'i[POSi]     things of my tables
Nc-pn--ns2-  asztal[N]+aid[PSe2i]               your tables
Nc-pn--ns2s  asztal[N]+aid[PSe2i]+e'[POS]       thing of your tables
Nc-pn--ns2p  asztal[N]+aid[PSe2i]+e'i[POSi]     things of your tables
Nc-pn--ns3-  asztal[N]+ai[PSe3i]                his/her tables
Nc-pn--ns3s  asztal[N]+ai[PSe3i]+e'[POS]        thing of his/her tables
Nc-pn--ns3p  asztal[N]+ai[PSe3i]+e'i[POSi]      things of his/her tables
Nc-pn--np1-  asztal[N]+aink[PSt1i]              our tables
Nc-pn--np1s  asztal[N]+aink[PSt1i]+e'[POS]      thing of our tables
Nc-pn--np1p  asztal[N]+aink[PSt1i]+e'i[POSi]    things of our tables
Nc-pn--np2-  asztal[N]+aitok[PSt2i]             your tables
Nc-pn--np2s  asztal[N]+aitok[PSt2i]+e'[POS]     thing of your tables
Nc-pn--np2p  asztal[N]+aitok[PSt2i]+e'i[POSi]   things of your tables
Nc-pn--np3-  asztal[N]+aik[PSt3i]               their tables
Nc-pn--np3s  asztal[N]+aik[PSt3i]+e'[POS]       thing of their tables
Nc-pn--np3p  asztal[N]+aik[PSt3i]+e'i[POSi]     things of their tables

b) Combinations on Case:
Category:   Case:          Example analysis:
01234567890
Nc-sn------ nominative     asztal[N]              The table is new.
Nc-sg------ genitive       asztal[N]+nak[GEN]     The leg of the table is red.
Nc-sd------ dative         asztal[N]+nak[DAT]     It gives a new look to the table.
Nc-sa------ accusative     asztal[N]+t[ACC]       I see the table.
Nc-si------ instrumental   asztal[N]+lal[INS]     Do something with the table.
Nc-sx------ illative       asztal[N]+ba[ILL]      She put it into the table.
Nc-s2------ inessive       asztal[N]+ban[INE]     He keeps it in this table.
Nc-se------ elative        asztal[N]+bo'l[ELA]    She took it out from this table.
Nc-st------ allative       asztal[N]+hoz[ALL]     Go to the table.
Nc-s3------ adessive       asztal[N]+na'l[ADE]    At the table there is a chair.
Nc-sb------ ablative       asztal[N]+to'l[ABL]    She came from the table.
Nc-s9------ terminative    asztal[N]+ig[TER]      Go as far as the table.
Nc-sw------ essive         asztal[N]+ul[ESS]      The plate served us as a table.

Nc-sm------ temporalis     hat[N]+kor[TEM]        We arrived at six.
Nc-sc------ causalis       asztal[N]+e'rt[CAU]    We do it for the table.
Nc-ss------ sublative      asztal[N]+ra[SUB]      Put it onto the table.
Nc-sh------ delative       asztal[N]+ro'l[DEL]    Talk about the table.
Nc-sq------ sociative      asztal[N]+ostul[SOC]   It is sold together with the table.
Nc-sy------ factive        asztal[N]+la'[FAC]     Soon it became a table.
Nc-sp------ superessive    asztal[N]+on[SUP]      The cat sits on the table.
Nc-su------ distributive   asztal[N]+onke'nt[DIS] We collected 10 forints in each table.
-------------------------------------------------------------------------------

2. Verb (V)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           main           m
                 auxiliary      a
- -------------- -------------- -
2 VForm          indicative     i
                 imperative     m
                 conditional    c
                 infinitive     n
- -------------- -------------- -
3 Tense          present        p
                 past           s
- -------------- -------------- -
4 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
5 Number         singular       s
                 plural         p
- -------------- -------------- -
6 Gender                        -
*********************************
7 Voice                         -
- -------------- -------------- -
8 Negative                      -
- -------------- -------------- -
9 Definiteness   no             n
                 yes            y
           l.s.  1s2s           2
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
=================================

Notes:
1. 1s2s is a special form for definitness, in which the speaker's person is
   first singular (I) and the target of the transitivity is second singular
   (you).
2. A Hungarian verb has
   3(Person)*2(Number)*4(Present indicative, Present imperativ,
   Present conditional,Past indicative)*2(Transitivity)
   +4(1s2s)+7(Infinitive)=59 different forms.

Combinations:
--- ---- ----- ----- ------ ------ -----------
Pos Type VForm Tense Person Number Definitness
--- ---- ----- ----- ------ ------ -----------
 V   m   i     [ps] [123]   [sp]   [yn]
 V   m   c     p    [123]   [sp]   [yn]
 V   m   m     p    [123]   [sp]   [yn]
 V   m   i     [ps] 1       s      2
 V   m   [cm]  p    1       s      2
 V   m   n     -    -       -      -
 V   m   n     p    [123]   [sp]   -
 V   a   i     p    [123]   [sp]   n
--- ---- ----- ----- ------ ------ -----------

Examples on combinations:
01234567890123
Vmip1s------n-   tanul[V]+ok[e1]                 I learn (a)
Vmip2s------n-   tanul[V]+sz[e2]                 you learn (a)
Vmip3s------n-   tanul[V]                        she learns (a)
Vmip1p------n-   tanul[V]+unk[t1]                we learn (a)
Vmip2p------n-   tanul[V]+tok[t2]                you learn  (a)
Vmip3p------n-   tanul[V]+nak[t3]                they learn (a)
Vmip1s------y-   tanul[V]+om[Te1]                I learn (the)
Vmip2s------y-   tanul[V]+od[Te2]                you learn (the)
Vmip3s------y-   tanul[V]+ja[Te3]                she learns (the)
Vmip1p------y-   tanul[V]+juk[Tt1]               we learn (the)
Vmip2p------y-   tanul[V]+ja'tok[Tt2]            you learn  (the)
Vmip3p------y-   tanul[V]+ja'k[Tt3]              they learn (the)
Vmis1s------n-   tanul[V]+tam[Me1]               I learnt (a)
Vmis2s------n-   tanul[V]+ta'l[Me2]              you learnt (a)
Vmis3s------n-   tanul[V]+t[Me3]                 she learnt (a)
Vmis1p------n-   tanul[V]+tunk[Mt1]              we learnt (a)
Vmis2p------n-   tanul[V]+tatok[Mt2]             you learnt  (a)
Vmis3p------n-   tanul[V]+tak[Mt3]               they learnt (a)
Vmis1s------y-   tanul[V]+tam[TMe1]              I learnt (the)
Vmis2s------y-   tanul[V]+tad[TMe2]              you learnt (the)

Vmis3s------y-   tanul[V]+ta[TMe3]               she learnt (the)
Vmis1p------y-   tanul[V]+tuk[TMt1]              we learnt (the)
Vmis2p------y-   tanul[V]+ta'tok[TMt2]           you learnt  (the)
Vmis3p------y-   tanul[V]+ta'k[TMt3]             they learnt (the)
Vmcp1s------n-   tanul[V]+ne'k[Fe1]              (if) I learnt (a)
Vmcp2s------n-   tanul[V]+na'l[Fe2]              (if) you learnt (a)
Vmcp3s------n-   tanul[V]+na[Fe3]                      (if) she learnt (a)
Vmcp1p------n-   tanul[V]+na'nk[Ft1]             (if) we learnt (a)
Vmcp2p------n-   tanul[V]+na'tok[Ft2]            (if) you learnt  (a)
Vmcp3p------n-   tanul[V]+na'nak[Ft3]            (if) they learnt (a)
Vmcp1s------y-   tanul[V]+na'm[TFe1]             (if) I learnt (the)
Vmcp2s------y-   tanul[V]+na'd[TFe2]             (if) you learnt (the)
Vmcp3s------y-   tanul[V]+na'[TFe3]              (if) she learnt (the)
Vmcp1p------y-   tanul[V]+na'nk[TFt1]            (if) we learnt (the)
Vmcp2p------y-   tanul[V]+na'tok[TFt2]           (if) you learnt  (the)
Vmcp3p------y-   tanul[V]+na'k[TFt3]             (if) they learnt (the)
Vmmp1s------n-   tanul[V]+jak[Pe1]               (let) me learn (a)
Vmmp2s------n-   tanul[V]+ja'l[Pe2]              (let) you learn (a)
Vmmp3s------n-   tanul[V]+jon[Pe3]               (let) she learn (a)
Vmmp1p------n-   tanul[V]+junk[Pt1]              (let) we learn (a)
Vmmp2p------n-   tanul[V]+jatok[Pt2]             (let) you learn  (a)
Vmmp3p------n-   tanul[V]+janak[Pt3]             (let) they learn (a)
Vmmp1s------y-   tanul[V]+jam[TPe1]              (let) me learn (the)
Vmmp2s------y-   tanul[V]+jad[TPe2]              (let) you learn (the)
Vmmp3s------y-   tanul[V]+ja[TPe3]               (let) she learn (the)
Vmmp1p------y-   tanul[V]+juk[TPt1]              (let) we learn (the)
Vmmp2p------y-   tanul[V]+ja'tok[TPt2]           (let) you learn  (the)
Vmmp3p------y-   tanul[V]+ja'k[TPt3]             (let) they learn (the)
Vmip1s------2-   tanul[V]+lak[Ie1]               I learn you
Vmis1s------2-   tanul[V]+talak[IMe1]            I learnt you
Vmcp1s------2-   tanul[V]+na'lak[IFe1]           (if) I learnt you
Vmmp1s------2-   tanul[V]+jalak[IPe1]            (let) me learn you
Vmn-----------   tanul[V]+ni[INF]                to learn
Vmnp1s--------   tanul[V]+nom[INRe1]             for me to learn
Vmnp2s--------   tanul[V]+nod[INRe2]             for you to learn
Vmnp3s--------   tanul[V]+nia[INRe3]             for her to learn
Vmnp1p--------   tanul[V]+nunk[INRt1]            for us to learn
Vmnp2p--------   tanul[V]+notok[INRt2]           for you to learn
Vmnp3p--------   tanul[V]+niuk[INRt3]            for them to learn
Vaip1s--------   fog[V]+ok[e1]                   I shall

3. Adjective (A)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           qualificative  f

- -------------- -------------- -
2 Degree         positive       p
                 comparative    c
                 superlative    s
- -------------- -------------- -
3 Gender                        -
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 instrumental   i
                 illative       x
                 inessive       2
                 elative        e
                 allative       t
                 adessive       3
                 ablative       b
                 terminative    9
                 essive         w
           l.s.  temporalis     m
           l.s.  causalis       c
           l.s.  sublative      s
           l.s.  delative       h
           l.s.  sociative      q
           l.s.  factive        y
           l.s.  superessive    p
           l.s.  distributive   u
           l.s.  essive_formal  f
* ***************************** *
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Animate        no             n
                 yes            y
- -------------- -------------- -
9 Formation                     -
- -------------- -------------- -
10Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
11Owner_Person   first          1
                 second         2
                 third          3
---------------- -------------- -
12Owned_Number   singular       s
                 plural         p
=================================

Notes:
1. Adjectives have the same suffixes that Nouns have plus Degree.
   Thus the number of possibble forms is 3*924=2772.
2. Types like ordinal and cardinal are grouped as numerals.

Combinations:
--- ---- ----- ---- ------------------------ ----  -------- --------- --------
Pos Type Degr  Num  Case                     Anim  OwnerNum OwnerPers OwnedNum
--- ---- ----- ---- ------------------------ ----  -------- --------- --------
 A  f    [pcs] [sp] [ngdaix2et3b9wmcshqypuf] [yn]  [sp]     [123]     [sp]
--- ---- ----- ---- ------------------------ ----  -------- --------- --------

Examples on combinations:
0123456789012
Afp-sn-------    nagy[A]                           big, great
Afc-sn-------    nagy[A]+obb[FOK]                  bigger
Afs-sn-------    leg[FF]+nagy[A]+obb[FOK]          biggest
Afs-sn----p1s    leg[FF]+nagy[A]+obb[FOK]+jaink[PSt1i]+e'[POS]
                                             sg that our greatest (heroes) have
Afp-sw-------    magyar[A]+ul[ESSMOD]              in Hungarian

4. Pronoun (P)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           personal       p
                 demonstrative  d
                 indefinite     i
                 possessive     s
                 interrogative  q
                 relative       r
                 reflexive      x
                 reciprocal     y
- -------------- -------------- -
2 Person         first          1
                 second         2
                 third          3
- -------------- -------------- -
3 Gender                        -
- -------------- -------------- -
4 Number         singular       s
                 plural         p

- -------------- -------------- -
5 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 instrumental   i
                 illative       x
                 inessive       2
                 elative        e
                 allative       t
                 adessive       3
                 ablative       b
                 terminative    9
                 essive         w
           l.s.  temporalis     m
           l.s.  causalis       c
           l.s.  sublative      s
           l.s.  delative       h
           l.s.  sociative      q
           l.s.  factive        y
           l.s.  superessive    p
           l.s.  distributive   u
           l.s.  essive_formal  f
- -------------- -------------- -
6 Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
7 Owner_Gender                  -
*********************************
8 Clitic                        -
- -------------- -------------- -
9 Referent_Type                 -
- -------------- -------------- -
10Syntactic_Type                -
- -------------- -------------- -
11Definiteness                  -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
- -------------- -------------- -
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person   first          1
                 second         2
                 third          3
- -------------- -------------- -
16Owned_Number   singular       s
                 plural         p
=================================

Combinations:

--- ---- ----- ---- ------------------------  -----------  ------------ ------------
Pos Type Pers  Num  Case                      Owner_Number Owner_Person Owned_Number
--- ---- ----- ---- ------------------------  ------------ ------------ ------------
 P  [px] [123] [sp] [ngdaix2et3b9wmcshqypuf]  [sp]         [123]        [sp]
 P    y     3   p   [ngdaix2et3b9wmcshqypuf]  [sp]         [123]        [sp]
 P    s  [123] [sp] [ngdaix2et3b9wmcshqypuf]  -             -
 P [diqr]  3   [sp] [ngdaix2et3b9wmcshqypuf]  [sp]          [123]       [sp]

Examples on combinations:
01234567890123456
Pp1-s------------         e'n[P]                   I
Pp1-p-s--------1-         enye'im[P]               my things (plural)
Pd3-s------------         ez[P]                    this
Pi3-s------------         minden[P]                every
Ps1-s-s--------1-         e'n[P]=enye'm+[POS]      my
Pq3-s------------         ki[P]                    who (in questions)
Pr3-s------------         aki[P]                   who (relative)
Px3-s------------         maga[P]                  himself
Pp3-p------------         ti[P]                    you
Pp3-s-p--------2-         tietek[P]                your something
Pp3-p-p--------2-         tieitek[P]               your somethings
Pp3-s-p--------2s         tieteke'[P]              your something
Py3-p------------         egyma's[P]               each other

5. Determiner (D)
Not applicable.

6. Article (T)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           definite       f
                 indefinite     i
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
*********************************
5 Clitic                        -
=================================

Notes:
We have tree articles: a, az and egy.  These may not have number and
case. The word 'az' may have but that is a pronoun in those cases.

Combinations:
--- ----
Pos Type
--- ----
T   [fi]
01
012345
Tf----              a[DET]              the
Tf----              az[DET]             the
Ti----              egy[DET]            a

7. Adverb (R)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           general        g
                 particle       p
           l.s.  causal         o
           l.s.  verbal         v
                 modifier       m
           l.s.  interrogative  q
- -------------- -------------- -
2 Degree                        -
*********************************
3 Clitic         no             n
                 yes            y
- -------------- -------------- -
4 Number         singular       s
                 plural         p
- -------------- -------------- -
5 Person         first          1
                 second         2
                 third          3
=================================

Combinations:
----  ------ ------  ------  ------ ------
Pos   Type   Degree  Clitic  Number Person
----  ------ ------  ------  ------ ------

 R    [gmp]   -        -        -     -
 R       q    -        y        -     -
 R       o    -        -        [sp]  [123]
----  ------ ------  ------  ------ ------

Examples on combinations:
012345
Rg----         most[R]                               now
Rg----         hogyan[R]                             how
Rgp---         belu:l[R]                             inside
Rgc---         belu:l[R]=belj+ebb[FOK]               deeper inside
Rgs---         leg[FF]+belu:l[R]=belj+ebb[FOK]       deepest inside
Rv----         olvas[V]+va[HIN]                      being read
Rm----         nem[R]                                no
Rp----         be[IK]                                into
Rq-y--         kutya[N]+-e[KSZE] (kutya-e?)          dog?
Ro--s1         a'll[V]+tomban[HIVe1]                 as I stand
Ro--s2         a'll[V]+todban[HIVe2]                 as you stand
Ro--s3         a'll[V]+ta'ban[HIVe3]                 as she stands
Ro--p1         a'll[V]+tunkban[HIVp1]                as we stand
Ro--p2         a'll[V]+totokban[HIVp2]               as you stand
Ro--sp         a'll[V]+tukban[HIVep]                 as they stand
Ro----         a'll[V]+va'n[HINN]                    as sy stands

Notes:
1. The modifier -e question word (the only Hungarian clitic) is attached
   to the preceding word with a hyphen.
2. Some adverbs may have degree, but these do not formulate a special
   class. Presently we cannot give a criteria for this.

8. Adposition (S)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           postposition   t
- -------------- -------------- -
2 Formation                     -
*********************************
3 Case                          -
- -------------- -------------- -
4 Clitic                        -
=================================

Notes:

1. Some of them require a suffix on the preceding noun (e.g. asztal[N]+
   on[SUP] tu'l[S]) while others do not (e.g. asztal[N] mellett[S]).
   This feature is not represented here.

Combinations:
---  ----
Pos  Type
---  ----
 S     t
---  ----
Examples on combinations:
01234
St---      miatt[S]                      because of
St---      keresztu:l[S]                 through

9. Conjunction (C)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           coordinating   c
                 subordinating  s
*********************************
2 Formation      simple         s
                 compound       c
- -------------- -------------- -
3 Coord_Type     sentence       p
                 words          w
- -------------- -------------- -
4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- -
6 Number                        -
- -------------- -------------- -
7 Person                        -
=================================

Notes:
1. 'form': Some conjuctions in Hungarian might occur alone (1) while
    others in pairs (2).

Combinations:

---  ---- ---------- ---------
Pos  Type Coord-Type Formation
---  ---- ---------- ---------
 C    [cs]   [pw]       [12]
---  ---- ---------- ---------

Examples on combinations:              Meaning:
01234567
Csp-1---          hogy[C]                that
Ccp-1---          e's[C]                 and
Ccw-2---          is[C]                  also
Ccw-2---          mind[C]                both

10. Numeral (M)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           cardinal       c
                 ordinal        o
                 fractal        f
                 collect        l
- -------------- -------------- -
2 Gender                        -
- -------------- -------------- -
3 Number         singular       s
                 plural         p
- -------------- -------------- -
4 Case           nominative     n
                 genitive       g
                 dative         d
                 accusative     a
                 instrumental   i
                 illative       x
                 inessive       2
                 elative        e
                 allative       t
                 adessive       3
                 ablative       b
                 terminative    9
                 essive         w
           l.s.  temporalis     m
           l.s.  causalis       c
           l.s.  sublative      s
           l.s.  delative       h
           l.s.  sociative      q
           l.s.  factive        y
           l.s.  superessive    p

           l.s.  distributive   u
           l.s.  essive_formal  f
           l.s.  multiplicative 6
*********************************
5 Form           digit          d
                 roman          r
                 letter         l
- -------------- -------------- -
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Class                         -
- -------------- -------------- -
9 Animate                       -
- -------------- -------------- -
10Owner_Number   singular       s
                 plural         p
- -------------- -------------- -
11Owner_Person   first          1
                 second         2
                 third          3
- -------------- -------------- -
12Owned_Number   singular       s
                 plural         p
=================================

Combinations:
--- ---- ---- ------------------------ ----  -------- --------- --------
Pos Type Num  Case                     Form  OwnerNum OwnerPers OwnedNum
--- ---- ---- ------------------------ ----  -------- --------- --------
 N  [cp] [sp] [ngdaix2et3b9wmcshqypuf] [dlr] [sp]    [123]     [sp]
--- ---- ---- ------------------------ ----  -------- --------- --------

Examples on combinations:
0123456789012
Mc-sn--------        egy[M]                  one
Mc-pn--------        nyolc[M]+ak[PL]         those eight (men)
Mc-pn-----p3s        nyolc[M]+ak[PL]+e'[POS] this belongs to those eight (men)
Mc-sa--------        sok[M]+at[ACC]          many (in accusative)
Mo-sn--------        hat[M]+odik[SORSZ]      6th
Mf-sn--------        hat[M]+od[TORT]         1/6
Mc-s6--------        egy[M]+szer[MUL]        one time
Mc-sw--------        hat[M]+an[ESSMOD]       six of them

11. Interjection (I)

= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Type           mood           m
                 other          o
- -------------- -------------- -
2 Formation                     -
=================================

12. Residual (X)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
                                -
=================================

13. Abbreviation (Y)
= ============== ============== =
P ATT            VAL            C
= ============== ============== =
1 Syntactic_Type                -
- -------------------------------
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
- -------------- -------------- -
5 Definiteness                  -
=================================

14. Particle (Q)
Not applicable.

3.4.1 Derivation

Without handling derivation a satisfactory morphological analysis is not possible for Hungarian. The HUMOR system, a general purpose morphological analyzer, handles it and the results can be converted to multext format since on syntactic level the morphological origin of stems are generally irrelevant. The resulting word class is defined by the rightmost derivational suffix. The suffix characters are literally attached to the word.

Here are the derivations that the analyser recognizes but instead of the origin we place only the resulting class to the output. (Suffix tags used in HUMOR are in upper case, actual suffixes are in lower case.)

Noun   -> Adjective
-------------------
-----   -------        --------------    ------------------------
Type    Form           Example           Comment
-----   -------        --------------    ------------------------
 BELI  : beli           ha'z+beli        property of living in the house
 FAJTA : fajta          ma's+fajta       of some other kind
 FELE  : fe'le          bu'tor+fe'le     similar to furniture
 FORMA : forma          toja's+forma     egg shaped
 SZERU : szeru"         tej+szeru"       milk+y
 IKEP  : i              ha'z+i           home (e.g. made)
 SKEP  : s, as, os,
         es, o:s        gyerek+es        child+ish
 UKEP  : u',
         u",
         ju',
         ju"            arc+u'           (red)-face+d
 FFOSZ : tlan,tlen,
         atlan,etlen,
         talan,telen                     'devoid of', '-less'
 MER   : nyi            kana'l+nyi       spoon+full

Noun,A -> Noun
--------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 COL   : sa'g,         bara't+sa'g            friend+ship
         se'g

Noun   -> Noun
--------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 DIM   : cska,acska,
        ecske,o:cske,
        ocska          utca'+cska             little street
 FEM   : ne'           Kova'cs+ne'            Mrs. Kova'cs

Noun   -> Verb
--------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 FI    : z,az,oz,ez,
        o:z            auto'+z                go by car

Verb   -> Adjective
-------------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 IFOSZT: atlan, etlen  felel+etlen            sg not being answered
 MIF   : o',
         o"            felel+o"               sy who answers
 MIB   : t,ott,ett,
         o:tt          felel+t                the answered (question)
 MIA   : ando',
         endo"         felel+endo"            sg that should be answered
 NIVALO: anivalo',
         enivalo',
         nivalo'       ne'z+nivalo'           sg that should be seen

Verb   -> Adverb
----------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 HIN   :  va,ve        olvas+va              (while ) reading (the book)

Numeral-> Adjective
-------------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 KIEM  : ik            hatod+ik               six+th
 LAGOS : lagos,leges   ma'sod+lagos           second+ary

Verb   -> Noun
--------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 IF    : a's,
         e's           olvas+a's              read+ing (gerund)
 DES   : hatne'k,
        hetne'k        olvas+hatne'k          the intention of reading

Adj    -> Verb
--------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 FAK   : i't           sze'p+i't              make it pretty (in compounds only)

 MI    : od,ed,o:d     va'llas+od+ik          becomes strong
 MIGY  : kod,ked,
         ko:d          okos+kod+ik            plays the smart (frequently)

Verb   -> Verb
--------------
-----   -------        --------------         ------------------------
Type    Form           Example                Comment
-----   -------        --------------         ------------------------
 MUV   : at,et,tat,tet  olvas+tat             makes him read
 GYAK  : gat,get,ogat,
         eget,o:get     olvas+gat             he reads frequently
 HAT   : hat,het        olvas+hat             he may read
 VISSZ : o'd,
         o"d            old+o'dik             dissolves
 SZENV : tatik, tetik   olvas+tatik           makes the book being read

Examples:

If ’szemtele’s’ (littering, action of throwing away litter) is not in the dictionary we derive it from the verb ’litter’ szemetel[V] + e’s[IF] (where IF=Verb2Noun).

Instead of giving the extra attribute to the verb expressing that it has a derivational suffix we simply give the result of the analysis+conversion: szemetele’s[N].

In Hungarian some derivation may follow the inflectional suffix. For these derivations the suffix+derivation together forms a compound derivation. The a new stem is generated from the stem + inflection + derivation segments and the resulting part of speech is determined by the derivation.

------------------     -----   --------------     ------------------------
Type                   Form    Example            Comment
------------------     -----   --------------     ------------------------
Nc-sn--ns1-   +FAM  :   e'k    ap +m+e'k          some people with my father
Nc-su--u---   +IKEP :   i      asztal+onke'nt+i   sg. done by each every table
Afc-sn--n---- +KIEM :   ik     nagy+obb+ik        the bigger one

There are adverbs that may get case endings. Since case inflections derivate adverb from nouns these constructions can be handled as derivations. That means the stem is the stem + inflection combination and the part of speech is adverb.

------------------     -----   --------------     ------------------------
Type                    Form   Example            Comment
------------------     -----   --------------     ------------------------
Ag----     + ablativ    t"l    akkor+t"l          since then

3.4.2 Compounding

Compounding is handled in a very similar way to derivation. The rightmost word class is always the resulting one. If it contains some derivation as well then the result is the word class that the derivation determines.

Examples:

1. Hungarian is similar to German but we have a competence limit: we do not put together more then 2 nouns. These two words might be compounds as well but they must be lexicalized forms. So e.g.: rendo”regyenruha (police uniform) is valid since it is put together from rendo”r (police) and egyenruha (uniform) where rendo”r=rend (order) + o”r (guard) and egyenruha=egyen (uni) + ruha (clothes).

2. Of course the situation is more complicated since nouns can be derived from verbs, adjectives or other nouns, and these derivations can be parts of compoundns: e.g. auto’buszmega’llo’ (bus stop) where meg+a’ll+o’ is derived from the verb a’ll (stand). Whereas mega’llo’ is lexicalized so dictionaries might contain it, the morphological rule is productive so compounding works well for constructions like tandi’jbefizete’s (paying the tution fee)= tan(tution)+ di’j(fee)+ be(in) +fizet(pay) +e’s(ment).

3. Compounding, such as derivation, will not be mirrorred in the multext format morphological analysis* a converter will produce an acceptable segmentation in the corpus. The word class (POS) for these compounds will be the word class of the resulting compounding (e.g. tandi’jbefizete’s/N).

COP project 106 multext-East  Deliverable D1.1 F— Romanian  130

3.5 Application to Romanian

The application to Romanian has been elaborated by D.Tufiş and A.M.Barbu.

The Romanian diacritic characters used have been encoded in the following way:

• ă as a(

• ş and ţ as s, and t,

• â and î as a> and i>

1. Noun (N)
In Romanian the following attribute-value pairs are applicable to Nouns:

1.1 Type
Attribute           value                Ro. example
Type                common  (c)          carte
                    proper  (p)          Ion

1.2 Gender
Attribute           value                Ro. example
Gender              masculine  (m)       ba(iatul
                    feminine   (f)       casa
                    neuter     (n)       fir (m.sg.),fire (f.pl.)
In Romanian the declension of a neuter noun always follows in singular
a masculine paradigm and in plural a feminine one. Specific
implementations could take advantage of this rule and by organizing
the paradigmatic space in partial paradigms (masc-sing, masc-pl,
fem-sing, fem-pl) to get rid of neuter value for the gender attribute.

1.3 Number
Attribute           value                Ro. example
Number              singular  (s)        fata(
                    plural    (p)        fete

1.4 Case
There are five functional cases in Romanian (nominative, genitive,
dative, accusative and vocative) but they are expressed by at most
three syncretic forms for each number: 1. nominative-accusative;
2. genitive-dative; 3. vocative. The ambiguity of a syncretic case
form may be solved at the syntactic level. Taking into account this
syncretism we considered the following attribute-value combinations.

Attribute           value                Ro. example
Case                direct(d)            omul
                    oblique (o)          omului
                    vocative (v)         omule
At the word level, the distinction between the three syncretic forms
is possible only for the definite nouns. Masculine indefinite nouns
are not differentiated by case (except, sometimes for vocative. The
feminine indefinite nouns have different forms in singular for
nominative-accusative, genitive-dative and vocative, but the
genitive-dative singular forms are also indistinguishable from the
plural forms:
fata(      - nominative-accusative, vocative singular, indefinite
fato       - vocative singular, indefinite
fete       - genitive-dative singular, nominative-accusative,
             genitive-dative, vocative plural, indefinite
ba(iat     - nominative-accusative, genitive-dative, vocative
             singular, indefinite
ba(iete    - vocative singular, indefinite
ba(iet,i   - nominative-accusative, genitive-dative, vocative,
             plural, indefinite

1.5 Definiteness
In Romanian, nouns can be marked for definiteness with the enclitic
definite article.
Attribute           value                Ro. example
Definiteness        yes (y)              omul
                    no (n)               om

1.6 Clitic
See discussion in 2.10
Attribute           value                Ro. example
Clitic              no (n)               sora
                    yes (y)              soru-mea

Combinations
Here, and in all the sections to follow, the trailing hyphens are deleted.
=========================
Tag          Example
=========================
Ncmsrn       frate
Ncmson       frate

Ncmsvn       frate
Ncmsry       fratele
Ncmsoy       fratelui
Ncmprn       frat,i
Ncmpon       frat,i
Ncmpvn       frat,i
Ncmpry       frat,ii
Ncmpoy       frat,ilor
Ncmpvy       frat,ilor
Ncfsrn       sora(
Ncfsvn       soro
Ncfson       surori
Ncfprn       surori
Ncfpon       surori
Ncfpvn       surori
Ncfsoyy      sora((-sii)
Ncfpry       surorile
Ncfpoy       surorilor
Ncfpvy       surorilor
Ncfsryy      sora((-sa)
Ncmsrn       creion
Ncmson       creion
Ncmsryy      creionu-(i)
Ncmsoy       creionului
Ncfprn       creioane
Ncfpon       creioane
Ncfpry       creioanele
Ncfpoy       creioanelor
Npfsr        Ioana
Npfso        Ioanei
Npmsrn       Bucures,ti
Npmsry       Bucures,tiul

2. Verb (V)
The following attribute-value pairs are applicable to verbs in Romanian:

2.1 Type
Attribute           value                Ro. example
Type                main (m)             a vedea
                    auxiliary (a)        a avea, a fi, a voi
                    modal (o)            a putea, a trebui
                    copulative (c)       a fi, a deveni

2.2 VForm
Traditionally, Romanian linguistics distinguishes between predicative
and non-predicative moods. This distinction may be easily mapped into

finite/non-finite dichotomy: indicative, subjunctive and imperative
are finite; infinitive, participle and gerund are non-finite (only
synthetic (non-compound) moods were mentioned; we use the opposition
synthetic-analytic to distinguish between concatenative (synthetic)
and compound (analytic) morpho-lexical phenomena).
Attribute           value                Ro. example
VForm               indicative(i)        vine
                    subjunctive (s)      vina(
                    imperative (m)       vino
                    infinitive (n)       veni
                    participle (p)       venit
                    gerund (g)           venind
As only synthetic forms were considered, the values conditional and
presumptive for the VForm attribute were left out in Romanian. Another
value for VForm which was left out is the Supine. It appears mostly
with a preposition, except for a few intransitive verbs when they are
subordinated to the impersonal verb a trebui (must). Only the
preposition allows for differentiating a supine from a
participle-masculine-singular.

2.3 Tense
The synthetic values for the attribute Tense, as listed in the table
below, apply to the indicative mood. The value p (present) is used
also for subjunctive and infinitive. All the other moods either have
no tense or have compound form tenses.
Attribute           value                Ro. example
Tense               present (p)          va(d
                    imperfect (i)        vedeam
                    past (s)             va(zui
                    pluperfect (l)       va(zusem

2.4 Person
Attribute           value                Ro. example
Person              first (1)            va(d
                    second (2)           vezi
                    third (3)            vede
The following features are pertinent to those moods which permit an
adjectival use, i.e. participle and gerund. However, the adjectival
use of gerund is extremely rare (o ma>na( tremurnda(
- a shaking hand) and therefore gender and number apply mainly for the
participle.

2.5 Number

Attribute           value                Ro. example
Number              singular (s)         ba(tut
                    plural (p)           ba(tut,i

2.6 Gender
Attribute           value                Ro. example
Gender              masculine(m)         ba(tut
                    feminine(f)          ba(tuta(
                    neuter(n)            ba(tut-ba(tute

2.7, 2.8, 2.9 Voice, Negation, Definiteness
As these feature values are either realized as compound forms or
irrelevant for Romanian, in our word-level encoding they were omitted.

2.10 Clitic
The cliticization phenomenon in Romanian is not restricted to
verb-pronoun relationship, but may also be observed with the (main)
verb and the auxiliary, the noun or adjective with pronoun, with noun
or adjective with copula, pronoun with auxiliary, preposition with
(indefinite) article, numeral or (indefinite) pronoun, negative adverb
with verb, auxiliary or pronoun, and some others (mainly created
through the contracted forms of the verb "a fi"-to be). We restrict
ourselves to considering only the graphically marked clicitizations.
In such cases, the two, three or (sometimes) four constituents of a
cliticized word-form are always separated by a hyphen. Omitting the
hyphen in such cases is an unacceptable error in written Romanian.
The examples below illustrate the specific types of cliticization we
will consider:
da(-mi-l !                (you give +to me+it/him = give it/him to me)
la(satu-ne-ai             (left + us+ you have = you have left us)
sparge-t,i-s-ar lampa     (break + your + s-part-conj + has lamp =
                           I wish your lamp would break)
sora(-mii                 (sister + my = to my sister)
fat,a-i                   (his/her face | the face is)
ros,u-i                   (his/her red | the red is )
m-au                      (me + they have )
i>ntr-o gaura(            (in + a hole)
i>ntr-o ora(              (in + one hour)
i>ntr-unele               (into + some of)
n-aud                     (not  + I hear = I do not hear)
n-am                      (not + I/we have = I have not )
nu-mi                     (not + to me )
The order in which different constituents of a cliticized form appear
is governed by precise morphological rules. For instance, the
auxiliaries always appear in the last position. The main verbs, except

those beginning either with the letter "a" or the letter i> and
the contracted forms (-s and -i) of the verb "a fi"- to be, always
appear in the first position, nouns and adjectives always precede the
cliticized pronouns, the negative adverbial particles nu- and n- only
appear in the first position, and so on.  However, in order to reduce
spurious ambiguity in morpho-lexical encoding, we considered the
attribute CLITIC as relevant only in those cases where the
clicitization phenomenon resulted in graphemic modification of the
cliticized term (as it is the epenthetic u in the gerund form shown in
the table below).
Attribute           value                Ro. example
Clitic              no (n)               am ridicat
                    yes (y)              ridica>ndu-l

Combinations
========================
Tag           Example
========================
Vmii1s        abandonam
Vmii2s        abandonai
Vmii3s        abandona
Vmii1p        abandonam
Vmii2p        abandonat,i
Vmii3p        abandonau
Vmis1s        abandonai
Vmis2s        abandonas,i
Vmis3s        abandona(
Vmis1p        abandonara(m
Vmis2p        abandonara(t,i
Vmis3p        abandona(
Vmil1s        abandonasem
Vmil2s        abandonases,i
Vmil3s        abandonase
Vmil1p        abandonasera(m
Vmil2p        abandonasera(t,i
Vmil3p        abandonasera(
Vmip1s        abandonez
Vmsp1s        abandonez
Vmip2s        abandonezi
Vmsp2s        abandonezi
Vmip3s        abandoneaza(
Vmip3p        abandoneaza(
Vmsp3s        abandoneze
Vmsp3p        abandoneze
Vmsp1p        abandona(m
Vmsp2p        abandonat,i
Vmm-2s        abandoneaza(
Vmm-2p        abandonat,i
Vmnp          abandona
Vmp--sm       abandonat
Vmp--sm---y   abandonatu

Vmp--sf       abandonata(
Vmp--pf       abandonate
Vmp--pm       abandonat,i
Vmg           abandona>nd
Vmg-------y   abandona>ndu
Va--1s        as,
Voip          trebuie
Vcip1s        sunt
Vcip3p        sunt
Vcip1s----y   -s
Vcip3p----y   -s

3. Adjective (A)

3.1 Type
Although it is not common practice in Romanian linguistics, one could
make the distinction between qualificative and determinative
adjectives. However, the attribute-value pairs proposed in this
section are appropriate for the qualificative adjectives.
Attribute           value                Ro. example
Type                qualificative (f)    frumos

3.2 Degree
The default value is positive; adjectives have also comparative and
superlative degrees, but in most of the cases they are expressed by
means of analytical forms (e.g. comp. mai bun (better), superl. cel
mai bun (the best)). A few adjectives have intrinsic etymological
comparative or superlative meanings (e.g. comparatives: anterior,
major; superlatives: optim, maxim, extrem etc.). The prefixes super-,
extra-, ultra- etc., are quite productive in forming the
quasi-analytic1 superlatives.  Adjectives are characterized by gender,
number and case.
Attribute           value                Ro. example
Degree              positive (p)         frumos
                    comparative (c)      ulterior
                    superlative (s)      extrem

3.3 Gender
Attribute           value                Ro. example
Gender              masculine (m)        bun
                    feminine (f)         buna(
                    neuter (n)           sg.bun/pl.bune

3.4 Number
Attribute           value                Ro. example
Number              singular (s)         bun
                    plural (p)           buni

3.5 Case
The adjectives present the same case syncretism as the nouns, except
for few adjectives that have an additional special form for
Genitive-Dative cases in plural (e.g. G.D.pl.  multor).
Attribute           value                Ro. example
Case                direct(r)            bunul
                    oblique (o)          bunului
                    vocative (v)         bunule

3.6 Definiteness
In noun-adjective construction, the definite article may attach
enclitically to either adjectives or modified nouns (never to both of
them). If present, the definite article attaches to the right of the
first word in the sequence.
Bunul  om  (The kind man)
Omul  bun. (The kind man)
Attribute           value                Ro. example
Definiteness        yes (y)              bunul
                    no (n)               bun

3.7 Clitic
See discussion on clitics in 2.10

Combinations
====================
Tag            Example
====================
Afpmsrn        bun
Afpmson        bun
Afpmsvn        bun
Afpmprn        buni
Afpmpon        buni
Afpmpvn        buni
Afpmsry        bunul
Afpmsoy        bunului

Afpmpry        bunii
Afpmpoy        bunilor
Afpfsrn        buna(
Afpfsvn        buna(
Afpfson        bune
Afpfprn        bune
Afpfpon        bune
Afpfpvn        bune
Afpfsry        buna
Afpfsoy        bunei
Afpfpry        bunele
Afpfpoy        bunelor
Afcmsrn        ulterior
Afcmson        ulterior
Afsmsrn        extrem
Afp            gri

4. Pronoun (P)
4.1 Type
In Romanian it is worth differentiating the negative pronoun from
other indefinite pronouns: a negative pronoun cannot be an argument
for a verb unless the verb itself is negated too (e.g. Nu am
va(zut pe nimeni / *Am va(zut pe nimeni).
Attribute           value                Ro. example
Type                demonstrative(d)     acesta
                    indefinite (i)       oricine
                    possesive(s)         (al) meu
                    int_rel (w)          ce
                    personal(p)          eu
                    reflexive(x)         se
                    negative (z)         nimeni
                    emphatic (h)         i>nsumi

4.2 Person
Attribute           value                Ro. example
Person              first (1)            eu
                    second (2)           tu
                    third (3)            el

4.3 Gender
Attribute           value                Ro. example
Gender              masculine (m)        el
                    feminine (f)         ea
                    neuter (n)           sg.acesta/pl.acestea

4.4 Number
Attribute           value                Ro. example
Number              singular (s)         eu
                    plural (p)           noi

4.5 Case
For the second person of the personal pronoun in both singular and
plural there is the vocative case too.The direct and oblique values
are needed for the syncretic causal forms of the other pronouns than
the personal ones.
Attribute           value                Ro. example
Case                nominative (n)       el
                    genitive (g)         (al) lui
                    dative (d)           lui
                    accusative (a)       (pe) el
                    vocative (v)         tu, voi!
                    direct(r)            acesta
                    oblique (o)          acestuia

4.6 Owner_Number
This attribute is meaningful for the Possessive pronouns and refers to
the grammatical number of the possessor.
Attribute           value                Ro. example
Owner_Number        singular (s)         meu
                    plural (p)           nostru
The gramatical number of the possesed object(s) is expressed by the
attribute Number (described above).

4.7 Owner_Gender
This attribute is irrelevant for Romanian.

4.8 Clitic
See discussion on clitics in 2.10.

4.9, 4.10, 4.11, 4.12, 4.13 Referent_Type, Syntactic_Type,
Definiteness, Animate, Clitic_s
These attributes are irrelevant for Romanian.

4.14 Pronoun_Form
For Romanian we need an attribute (called Pronoun_Form) to make the
distinction between strong and weak forms of the same pronoun. All the
weak forms can be adjoined to the adjacent words both proclitically or
enclitically. In such cases the junction is always graphically marked
by a hyphen between the pronoun and the neighboring word. The hyphen
also marks possible elisions from either pronoun or the adjacent word.
Although in traditional grammar books the demonstrative, int_rel and
indefinite pronouns are not characterised by person, in our
dictionaries they are recorded (for reasons beyond morpho-lexical
encoding) as 3rd person (the same as nouns). However, for the
automatic tagging this value has been marked as irrelevant.
Attribute           value                Ro. example
Pronoun_Form        strong (s)           lui
                    weak (w)             i>i, i-

Combinations
====================================
Tag                    Example
====================================
Pp1msn--------s        eu
Pp1msd--------w        mi
Pp1msd--------s        mie
Pp1msd--y-----s        mi-
Pd-msr                 acesta
Pd-mso                 acestuia
Pi-mpr                 tot,i
Ps1fsrs                mea
Pw-mso                 ca(rui
Pn-msr                 nimeni
Ph1msr                 i>nsumi
Ph1fsr                 i>nsa(mi
Px3msa--------s        sine
Px3msa--------w        se
Px3msa--y-----w        s-

5. Determiner (D)

5.1 Type
The need for a negative value of the determiners' Type attribute is
argued on the same lines as in the section on pronoun' s Type.  In
Romanian the negative determiner is expressed by the unit nici +
indefinite article (e.g. nici un, nici o).  In Romanian, there are
specific forms for the so-called emphatic determiner, which may

accompany both a noun and a personal pronoun: fata
i>nsa(s,i (the girl herself), also ea
i>nsa(s,i (she herself).
Attribute           value                Ro. example
Type                demonstrative (d)    acest
                    indefinite (i)       orice
                    possessive (s)       meu
                    int_rel (w)          ce
                    negative (z)         nici un
                    emphatic (h)         i>nsus,i

5.2 Person
This attribute is meaningful for the Possessive determiners and refers
to the grammatical person of the possessor.
Attribute           value                Ro. example
Person              first (1)            meu
                    second (2)           ta(u
                    third (3)            sa(u

5.3 Gender
Attribute           value                Ro. example
Gender              masculine (m)        meu
                    feminine (f)         mea
                    neuter (n)           sg.meu/pl.mele

5.4 Number
Attribute           value                Ro. example
Number              singular (s)         meu
                    plural (p)           mei

5.5 Case
Attribute           value                Ro. example
Case                direct(d)            aceasta
                    oblique (o)          acestei

5.6 Owner_Number
This attribute is meaningful for the Possessive determiners and refers
to the grammatical number of the possessor.
Attribute           value                Ro. example
Owner_Number       singular (s)          meu

                   plural (p)            nostru

5.7 Owner_Gender
This attribute is irrelevant for Romanian.

5.8 Clitic
See discussion on clitics in 2.10.
Attribute           value                Ro. example
Clitic              no (n)               mama mea
                    yes (y)              maica(- mea

5.9 Modific_Type
As mentioned in the corresponding section on Pronoun, the Modific_Type
attribute is relevant for some determiners too. The prenominal
determiner always precedes the noun (e.g.acest ba(iat - this
boy), whereas the postnominal determiner appears only after the noun
(e.g. ba(iatul acesta - this boy).
Attribute           value                Ro. example
Modific_Type        prenominal (e)       acest
                    postnominal (o)      acesta

Combinations
=======================
Tag           Example
=======================
Dd-mso---e    acestui
Dd-mso---o    acestuia
Di-mpr        tot,i
Ds1fsrs       mea
Dw-msr        care
Dw-mso        ca(rui
Dz-msr        nici_un
Dh1msr        i>nsumi
Dh1fsr        i>nsa(mi
Although in traditional grammar books the demonstrative, indefinite
and int_rel determiners are not characterised by person, in our
dictionaries they are recorded (for reasons beyond morpho-lexical
encoding) as 3rd person (the same as nouns). However, for the
automatic tagging this value has been marked as irrelevant.

6. Article (T)

6.1 Type
Although it presents only a few items, the article in Romanian has
four types, unlike in most of the European languages. Beside the two
recommended types: definite and indefinite which have the generally
known semantic value, Romanian uses two additional types of articles,
which are semantically subordinated to the definite article but which
have special forms and meanings:
- the possessive article (also called genitival article) is an element
  in the structure of the possessive pronoun, of the ordinal numeral
  (e.g. al meu (mine) and al treilea (the third)), and of the indefinite
  genitive forms of the nouns (e.g. capitol al ca(rt,ii
  (chapter of the book)).
- the demonstrative article links a definite noun to its determinants,
  links a numeral or an adjective to a noun, and it is a constituent
  part of the relative superlative (e.g.  fata cea mare (the elder
  girl), cel lenes, (the lazy), respectively prietenul cel mai bun
  (the best friend)).
Notice that the definite article has only enclitic forms, except for
one proclitical form (lui + proper noun: lui Ion).
Attribute           value                Ro. example
Type                definite (f)         lui
                    indefinite (i)       un
                    possessive (s)       al
                    demonstrative (d)    cel

6.2 Gender
Attribute           value                Ro. example
Gender              masculine (m)        un
                    feminine (f)         o
                    neuter (n)           sg.cel/pl.cele

6.3 Number
Attribute           value                Ro. example
Number              singular (s)         un
                    plural (p)           nis,te

6.4 Case
Attribute           value                Ro. example
Case                direct(r)            cel

                    oblique (o)          celui

6.5 Clitic
The inflected forms of the foreign-origin words (mainly nouns) not
fully assimilated, are usually written with a hyphen between the
base-form and the inflectional ending.  In our encoding, we classified
these endings (which are supposed to be split by the segmenter) as
clitic articles (clitic attribute is always "y") which can be either
definite (type=f, "-istul") or indefinite (type=i, "ist") and are
characterised by gender (gender=m, "ist"; gender=f, "ista("),
number (number=s, "ist"; number=p, "is,ti") and case (case=r,
"istul"; case=o, "istului").

Combinations
=====================
Tag          Example
=====================
Tfmso        lui
Tffso        lui
Timsr        un
Tsmpr        ai
Tdfso        celei
Timsry       -ist
Timsoy       -ist
Tfmsry       -istul
Tfmsoy       -istului

7. Adverb (R)

7.1 Type
The distinction proposed here considers the principal syntactic
properties of the adverbs.  For Romanian, the general type includes
most of the pronominal adverbs (demonstrative: aici (here),
indefinite: oriunde (anywhere)). As argued before for pronouns and
determiners, a distinct negative value is needed for adverbs as well
(nica(ieri - nowhere, niciodata( - never). The particle
type covers those adverbs which can dislocate verbal compound forms
(ex. Ea a tot ca>ntat -- She has ever sung) or mark degrees
(ex. circa (about), foarte (very), prea (too)). Such adverbs are cam,
mai, prea, s,i, tot, foarte etc.  A useful distinction in
Romanian considers the adverbs which can have predicative role, that
is they can govern a subordinate sentence (ex. Fires,te
ca( o s,tiu -- Certainly I know it). Here (for uniformity
within a multilingual environment), they are squeezed into the
modifier class.  No formal distinction is made between the
interrogative adverbs and the relative ones.

Attribute           value                Ro. example
Type                general (g)          bine, acolo
                    particle (p)         mai, cam
                    negative (z)         nica(ieri
                    modifier (m)         fires,te, poate
                    int_rel(w)           cum

7.2 Degree
In Romanian, the comparative and superlative of adverbs is formed
analytically with mai (put,in), cel mai (put,in), foarte:
ex. mai repede (faster), cel mai devreme (at the
earliest). Nonetheless there are some adverbs with comparative or
superlative meaning (ex. optim, ulterior,. definitiv). These adverbs
can be used for expressing the absolute superlative of other adverbs
or adjectives: ex. extrem de bine, formidabil de frumos.
Attribute           value                Ro. example
Degree              positive (p)         bine
                    comparative (c)      ulterior
                    superlative (s)      extrem

7.3 Clitic
See discussion on clitics in 2.10.

Combinations
=====================
Tag        Example
=====================
Rgp        repede
Rgs        extraordinar
Rgc        ulterior
Rp         mai
Rz         nica(ieri
Rm         probabil
Rw         cum

8. Adposition (S)

8.1 Type
It is the preposition type which is only pertinent to Romanian,
although some intercalating adpositions may be seen as a sort of
circumposition, for instance i>ntre...s,i...
(between..and...).

Attribute           value                Ro. example
Type                preposition (p)      la, pe, i>n

8.2 Formation
In Romanian there is a distinct class of compound prepositions. Each
of them forms a formal and semantic unit, although graphically they
stay unfused, e.g. de la, pe la, de pe, etc.
Attribute           value                Ro. example
Formation           simple (s)           la, pe, i>n
                    compound (c)         de la

8.3 Case
This attribute marks the subcategorisation properties of the adpositions.
Attribute           value                Ro. example
Case                genitive (g)         i>naintea
                    dative (d)           datorita(
                    accusative (a)       la

8.4 Clitic
See discussion on clitics in 2.10.

Combinations
======================
Tag         Example
======================
Spsa        i>n
Spsay       i>ntr-
Spsd        datorita(
Spca        de_la

9. Conjunction (C)

9.1 Type
The distinction between coordinating and subordinating conjunctions is
pertinent to Romanian as well.
Attribute           value                Ro. example
Type                coordinating (c)     s,i, dar
                    subordinating (s)    ca(, daca(

9.2 Formation
Likewise prepositions, we can distinguish  two kinds of conjunctions in
Romanian:
- simple conjunctions: e.g. s,i,dar,des,i  etc.
- conjunctions formed periphrastically, with some word/phrase combined
  by a conjunction: din moment ce, fa(ra( sa(,
  fat,a( de cum etc.
Attribute           value                Ro. example
Formation           simple (s)           deoarece
                    compound (c)         de_vreme_ce

9.3 Coord_Type
In Romanian, there are three kinds of conjunctions depending on their
usage: as such or together with other conjunctions or adverbs:
- simple, between conjuncts: Ion s,i Maria (John and Mary);
- repetitive, before each conjunct: ori  Ion ori  Maria ori...
  (either John or Mary or...)
- correlative, before a conjoined phrase, it requires specific
  coordinators between conjuncts: ata>t mama ca>t s,i
  tata (both mother and father).
With respect to the place of the conjunctions, most of them stay
before the conjunct, except for: as,adar, deci (so), dar,
i>nsa( (but), daca( (if), which also appear with
expressive value inside the conjoined sentence.
Attribute           value                Ro. example
Coord-Type          simple (s)           s,i,deoarece
                    repetit (r)          fie...fie...
                    correlat (c)         ati>t...ca>t s,i

9.3 Sub_Type
In Romanian, each conjunction requires another mood, so that the
diversity may be controlled by subcategorisation rules.
Attribute           value                Ro. example
Sub-Type            negative (z)         nici
                    positive (p)         dar
This attribute distinguishes among the positive and negative

conjunctions, providing means to control verbal double negation, (as
in case of the negative pronouns, determiners and adverbs): nici NU am
venit, nimeni NU vorbes,te, nici_un tren N-a trecut,
nica(ieri N-am va(zut

9.5 Clitic
See discussion on clitics in 2.10.
Attribute           value                Ro. example
Clitic              no (n)               ca( as,a
                    yes (y)              c-as,a

Combinations
======================
Tag          Example
======================
Ccsps        s,i
Ccrps        fie...fie
Csrzs        nici...nici
Csspc        de_vreme_ce

10. Numeral (M)

10.1 Type
Traditional Romanian grammars usually distinguish seven numeral types,
where five of them have specific forms and the other two are obtained
by composition.
The first group is made up by the following numeral types: cardinal
(trei-three), ordinal (al treilea-the third), fractional (treime-one
third), multiple (i>ntreit-trine), collective
(ama>ndoi-both).
The second group contains the numeral types which are composed by
means of other parts of speech: distributive (ca>te trei-...each
three...), adverbial (de trei ori-thrice) and again the collective
numeral which also has compound forms (tot,i trei-all three).
Nonetheless, as the numerals of the second group have a weak syntactic
cohesion, namely each composition element may be regarded as an
element of the sentence, with its own grammatical function, these last
numeral types are irrelevant for the morphosyntactic annotation.
Attribute           value                Ro. example
Type                cardinal (c)         trei
                    ordinal (o)          (al) treilea
                    fractal (f)          treime

                    multiple (m)         i>ntreit
                    collect(l)           tustrei
In Romanian (as in many other languages) several numerals have noun
behaviour (some grammarians classify such numerals as nouns) with
gender and declension of their own, which they preserve even in the
composition of the superior order numerals; these are, for instance,
suta( (hundred), mie (thousand), milion (million) and miliard
(billion). In a sentence most numerals may fulfill the function of
other parts of speech like noun, determiner or adverb.

10.2 Gender
Attribute           value                Ro. example
Gender              masculine (m)        doi, primul
                    feminine  (f)        doua(, prima
                    neuter (n)           (un) milion, (doua() milioane

10.3 Number
Attribute           value                Ro. example
Number              singular (s)         primul
                    plural (p)           primii

10.4 Case
Attribute           value                Ro. example
Case                direct(r)            primul
                    oblique (o)          primului

10.5 Form
Attribute           value                Ro. example
Form                digit (d)            1690
                    letter (l)           unsprezece
                    both (b)             5 mii
                    roman (r)            XIV

10.6 Definiteness
By virtue of their noun or adjective value, some numerals may take the
enclitic article (prim/primul - first/the first). Consequently for the
Romanian, definiteness attribute helps distinguish the enclitic forms
from the other forms.
Attribute           value                Ro. example
Definiteness        yes (y)              primul
                    no (n)               prim

10.7 Clitic
See discussion on clitics in 2.10.

Combinations
=======================
Tag           Example
=======================
Mcmprl        doi
Mcmpol        doi
Momsrl        doilea
Momsol        doilea
Mlmpr         ama>ndoi
Momsrlyy      primu-i
Mffpoly       treimilor

11. Interjections (I)
In Romanian there are no relevant  subcategories of interjections.
====================
Tag        Example
====================
I        oh,ah,au
====================

12. Residual (X)
No attributes are defined for this category.
==========================
Tag        Example
==========================
X        show, a+b, retro-
==========================

13. Abbreviation (Y)
The Syntactic_Type attribute is useful for specifying the grammatical
category of an abbreviation. Although the values for this attribute
could range over the part of speech categories in the language, in
Romanian most of the abbreviations falls into noun class.

13.1 Syntactic_Type

Attribute           value                Ro. example
Syntactic_Type      nominal(n)           d-na (doamna)
                    verbal (v)           v. (vezi)
                    adjectival (a)       ant. (anterior)
                    adverbial(r)         f. (foarte)

13.2 Gender
Attribute           value                Ro. example
Gender              masculine (m)        d-lui
                    feminine (f)         d-na
                    neuter(n)            apt.

13.3 Number
Attribute           value                Ro. example
Number              singular (s)         d-na
                    plural (p)           d-nele

13.4 Case
Attribute           value                Ro. example
Case                direct (r)           d-na
                    oblique (o)          d-nei

13.5 Definiteness
Attribute           value                Ro. example
Definiteness        yes (y)              d-nele
                    no (n)               d-ne

Combinations
======================
Tag           Example
======================
Ynmsry        d-ul
Ynfsoy        d-nei
Ynnsry        apt.

14. Particle (Q)

14.1 Type
Attribute           value                Ro. example

Type                negation(z)          nu,n-
                    infinitive (n)       a
                    subjunctive (s)      sa(
                    aspect               fi
                    future               o

14.2 Formation
This attribute is not irrelevant for Romanian.

14.3 Clitic
See  discussion on clitics in 2.10.
Attribute           value                Ro. example
Clitic              yes (y)              n-am
                    no (n)               nu am

Combinations
=================
Tag       Example
=================
Qz        nu
Qz-y      n-
Qn        a
Qs        sa(
Qa        fi
Qf        o

COP project 106 multext-East  Deliverable D1.1 F— Slovene  153

3.6 Application to Slovene

The application to Slovene has been elaborated by Tomaž Erjavec and Peter Holozan.

Acknowledgements: The authors thank Velimir Gjurin, France Žagar, Vladimír Petkevič, Lydia Sinapova and David Stermole for their much appreciated comments and suggestions. All errors of course remain our own.

All Slovene diacritical characters used have been encoded in the following way:

1. The Slovene ’hachek’ diacritic (ˇ ) is marked by the corresponding nondiacritic counterpart followed by ’<’. The following possibilities exist:

     c< s< z<
1. Noun (N)
1.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           common         c stol              obc<i samostalnik
                 proper         p Janez             lastni samostalnik
- -------------- -------------- - ----------------- -------------------------
2 Gender         masculine      m stol              mos<ki spol
                 feminine       f miza              z<enski spol
                 neuter         n sonce             srednji spol
- -------------- -------------- - ----------------- -------------------------
3 Number         singular       s stol              ednina
                 plural         p stoli             mnoz<ina
                 dual           d stola             dvojina
- -------------- -------------- - ----------------- -------------------------
4 Case           nominative     n stol,  medved     imenovalnik
                 genitive       g stola, medveda    rodilnik
                 dative         d stolu             dajalnik
                 accusative     a stol,  medveda    toz<ilnik
                 locative       l stolu             mestnik
                 instrumental   i stolom            orodnik
* ***************************** * ----------------- -------------------------
5 Definiteness                  -
- -------------- -------------- -
6 Clitic                        -
- -------------- -------------- -
7 Animate                       -
- -------------- -------------- -
8 Owner_Number                  -
- -------------- -------------- -
9 Owner_Person                  -
---------------- -------------- -
10Owned_Number                  -
=================================

1.2 Combinations
*** **** **** **** **** =====================================================
PoS Type Gend Numb Case Examples
*** **** **** **** **** =====================================================
N   p   [mfn]  s    n  Janez,      Micka,    Kosovo
N   p   [mfn]  s    g  Janeza,     Micke,    Kosova
N   p   [mfn]  s    d  Janezu,     Micki,    Kosovu
N   p   [mfn]  s    a  Janeza,     Micko,    Kosova
N   p   [mfn]  s    l  Janezu,     Micki,    Kosovu
N   p   [mfn]  s    i  Janezom,    Micko,    Kosovom
N   p   [mf]   p    n  Kristusi,   Slovenke
N   p   [mf]   p    g  Kristusov,  Slovenk
N   p   [mf]   p    d  Kristusom,  Slovenkam
N   p   [mf]   p    a  Kristusi,   Slovenke
N   p   [mf]   p    l  Kristusih,  Slovenkah
N   p   [mf]   p    i  Kristusi,   Slovenkami
N   p   [mf]   d    n  Kristusa,   Slovenki
N   p   [mf]   d    g  Kristusov,  Slovenk
N   p   [mf]   d    d  Kristusoma, Slovenkama
N   p   [mf]   d    a  Kristusa,   Slovenki
N   p   [mf]   d    l  Kristusih,  Slovenkah
N   p   [mf]   d    i  Kristusi,   Slovenkama
N   c    n     s    n  iskanje         ('gerund' classified a Noun)
N   c   [mfn]  s    n  stol,    lipa,   sonce
N   c   [mfn]  s    g  stola,   lipe,   sonca
N   c   [mfn]  s    d  stolu,   lipi,   soncu
N   c   [mfn]  s    a  stol,    lipo,   sonce
N   c   [mfn]  s    l  stolu,   lipi,   soncu
N   c   [mfn]  s    i  stolom,  lipo,   soncem
N   c   [mfn]  p    n  stoli,   lipe,   sonca
N   c   [mfn]  p    g  stolov,  lip,    sonc
N   c   [mfn]  p    d  stolom,  lipam,  soncem
N   c   [mfn]  p    a  stole,   lipe,   sonca
N   c   [mfn]  p    l  stolih,  lipah,  soncih
N   c   [mfn]  p    i  stoli,   lipami, sonci
N   c   [mfn]  d    n  stola,   lipi,   sonci
N   c   [mfn]  d    g  stolov,  lip,    sonc
N   c   [mfn]  d    d  stoloma, lipama, soncema
N   c   [mfn]  d    a  stola,   lipi,   sonci
N   c   [mfn]  d    l  stolih,  lipah,  soncih
N   c   [mfn]  d    i  stoloma, lipama, soncem
*** **** **** **** **** =====================================================

2. Verb (V)
2.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           main           m delati            glavni glagol
                 auxiliary      a imeti             pomoz<ni glagol
                 modal          o hoteti            naklonski glagol
                 copula         c biti              pomoz<nik 'biti'
- -------------- -------------- - ----------------- -------------------------
2 VForm          indicative     i delam             sedanjik
                 imperative     m delaj             velelnik
                 conditional    c bi                pogojnik od 'biti'
                 infinitive     n delati            nedoloc<nik
                 participle     p delal,(po)delan   delez<nik
                 supine         u delat             namenilnik
- -------------- -------------- - ----------------- -------------------------
3 Tense          present        p delam, delaj      sedanjik, velelnik
                 future         f bom               prihodnjik od 'biti'
                 past           s delal             opisni delez<nik na -l
- -------------- -------------- - ----------------- -------------------------
4 Person         first          1 delam             prva oseba
                 second         2 delas<            druga oseba
                 third          3 dela              tretja oseba
- -------------- -------------- - ----------------- -------------------------
5 Number         singular       s delam             ednina
                 plural         p delamo            mnoz<ina
                 dual           d delata            dvojina
- -------------- -------------- - ----------------- -------------------------
6 Gender         masculine      m delal             mos<ki spol
                 feminine       f delala            z<enski spol
                 neuter         n delalo            srednji spol
********************************* ----------------- -------------------------
7 Voice          active         a delam, delal      tvorni nac<in
                 passive        p (po)delan         trpni nac<in
- -------------- -------------- - ----------------- -------------------------
8 Negative       no             n ho"cem            nezanikan
                 yes            y no"cem            zanikan
- -------------- -------------- - ----------------- -------------------------
9 Definiteness                  -
- -------------- -------------- -
10Clitic                        -
- -------------- -------------- -
11Case                          -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
=================================

Notes:
1. The verb 'to be' in all its functions is Type=c;
  Auxiliary verbs Type=a do not include the 'to be' and modal verbs.
2. The "past participle" is actually used for making all the compound
  active tenses (future, past, pulperfect) and is encoded
  analytically as  Type=participle, Tense=past, Voice=Active.
3. The passive participles is encoded analytically as
  Type=participle, Tense='-', Voice=Passive.
  This encoding will only be used for the passive participle in the
  predicative position, e.g. 'on je bil tepen'/he was beaten. The
  case marked passive participles in the attributive position (e.g.
  'tepen pes'/beaten dog) or those in non-nominative case (e.g.
  'hrano imam skuhano'/I have the food cooked[acc]) will be
  classified as (qualitative) adjectives.
4. The adjectival (e.g. 'stokajoc<'/moaning) and adverbial (e.g.
  'lez<e'/lying down) participles are classified as adjectives and
  adverbs respectively.
5. Negative is always marked as 'n' except for two verbs:
  'noc<em'/to not_want, 'nisem'/to not_be

2.2 Combinations
*** **** **** **** **** **** **** ---- ---- =======================
PoS Type VFrm Tens Pers Numb Gend Voic Neg  Examples
*** **** **** **** **** **** **** ---- ---- =======================
V   c     n    -    -    -    -    -   -   biti
V   c     u    -    -    -    -    -   -   bit
V   c     c    -    -    -    -    -   -   bi
V   c     m    p  [12]  [pd]  -    -   -   bodimo, bodiva, bodite, bodita
V   c     i    f  [123] [spd] -    -   -   bom, bos<, bo, bomo, bova,...
V   c     i    p  [123] [spd] -    a [ny]  sem, si, je, smo, ...
V   c     p    s    -   [spd][mfn] a   -   bil, bila, bili, ...
V   o     n    -    -     -    -   -   -   hoteti
V   o     i    p   [123][spd]  -   - [ny]  hoc"ceva, noc<es, ...
V   o     m    p   [12] [pds]  -   -   -   hoti, hotimo, hotita, ...
V   o     p    s    -   [spd][mfn] a   -   smel, smela, smeli,...
V   o     p    -    -   [spd][mfn] p   -   dovoljen, dovoljeni, ...
V   m     n    -    -    -    -    -   -   delati
V   m     u    -    -    -    -    -   -   delat
V   m     m    p   [12] [pd]  -    -   -   dejaj, delajte, ...
V   m     p    -    -   [spd][mfn] p   -   dovoljen, dovoljeni, ...
V   m     p    s    -   [spd][mfn] p   -   dovolil, dovolili, ...
V   m     i    p   [123][spd] -    - [ny]  nimam, delam, delava, delamo
*** **** **** **** **** **** **** ---- ---- =======================

3. Adjective (A)
3.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           qualificative  f lep               kakovostni
                 possessive     s stric<ev          svojilni
           l.s.  ordinal        o slovenski         vrstni
- -------------- -------------- - ----------------- -------------------------
2 Degree         positive       p lep               osnovnik
                 comparative    c leps<i            primerjalnik
                 superlative    s najleps<i         presez<nik
- -------------- -------------- - ----------------- -------------------------
3 Gender         masculine      m lep               mos<ki spol
                 feminine       f lepa              z<enski spol
                 neuter         n lepo              srednji spol
- -------------- -------------- - ----------------- -------------------------
4 Number         singular       s lep               ednina
                 plural         p lepi              mnoz<ina
                 dual           d lepa              dvojina
- -------------- -------------- - ----------------- -------------------------
5 Case           nominative     n lep, lepi         imenovalnik
                 genitive       g lepega            rodilnik
                 dative         d lepemu            dajalnik
                 accusative     a lep, lepega       toz<ilnik
                 locative       l lepemu            mestnik
                 instrumental   i lepim             orodnik
* ***************************** * ----------------- -------------------------
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Animate                       -
- -------------- -------------- -
9 Formation                     -
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
---------------- -------------- -
12Owned_Number                  -
=============================== =

Notes:
The three deverbative adjectival participles (part. -l, part -n/t,
part. c<) are also classified as qualificative adjectives.  e.g.
'otekel (obraz), premagan, (sovrag) govorec< (pes)' are marked Afpmsn.

3.2 Combinations
*** **** **** **** **** **** ================================================
PoS Type Degr Gend Numb Case Examples
*** **** **** **** **** **** ================================================
 A   s     p [mfs][spd][ngdali] stric<ev, galjotove, ...
 A   o     p [mfs][spd][ngdali] dlakocepski, slovenskega, ...
 A   f  [spd][mfs][spd][ngdali] lep, lepi, lepa, leps<i, najleps<i, ...
*** **** **** **** **** **** ================================================

4. Pronoun (P)
4.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           personal       p jaz               osebni
                 demonstrative  d ta                kazalni
                 indefinite     i nekdo             nedoloc<nostni
                 possessive     s moj               svojilni
                 interrogative  q kdo               vpras<alni
                 relative       r kdor              oziralnostni
                 reflexive      x se, svojih        povratni
                 negative       z nihc<e            nikalni
                 general        g vsak              celostni
- -------------- -------------- - ----------------- -------------------------
2 Person         first          1 moj               prva oseba
                 second         2 tvoj              druga oseba
                 third          3 njegov            tretja oseba
- -------------- -------------- - ----------------- -------------------------
3 Gender         masculine      m moj               mos<ki spol
                 feminine       f moja              z<enski spol
                 neuter         n moje              srednji spol
- -------------- -------------- - ----------------- -------------------------
4 Number         singular       s moj               ednina
                 plural         p moji              mnoz<ina
                 dual           d moja              dvojina
- -------------- -------------- - ----------------- -------------------------
5 Case           nominative     n moj               imenovalnik
                 genitive       g mojega            rodilnik
                 dative         d mojemu            dajalnik
                 accusative     a moj/mojega        toz<ilnik
                 locative       l mojemu            mestnik
                 instrumental   i mojim             orodnik
- -------------- -------------- - ----------------- -------------------------
6 Owner_Number   singular       s moj               ednina svojine
                 plural         p nas<              mnoz<ina svojine
                 dual           d najin             dvojina svojine
- -------------- -------------- - ----------------- -------------------------
7 Owner_Gender   masculine      m njegov            mos<ki spol svojine
                 feminine       f njen              z<enski spol svojine

                 neuter         n njegov            srednji spol svojine
********************************* ----------------- -------------------------
8 Clitic         no             n mene              beseda
                 yes            y me                naslonka
- -------------- -------------- - ----------------- -------------------------
9 Referent_Type  personal       p sebe, se, si      osebni povratni
                 possessive     s svoj              svojilni povratni
- -------------- -------------- - ----------------- -------------------------
10Syntactic_Type nominal        n kdo               samostalnis<ki
                 adjectival     a kateri            pridevnis<ki
- -------------- -------------- - ----------------- -------------------------
11Definiteness                  -
- -------------- -------------- -
12Animate                       -
- -------------- -------------- -
13Clitic_s                      -
- -------------- -------------- -
14Pronoun_Form                  -
- -------------- -------------- -
15Owner_Person                  -
- -------------- -------------- -
16Owned_Number                  -
=================================
Notes:
1. The Type taxonomy for pronouns different from the one of Slovene
   grammars. Mostly Sl. grammar classes were merged. In particular,
   the indefinite and relative types contain many of pronoun types
   from Slovene grammars:
   (nedoloc<nostni = nedoloc<ni+poljubnostni+mnogostni+istostni+drugostni)
   (oziralnostni = oziralni + oziralno poljubnostni)
2. Type=relative includes so called 'relative general' (oziralno poljubnostni)
   pronouns. They can consist of two words, where the second is 'koli'.
   Variant spelling exists: 'kolikorkoli', 'kolikor koli'.
   Possible solution: they can be treated as compounds.
3. Type=reflexive ecompasses all reflexive pronouns (sebe, se, sebi,
   si, svoj) as well as 'se' and 'si' in all their roles.
   Personal and possessive reflexives are further distinguished via
   the Referent_Type attribute.
   'se' and 'si' in their roles as an obligatory but dislocated
   constituent' of reflexive verbs have the unique MSD Px------y.
   Note: 'se' can also appear as a suffix: e.g. 'zase'/for_oneself with
   variant spelling 'za se'.
4. Owner_Gender is relevant only for third person singular
   possessive pronouns, and distinguishes only masculine and feminine
   forms. It will be '-' in all other cases.
5. As in Slovene grammars, pronouns are distinguished between having a nominal
   and adjectival function. All pronoun types except demonstrative and
   possessive can be nominal, and all except personal can be
   adjectival.
   In the future version of the specifications, this attribute should
   be given a lower position, as it is used by three languages.

6. Referent_Type is used to distinguish personal reflexives (which include
  'se' in all its functions) from the possesive reflexives ('svoj').

4.2 Combinations
*** **** **** **** **** **** **** **** ---- --- --- ================
PoS Type Pers Gend Numb Case PosN PosG Cltc Ref Syn Examples
*** **** **** **** **** **** **** **** ---- --- --- ================
P    x    -    -    -    -    -    -    y   -   -   se, si
P    x    -    -    -  [gdali]-    -  [ny]  p   n   sebe, se, sabo,...
P    x    - [mfn][spd][ngdali]-    -    -   s   a   svoj, svoja, ...
P    p  [123]  - [spd][ngdali]-    -  [ny]  -   a   njega, ga, jih, ...
P    s  [12][mfn][spd][ngdali][spd]-    -   -   a   moj, vajinima, ...
P    s   [3][mfn][spd][ngdali]s  [mfn]  -   -   a   njegov, njen, njegova,...
P    s   [3][mfn][spd][ngdali][pd] -    -   -   a   njihova, njuna, ...
P    d    -    -    -    -    -    -    -   -   a   le-tega, onega, ...
P    d    - [mfn][spd][ngdali]-    -    -   - [an]  le-ta, le-tega, onega, ...
P    i    -    -    -    -    -    -    -   -   a   koliko, c<igar
P    i    -    -    - [ngdali]-    -    -   -   n   kaj, marsikaj, ...
P    i    - [mfn][spd][ngdali]-    -    -   - [an]  en, tegale, drug, ...
P    q    -    -    -    -    -    -    -   -   a   koliko
P    q    -    -    - [ngdali]-    -    -   -   n   kaj, c<cesa, kdo, ...
P    q    - [mfn][spd][ngdali]-    -    -   -   a   kaks<sen, kateri, ...
P    r    -    -    -    -    -    -    -   -   a   kolikor, c<igar
P    r    -    -    - [ngdali]-    -    -   -   n   kar, c<esar, kdorkoli, ...
P    r    - [mfn][spd][ngdali]-    -    -   -   a   kakrs<en, kateri
P    z    -    -    -    -    -    -    -   -   a   nic<, nikogar, noben
P    z    -    -    - [gdali] -    -    -   -   n   nikogar, nikomur, ...
P    z    - [mfn][spd][ngdali]-    -    -   -   a   nikakrs<en, noben
P    g    -    -   -     -     -   -    -   -   a   vsakogar
P    g    -    -   -  [ngdali] -   -    -   -   n   vsakdo, vsakogar, ...
P    g    - [mfn][spd][ngdali] -   -    -   -  [an] ves, vsega, vsa, vsak, ...
**** **** **** **** **** **** **** **** --- --- --- ================

5. Determiner (D)
Not applicable.

6. Article (T)
Not applicable.
Notes:
This category (Slovene: 'c<len') exists for Slovene, but it
encompasses only two words ('ta'/this, 'en'/one) and even these are
used only colloquially.  Furthermore, the context in which these two
words can appear as articles is identical to the one in which they
appear as a pronoun or numeral respectively. As it would be impossible
to disambiguate between their two meanings, this category will not be
used for Slovene.

7. Adverb (R)
7.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           general        g potihoma          prislov
- -------------- -------------- - ----------------- -------------------------
2 Degree         positive       p malo              osnovnik
                 comparative    c manj              primernik
                 superlative    s najmanj           presez<nik
********************************* ----------------- -------------------------
3 Clitic                        -
- -------------- -------------- -
4 Number                        -
- -------------- -------------- -
5 Person                        -
=================================
Notes:
1. The adverbial (e.g. part -e: 'lez<e'/lying down)
   participles are classified as and adverbs.
2. Some pronouns can also have an adverbial function. This issue has not
   been considered here.

7.2 Combinations
*** **** **** ==============================================================
PoS Type Degr Examples
*** **** **** ==============================================================
 R    g    p  bljuvaje
 R    g    c  kasneje
 R    g    s  najprijazneje
*** **** **** ==============================================================

8. Adposition (S)
8.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           preposition    p nad               predlog
- -------------- -------------- - ----------------- -------------------------
2 Formation      simple         s nad, pod          (enostaven) predlog
                 compound       c zase, nanj        predlog s pripon. zaimkom
********************************* ----------------- -------------------------
3 Case
  (req.by prep.) genitive       g brez              rodilnik
                 dative         d k                 dajalnik
                 accusative     a po                toz<ilnik
                 locative       l pri               mestnik
                 instrumental   i s                 orodnik
- -------------- -------------- - ----------------- -------------------------
4 Clitic                        -
=================================
Notes:
Slovene has prepositions only, with some arguable exceptions
(e.g. 'navkljub'/in spite of), which can be pre- or postpositions.
These would need further study.

8.2 Combinations
*** **** **** ---- =========================================================
PoS Type Form  Case Examples
*** **** **** ---- =========================================================
 S    p    c    -   zase, nase, nanj, nanjo, nanju, ...
 S    p    s    g   brez, izpred, od, ...
 S    p    s    d   k, navzlic, ...
 S    p    s    a   s, po, skozi, ...
 S    p    s    l   po, na, v, ...
 S    p    s    i   pod, s, med, ...
*** **** **** ---- =========================================================

9. Conjunction (C)
9.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           coordinating   c in                priredni veznik
                 subordinating  s da                podredni veznik

********************************* ----------------- -------------------------
2 Formation      simple         s in, da, ...       enodelni, dvodelni
                 compound       c kljub_temu_da     vec<besedni
- -------------- -------------- - ----------------- -------------------------
3 Coord_Type                    -
- -------------- -------------- -
4 Sub_Type                      -
- -------------- -------------- -
5 Clitic                        -
- -------------- -------------- -
6 Number                        -
- -------------- -------------- -
7 Person                        -
=================================
Notes:
1. According to Slov. grammar, some conjunctions are 'two-part', but these
   can often be either single or two-part conjunctions with identical first
   and second conjunct, e.g. 'ali ... ali'; such conunctions are treated
   as ordinary, 'one-part' conjunctions, i.e. as Formation=simple.
2. Conjunctions are also classified into 'multi-word' conjunctions,
   e.g. 'kljub temu da'/in spite of. These conjuctions constitute one
   lexical 'word'; should be merged by the text tokeniser e.g.:
   kljub_temu_da and marked as Formation=compound.

9.2 Combinations
*** **** ---- ===============================================================
PoS Type Form Examples
*** **** ---- ===============================================================
 C    c    s   in, ali
 C    s    s   da, ki, ali
 C    s    c   medtem_ko
*** **** ====================================================================

10. Numeral (M)
10.1 Lexicon
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type           cardinal       c en                glavni s<tevnik
                 ordinal        o prvi              vrstilni s<tevnik
                 multiple       m enojen            mnoz<ilni s<tevnik
                 special        s dvoje, ...        ostali s<tevniki
- -------------- -------------- - ----------------- -------------------------
2 Gender         masculine      m en                mos<ki spol
                 feminine       f ena               z<enski spol

                 neuter         n eno               srednji spol
- -------------- -------------- - ----------------- -------------------------
3 Number         singular       s en                ednina
                 plural         p tri               mnoz<ina
                 dual           d dva               dvojina
- -------------- -------------- - ----------------- -------------------------
4 Case           nominative     n en                imenovalnik
                 genitive       g enega             rodilnik
                 dative         d enemu             dajalnik
                 accusative     a en / enega        toz<ilnik
                 locative       l enemu             mestnik
                 instrumental   i enim              orodnik
********************************* ----------------- -------------------------
5 Form           digit          d 1984              arabska s<tevilka
                 roman          r MCMXXCIV          rimska stevilka
                 letter         l tisoc<devetsto    s<tevnik
- -------------- -------------- - ----------------- -------------------------
6 Definiteness                  -
- -------------- -------------- -
7 Clitic                        -
- -------------- -------------- -
8 Class                         -
- -------------- -------------- -
9 Animate                       -
- -------------- -------------- -
10Owner_Number                  -
- -------------- -------------- -
11Owner_Person                  -
- -------------- -------------- -
12Owned_Number                  -
= ============== ============== =
Notes:
Numerals in Slovene can function as nouns, adjectives or adverbs, and
are in grammars described as subtypes of these categories. The above
classification runs counter to the established practice and is missing
an important syntactic distinction.

10.2 Combinations
*** **** **** **** **** ---- ===============================================
PoS Type Gend Numb Case Form Examples
*** **** **** **** **** ---- ===============================================
 M    -    -    -    -    d  1984, ...
 M    -    -    -    -    r  MCMXXCIV, ...
 M    c [mfs][spd][ngdali]l  en, eden, dva, trije, s<tirje, pet, ...
 M    o [mfs][spd][ngdali]l  prvi, drugi, tretji, c<etrti, peti, ...
 M    m [mfs][spd][ngdali]l  enojen, dvojen, trojen, ...
 M    s [mfs][spd][ngdali]l  dvoje, troje, dvoj, dvakraten, ...
 M    c [mfs][spd][ngdali]l  en, eden, ena, eno, trije, tri, troje, dva, ...
*** **** **** **** **** ====================================================

11. Interjection (I)
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type                          - ah, aha, amen,    medmet
- -------------- -------------- - ----------------- -------------------------
2 Formation                     -
=================================

12. Residual (X)
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
P ATT                           - sic, 2+2=5, ...   Ostanki
=================================

13. Abbreviation (Y)
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Syntactic_Type                - dr., ipd, ZSMS    okrajs<ava
- -------------- -------------- - ----------------- ------------------------
2 Gender                        -
- -------------- -------------- -
3 Number                        -
- -------------- -------------- -
4 Case                          -
- -------------- -------------- -
5 Definiteness                  -
=================================
Notes:
Abbreviation often functions as a proper noun or adjective (NATO, Elra).

14. Particle (Q)
= ============== ============== = ================= =========================
P ATT            VAL            C Example           Slovene term
= ============== ============== = ================= =========================
1 Type                          - spet              c<lenek
- -------------- -------------- - ----------------- -------------------------
2 Formation                     -
- -------------- -------------- -
3 Clitic                        -
=================================

References

Bel N., Calzolari N. and Monachini M. eds. (1995): “Common Specifications and Notation for Lexicon Encoding and Preliminary Proposal for the Tagsets”, MULTEXT Deliverable D1.6.1B, Pisa.

Leech J. and A. Wilson (1994): “Morphosyntactic Annotation”, EAGLES Interim Report, Pisa.

Monachini M. and N. Calzolari (1995): “Synopsis and Comparison of Morphosyntactic Phenomena Encoded in Lexicons and in Corpora and Application to European Languages”, EAGLES document EAG-LSG-T4.6/CSG-T3.2, Pisa.

MULTEXT (1993): “MULTEXT Technical Annex”, Aix-en-Provence.

MULTEXT WP1.6 Report A2 (1994): “Common Specifications and Notation for Lexicon Encoding”, MULTEXT Report Milestone A2.